A Complete Bibliography of Publications in the Journal of Systems and Software (1979–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

29 August 2024 Version 3.01

Title word cross-reference

+ [WH99]. ^{,} [FW90]. β [LM94]. F [CJP98]. K [Nei97, DT90, DS94]. L [VH89]. N [EL88, Pha94]. R [Kor99b]. Z [FF96, FRF98].

-Coteries [Nei97]. -Distributed [LM94]. -machine [VH89]. -metric [CJP98]. -modular [EL88]. -nets [Kor99b]. -trees [WH99]. -Version [Pha94, EL88].

/M/1/Fifo [MR86].

1 [Bel91, Lit90, WL99]. 1-2-3 [Lit90]. 1679 [Coo81]. 1980s [Gla92d]. 1990s [Ano92f, Gla92e]. 1996 [BT97].

2 [WM90]. 2167A [Wal91].

3 [Lit90].

4GL [Ano87e, Dol97]. 4GLs [Gla91a].

’99 [LS99].

Abbott [BYY87]. Abstract [Bel91, Car96, Vla98]. Abstraction [BW83, MM81, MG81, Nit96, Nit98]. Abstractions [How80, VP92, KB98]. Abuse [Got92b]. Academic [BHR89, Lai99]. Academic/Industrial [BHR89]. Acceptance [Gla93d]. Access [CH83, Hać86a, HB83, Hen95, MO90, UH86, Ulu97]. Accumulation [Tör90]. achieve [Ano87f, RVM99]. Achieving [Bol97a, KWME99, Lam97, Ber94]. Acquisition [SL96, Tar92, Eri92, Ozk97]. across [FF95]. ACSE [Lai95, LL97a]. action [Rom98]. Active [PW87, WNHM86, BG98, ÇZUB99, DMV98, KR98, yL98, MA94, SBB98, Ulu98]. Actor [RMC93]. Actor-Based [RMC93]. Actors [Chu97]. ADA [Bel91, CW90, Bak88, CT94, Coh81, DS92, Dil91, EOM95, MA89, PW92, Rom98, SC88, Tom89, Wal91, Whe81]. Ada-like [CT94]. Adaptation [PW92, WKM94]. Adaptations [GRS92]. Adapting [CR89]. Adaptive [Blu93, CW97, SF92, LC98]. Adding [RUV92]. Addressing [Jef96, WJ99, HR95]. adhocracy [Ano93f, Gla93b]. Adjustment [UH96]. Administration [DR84]. Adopting [Har97]. adoption [BM89, WD99]. Advanced [GK91b]. Advantage [Gla97f]. advantages [VC97]. advice [KJ99]. aerospace [Tha80]. affecting [DLT99]. Affinity [WWC97]. Affirm [MD81]. after [Gla98f]. Age [Rei87]. Agency [FK92]. agent [PA99, Shu99]. agent-based [Shu99]. agents [WM99]. Aggregate [Mot96]. AIDA [EOM95]. Aided [Amm91, BTT84, Mey88a]. Aisim [Mer87]. Albrecht [Dol97]. Algebra [Bra96, MR84]. Algebraic [BCFG86, DGM93, KH96, Krä91b]. Algorithm [AT97, CBK96, DS94, Fra86, HJ90a, HJ91, Hać93, HL94a, Hać94, Hen88, MTW97, NM93, RY93, Tan96, TW95, WTS95, Cha93, JM96, Ng99, TW98]. Algorithms [BP86, Bha84, CZ91, DHP86, FG93, Fis91, HJS91, Hać91, MS97, MPS86, MS90, Sca88, Sta90, TM98, Tri86a, Ver89, YP94, ZR87, Amm89, BdADH94, CDS99, DLT99, Hać92, KH97, MC98]. All-uses [FWH97]. alleviating [WWC98]. Allocating [ML95, TC93]. Allocation [AH81, Cho95, Rah92, SG89, KKiMT96, Leu97, PM99, TM98]. alone [ST89]. Alterations [SB95]. alternating [GAK92].

Alternative [KML94]. Alvey [Qui94].

American [Gla97f]. Among

[HB83, LIC92, Sho91, CLLC96, Dol97]. Analyses [Eva97, OKOM97]. Analysis [AS96, Ber93, BC91, CH94, CR90, CMP85, DLG96, DG80, EHS93, Eli92, FM93, FN85, FAI94, FAI97, Gla92a, GDF86, Hag91, Har88b, HH87, HB89, IMM95, Jef91, JP94, JO83, KSH92, Ken84, KS96, KP93, KP97b, Lak93, LL97b, LH83, Liu93, MTG92, McF92, MW95, MM93c, Mue86, Myr90, Nel81, Par86, PdF97, PH86, RCSD93, Sca88, Sch91, Sel93, SB95, TOYI95, Tia96, TTP97, Tsu85, WH91a, WH91b, WKM94, YNDS88, ZEB88, ZX94, vdBK94, CLL99, CJP98, CKyL98, DS98, FTAM99, GAK92, Hać92, JM96, Kor99a, LHC95, LM96, Lut96, MR99, NS92, Row86, SC88, Wei79]. Analysts [TDB97]. Analytic [BDM^{+}93, FSA87, FWP93, Lee93]. analytical [RST98]. Analyzer [FLN91, BB89, EOM95]. Analyzing [His98, MW95, Mot96, Sta90]. and-or [Vla98]. Anecdote [Gla98a]. animation [MSHB98]. Annotated [Not85a]. announcements [Ano92d]. Anomalies [WKM94]. Anomaly [CA87b, CA87a, Hsi91b, KMM89, Zwe90]. anonymous [Gla95g]. answers [VL94]. anything [Gla97g]. Apollo [BP86]. appear [Gla96h]. Applicability [WH91a]. Application [AF96, BFG97, BYY87, Ber93, BL98, CLO95, DK94, EHS93, Fri90, Gla92f, GV92, Gla99c, HL94a, HZ84, HS95, Lan98b, LVB^{+}93, MM93a, MB84, Nit98, SL96, TZ92, TM97, YN91, ZC97, Ano92g, CTZ92, Dav99, DFCR96, FJ98, HBJ^{+}99, WHMP99]. Application-Specific [DK94]. Applications [Ano86d, CR85, Goe80, Gom89, HH97, HFK92, KP97a, MD91, MK90, Sta93b, Zho94, DY99, ISS98, JCJ99, KHL^{+}99, Ozk97, PL94, SPK99, SBB98, TL99]. Applying [BS93, KS96, LL98, Moo98, Rog89, Ano93e].

Approach [AQ90, Bar92, BW83, BAH96, BST93, CB89b, Car96, DA86, DK97, DLS94, Dil91, Dye87, Dye93, Fra90, HZ84, HOT97, JvB83, KL95, KAL97, KSW93, KCK^{+}98, Lam97, Lan98a, LF96, Mai96, MC91, MWH98, MR83, Mue86, NS83, PM90a, PdF97, Pow86, Rv91, Rv93, She90, SCK95, SCK86, Tia96, TM97, WLPL95, WWF94, GP98, HK98, Has98, JF99, JMM99, Kam89, LM96, Liu98, LJ99, LJM96, Mur99, PL94, PMB99, PP94, SMCL96, WC99]. Approaches [KO95, KML94, RBCM91, VP92]. appropriate [Ozk97]. approximate [VL94]. Approximation [MR86]. Approximations [vD93]. April [BT97]. Arabic [AA98]. Architectural [Lea95]. Architecture [Amb87, DY99, HJ90b, IMM95, JO83, KP97a, RC89, TL96, PM94, SMS94, Vla98, Wei79]. Architecture-driven [DY99]. Architectures [AT97, CFK91, Gom95, Ulu95, RR98]. Areas [Gla92a]. ARP [SSK98]. array [MJ89]. array-theoretic [MJ89]. Arrays [RT86, Row86]. Arrivals [BFC92]. Art [FG94, Hat99, Sto92]. article [Ano87d]. Artifact [HMG96]. artificial [PP94]. ARTS [DF84]. Asia [Zuc90a, Rei90b]. ASN.1 [LL99]. Aspects [Sah94, HL94b]. Assembly [BK85, SW88]. Assertion [YRN80]. Assertions [FAS94, JM90, MGJT87]. Assess [KK81]. Assessing [GC94, Vis99a, Liu98]. Assessment [Cav84, Gla94a, Gla96a, Gla97a, JM90, Pre95, VP92, Vis99b, Gla95c, Gla98b, Gla99a, Gla99b, Gla00, MR99, NL99]. ASSET [OBS79, Rei90a]. Asset-R [Rei90a]. assignment [VVS99]. Assistance [GK91b]. assistant [WT89]. Assisted [Bar92, GH83]. Associative [Hsi91a]. assurance [CW89, Iso98, NDM80, RST98]. Assuring [Hon90]. Asynchronous [FC96, FG93, Hać91, HMG96, KM92]. ATM [SSK98]. ATNet [BMSB94]. Atomic [WM96, Rom98]. Attack [DG87]. Attitudes [TKS95, JH99]. Attribute [Mot96]. Attributes [GR97]. Audits [Ber81]. Australia [CFSS98]. authentication [JC98]. Author [Ano80a, Ano81a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89a, Ano89b, Ano90a, Ano91a, Ano92a, Ano93a, Ano94a, Ano95a, Ano96a, Ano97a]. Authorization [FM87]. Automata [SP94]. Automated [Arc81, Ber91, DF84, FAS94, Mer87, NS83, PBC93, Rec93, SKF95, SC88]. Automatic [AM85, DGM93, Gla90b, Hab85, DF98, Gla97i, JF99, Lai95, LL99]. Automating [SG89]. Automation [BMP97, HZ83, Gla95h]. autonomous [WM99]. Autonomy [Lue92]. Availability [Tsu85]. Available [LS97]. Avionics [Lam97]. avoiding [PV94]. Axis [Sah94].

B [WH99]. Back [Bux90, ZK85]. Back-End [ZK85]. Background [Sei89]. Bad [BAH96, Gla89h]. Balancing [HJ90a, HJ91, RCSD93, DLT99]. BaLinda [YFY96]. Band [RT86]. bandwidth [Ng99]. Base [GRS92, GSC91, MP90, RT93, Won93, HR95, HL94b, PL94, PM94, RC89, SW96, ZS88]. Base- [MP90]. Based [Art87, Bar86, Car96, DS92, DK94, Dye93, EL94, FM90a, FWD97, Fra90, HL93, HFK92, KH81, KB96, Ken84, KAL97, KN97, Lan98a, LL97b, Liu93, MG81, MW95, MSI90, MIH92, Mos84b, MP90, MP95, NM93, OG80, PBC93, PdF97, Pre95, RY93, Rv93, Ros87, RMC93, SGL93, Sam93, STJ83, She90, Tak97, TW95, YY93, ZX94, AH88, AH90, Amm89, BK95, Çam99, CJP98, CKyL98, Chr99, De 97, DB95, Dil91, Fic89, FPW96, Gie79, Gla95i, HHZ92, HL94b, HHL^{+}97, JVP^{+}98, KWME99, KSFT89, Kuo94, LL99, Moo98, MHSM99, Mos84a, NL99, Nec96, Pla95, PA99, PS90, PP94, Shu99, WSM^{+}95, ZLZ^{+}96]. Bases [KZ91, Uck91, BF96, MP94, MA94, She89]. Basic [Boe83, GMP94, De 97]. Basis [Lea95, McF92, WM90, RG79, TFLW99]. Battle [RB93a]. Be [Gla91h, Mat86, Gla96h, ZCd96]. become [Gla89c]. becoming [Gla92c]. beginning [Gla98k]. Behavior [MD91, Nit96, Sak84, WSR^{+}83, SC88, War89]. Behavioral [BW83, Mil96a, OHBR90]. Behaviors [FZHS95]. Behaviour [Nit98]. Belbin [HS99]. believe [FF89]. Benchmark [Haç89a, ÇZUB99]. Bend [Gla96f]. benefits [BB89]. best [Gla91b, Gla92d]. Better [Gla93i, Gla98d]. Between [Chr86, HD84, Lan90, Gla89i, Gla92g, HZ79]. Beyond [Gla95d]. Bibliography [Not85a]. BIDDLE [YY93]. Binding [Gan91, GJ88, ZS88]. Biographies [Ano81b, Ano81c, Ano84b, Ano85b, Ano87b, Ano87c, Ano88b, Ano88c, Ano89c, Ano89d, Ano89e, Ano89f, Ano90b, Ano90c, Ano92b, Ano92c, Ano94b, Ano94c, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano98a, Ano98b, Ano98c]. Biography [Ano79, Ano80b, Ano80c]. Blit [Car83]. Block [HL83]. blood [Ken80]. Blueprints [SG91]. BOOM [RA96, Gla97d]. Born [CHB94]. Boston [Bra89]. bounds [PNK96]. Box [HZ84, WL99]. Boy [Gla90g]. brainstorming [Gla97h]. Branch [Ber93, BM96, BMP97]. Break [Spa92]. Break-ins [Spa92]. Breakdown [Tau80]. brightest [Gla91b]. brilliant [Gla89d]. Broadcasts [Ram90]. Broadening [McF92]. Brooks [Ano87d]. BSD [WLC95]. Budget [Leu92]. Buffer [DMV98, Hać91, CB89a]. Buffered [MF90]. Build [Pfl95]. Building [Bar92, HL94a, HO96, Hać92]. builds [FW90]. Built [Gla89f]. Built-up [Gla89f]. bullet [Ano87d]. Bus [GDF86, MBCD86]. Business [DLG96, HH97, TL96, WM90, DIP98, GV99, MHSM99]. bust [Gla97d]. Buyer [Hon90].

C [AP97, CDM98, CMP85, DJL93, FLN91, LH98, Lok96]. Cache [Hać91]. Calibrating [Gul91]. Call [Ano93b, Ano93c, Ano93d, Ano96m, ZM96, CV95, Gla95g]. called [Gla89d]. Calls [Ano92d]. CAMS [LJM96, SGL93]. Can [BB81, Gla91h, Mat86, Gla89c, Gla98d, Gla93a]. Candidate [BC94]. Capabilities [MR84, Zel96]. capability [GAW92]. capacity [VVS99]. capture [Iso98]. capture-and-recapture [Iso98]. Cards [Bri92, JT97]. Cares [Gla98j]. Case [AH90, BMP97, DGM93, EC98, Eli92, FWD97, Gla96j, Gor91, JVP^{+}98, PW92, Rv92, Sed93, SW94b, SB88, Wic92, AH88, Bar94, BP80, DFCR96, HLAB99, KC98, Lin99, MT98, RR98, Sal80, SSvdW99, War89, WHMP99, APL95, Gla91a, IYKO95, IKCN91, PC98a, PKK98, RBM95, TM97]. Case-Based [FWD97]. case-supported [Bar94]. cases [NS92]. casual [RB99]. Catastrophic [DG92]. categorization [GKP98]. Categorizing [OW84]. cause [Gla96h]. Caused [BAH96, FAI97]. Causes [LP95]. CD [Lai97c]. Centered [FG94]. centric [LS99]. century [Gla99d]. Cépage [Mey88a]. Certification [WH97]. Challenges [JTW98]. Change [HR96, LS98, Cha95]. CHANGE-POINTs [Cha95]. changes [Ber98]. changing [CR89]. Chaos [MTW97]. character [AA98]. Characteristics [CM93, HBJ^{+}99, TT98]. Characters [CW97]. Charts [BCD92]. Chasm [Gla91h]. cheater [HLC99]. Checking [Kun91a]. Checkpointing [YP94]. Checksum [Bro87]. cheek [Gla91g]. Chinese [CW97, Gla90g]. choosing [CTA94]. Chorus [Ban86]. Christian [CR89]. chunk [Hsi91b]. CICS [FSA87, FF96]. Circuit [PH93]. Circuit-Switched [PH93]. CIS [Gil88]. claims [FF96, Gla96g]. Clarifying [Gla91c]. clarity [LJ99]. Class [CH94, MBCD86, MM93b, Rom99]. Classes [BBG86, Hać88, LH98]. classical [SSK98]. Classification [Esk89, Lak97, PS90, Tak97, Tri86a, JCJ99, Kam95, SH98, ZMAER99]. Classifying [WWC98]. Classroom [MC91]. Cleaning [CC99b, Gla98f]. Client [Gla97d, SMS94]. client-server [SMS94]. Client/server [Gla97d]. close [Gla95a]. Closed [MR86]. Closely [HJ90b]. Closely-Coupled [HJ90b]. Closing [CFSS98]. Closure [Fra86]. Cluster [Gla92f, Ano92g, Shu99]. Clustered [WWC97]. Clustering [BP91, MW95, RY93]. CMM [Chr99, RVM99]. CMM-based [Chr99]. COBOL [AP97, Ano87h, BB89, Gla97b]. Cocktail [Gla90b]. COCOMO [Gul91]. Code [AC97, AF96, BAEH96, CR90, Dol97, Kal92, Lak93, Lue92, OC90, BM98, CDM98, Gla97i]. Cognitive [AS96, Let87, SFM99, KC98, LJ99]. coherent [PM94]. Cohesion [Dha95]. collaboration [BHR89, GAWC91, GAW92]. collaborations [MBL^{+}99]. collected [Mar81]. Collecting [OW84]. Collection [BBC^{+}88, YNDS88, Yua90]. Collective [SM92b]. Colored [SBM94]. Combinators [SD94]. Combining [HK98]. Coming [Fis81]. Commanders [Sch81]. Comments [CA87a, DF98]. Commit [WM96]. committee [Gla96i]. Common [AJMP96, GDF86, MP90]. Communicating [GHC91, Jia99, Moy96, Lai99]. Communication [ASSA96, Gla91h, Jma96, LL98, MF90, MWH97, MWH98, YCGH92, KH97, Rog89, WWC98]. Communications [Mor86, JS99]. Community [AM94]. Commutative [Hsi91a]. Company [Sed93].

Comparative

[GKP98, Gla92a, MRW^{+}94, PT91, TOYI95, Wil89, GAK92, Kam95, PKK98]. compared [Lit80]. Comparing [MF90, Mos84b, Mos84a]. Comparison [Bla87, DHP86, FWD97, HJ90b, HG91, Moy96, Ver89, FWH97, MO84, TT98]. Competencies [TB95]. Compilation [Fri83, HL94a]. Compiler [Ros87]. Compilers [Mos84b, Mos84a]. completeness [RPL97]. Complex [Dam96, PdC94, PdF97, Sca88, Gie79, Lai97d, NC88, SW95a]. Complexity [AR90, BK85, CS85, DS92, Eva83, Gon95, HC87, HS95, HB89, HL98, KML94, Mac91, MTG92, MM92, MK90, MK93, Rey84, Tak97, TZ92, Zei88, CA88, MT98, Moo98]. Component [BDM^{+}93, FPW96]. component-based [FPW96]. Components [BAEH96, DJL93, Eva97, TL96]. Composite [DGS88, HS95]. composition [SZ98]. Compositional [UH96]. Comprehending [Sca88]. comprehensibility [FRF98]. Comprehension [AS96, BBP96, Let87, RBCM91]. Comprehensive [Zvi93, KR98]. Compression [Cha91, CBK96, CW97]. Computation [KD91, MJ89]. computational [Vla98]. Computations [AQ90, BFR96, BP91, WWC98]. Computed [DS98]. Computer [Amm91, Bar92, BTT84, BLPB92, CZ91, CM92, DG87, DV94, DHP86, FM90b, FS91, FK92, GK91a, GK91b, Gla90g, Gla92a, Gla96c, Gla97e, Got92a, Got92b, Hay86, Kal92, KL90, KNT86, LIC92, Lue92, MC91, Mat86, MvS95, RA91, SL80, Sch81, Spa92, TLPH95, YN91, Zvi93, Fle95, FF89, Gie79, Gla89c, Gla96e, Har98, Mar81, Mey88a, ST89, Sny79, TVK94, TVK95, VB99, WSM^{+}95]. computer-aided [Mey88a]. Computer-Assisted [Bar92]. computer-based [WSM^{+}95]. Computers [IMM95, OS87, SM92b, CC99b]. Computing [Eng81, FK92, Gla91h, KN97, Mor86, Pow86, Rv91, Rv92, Rv93, Sch97, ZR94, BS96, Gla95i, Mar81, SPK99, Shu99, TE99, TW98, TM98]. Concept [FS91, Gla89d, HLC99]. Concepts [CHB94, Sku91, TKS95, JNY84, SPK99, Rog94]. Conceptual [BF90, Del92, FM87, Kun95, RA91, Sak84, DB95, RB99, TFLW99]. Concurrency [wLyLlH97, RMC93, SW96, CKyL98, yLcY98, PMB99, WSM^{+}95]. Concurrent [ASSA96, Ara95, KC96, KD91, Kus90, LVB^{+}93, Ng93, SW93, She90, SBM94, Sta90, SP94, UH96, WH97, CLLC96, DY99, FW90, TS89]. concurrently [ST89]. Conditionals [HC86]. Configuration [Bro87, CZdV98, Rav81]. Configurations [ZX94]. conformance [NS92]. Congestion [Hać94, PV94]. Connections [Cho95, SSK98]. Cons [Gla90f]. Conservation [Leh80]. Considerations [RA91, Rog89, Won93, Car99, Gie79]. Consistency [CC99a, Kun91a, Liu95, VT99]. Consortium [DB86]. Constraint [Car96, UW95]. Constraint-Based [Car96]. Constraint/Rule [UW95]. Constraints [LH95, Pha94, UH96, WFZ96, ZR87, VT98, VT99]. construct [Jav88]. Constructing [Fer93, KH96]. Construction [OH94, vC80, MSHB98, RG79, SFM99]. Contemporary [Rey80, Gla95i]. Contents [AH81, Ano97m, Ano97n, Ano97o, Ano98f]. Context [AS96, HP90, HP92, SMS94, Tom89]. Context-Free [HP90, HP92]. contextual [NL99]. Contingency [Lan98a, NDM80]. Contingent [vS96]. Continuing [Bra89]. continuous [Tia99]. contract [NL99]. contract-based [NL99]. Contribution [KAL97, PV94, War89]. Control [ANB93, Bha84, CL94, CH83, CW90, FSA87, FZ93, Gla97c, Hać94, HB83, HUMT92, HU96, wLyLlH97, MO90, RUV92, San95, TM97, BG98, Çam99, CKyL98, Ken80, yLcY98, MA94, SW96]. Controllable [KMO91]. controlled [Vis99a]. controlling [Ebe99]. Controversy [Ano92e, Bab91, Blu89, Bri92, Ebe94, Gla91c, Har95a, Hei95, Pre90, Pul90, Qui94, RA91, Rei90b, SM92b, Tau92, Thi94, VPM93, Zuc90b, Zuc90a]. Conversion [Sny91, TE99]. converters [JS99]. COOL [Bra96]. Cooperation [HMG96]. Cooperative [NG91, SM92a, WM99]. Coordinating [Sch81]. Coordination [HMG96]. copy [HMC98]. CORBA [LFY^{+}99]. Corner [Ano92e, Bab91, BS93, Blu89, Bol97a, Bri92, Gla89f, Gla90b, Gla90a, Gla90g, Gla90c, Gla90d, Gla90e, Gla90f, Gla90h, Gla91a, Gla91h, Gla91c, Gla91e, Gla92b, Gla92e, Gla92f, Gla92h, Gla93g, Gla93e, Gla93d, Gla95d, Gla96d, Gla96e, Gla96h, Gla96j, Gla97f, Gla97c, Gla98f, Gla98j, Got92a, Har95a, Har95b, Pul90, RA91, Rei90b, SM92b, Tau92, VPM93, Zuc90b, Zuc90a, ZWM96, Ano86b, Ano87d, Ano90d, Ano91c, Ano91b, Ano92f, Ano92g, Ano92h, Ano92i, Ano93e, Ano93f, Ano94e, Ano94d, Ano94f, Ano95h, Ber94, BS96, Fle95, Gla86, Gla88a, Gla88b, Gla88c, Gla89a, Gla89b, Gla89c, Gla89h, Gla89d, Gla89g, Gla89e, Gla91d, Gla91b, Gla91i, Gla91g, Gla91f, Gla92d, Gla92c, Gla92g, Gla93a, Gla93b, Gla93f, Gla93c, Gla93h, Gla94c, Gla94d, Gla94g, Gla94e, Gla94f, Gla94h, Gla94b, Gla94i, Gla95c, Gla95a]. corner [Gla95f, Gla95e, Gla95b, Gla95g, Gla96b, Gla96c, Gla96f, Gla96i, Gla96g, Gla96k, Gla97d, Gla97e, Gla97i, Gla97h, Gla97k, Gla97g, Gla97j, Gla98a, Gla98g, Gla98h, Gla98k, Gla98c, Gla98d, Got93, Gul92, Hoa94, HY94, yL98, Len95, Pau92, Pla95, Pre90, Sai98, SW95a, Thi94, ZS95, Gla95j, Gla98i, Qui94]. Corporate [NB93]. Correct [Eva95, Ree85]. Correction [DT90]. Correctness [Bri92, MM93b]. Corrigendum [Gla99b, Gla00, Li99]. Cost [AH90, EHS93, Hag91, KT85, LP95, LM94, Leu92, MHSM99, OG80, WFZ96, vS83, BW80, VH89, Wey99, WM95]. cost-effective [Wey99]. Costs [AQ90]. Coteries [Nei97]. Could [OT92]. Counting [BK92, Gla99d]. couple [Ano94e, Gla94c]. Coupled [FG93, HJ90b]. Coupler [Cla86]. Coupling [Dha95, Loh84, OHK93, RY93]. Couriers [Bri92]. Course [BHR89, MC91, TE99]. Courses [CFSS98, MR99]. Coverage [FLN91]. Cray [CM86]. Creating [AC97]. creativity [Ano94e, Gla94c, Gla96c]. crisis [Gla96h, Gla97h, Gla98k]. Criteria [FN86, Ham81, OG80, SKF95, SKS96]. Criterion [Pas96]. Critical [DB86, GC94, JM96, LSD95, LvSL81, Ost92, Ozk97, SBB98]. criticism [Iso98]. Critique [Lit90, SI94, BM89, Gla95i]. Cryptographic [LL97b]. Cryptologic [Sny79]. CSMA [Lai97c]. CSMA/CD [Lai97c]. Current [ZS95, JH99, Ano95h]. Curriculum [Jef91, CR89, LS99]. Curve [BB81]. Customer [HHSR94, SBA97, FSA87]. Cycle [AJMP96, Bas80, FF95, HZ83, Leh80, RUV92, TD80, Dav88, Gla94d, Got93, OBS79, Tia99].

Damage [JM90]. Dangerous [Gla86]. Dass [DR84]. Data [AH81, Bel91, BBC^{+}88, Bla87, Cha91, CW97, DR84, Dam96, FZ93, GRS92, GSC91, JO83, KZ91, Ken84, KSW93, LHC96, Las90, Liu93, Mar84, MG81, MP90, Mot96, MK93, Myr90, OW84, PM90a, RT93, SG91, SW95b, SB88, TL96, TC93, Tan96, TK91, Uck91, UW95, Vel87, VP92, WSN92, WH91a, Won93, YRN80, YY93, YNDS88, vS83, BF96, De 98, HR95, HL94b, JF99, Jen99, Kuo94, LHC95, MP94, MA94, OL99, Ozk97, PL94, PM94, RC89, SW96, She89, Wei79, ZS88]. Data-Driven [YY93]. Data-Intensive [TL96]. Data/Knowledge [Mot96]. Database [Bar86, BW96, Bha84, BM83, Fri90, FM87, HB83, Mar84, NY84, NS87, PL83, SW94a, UH86, Ulu95, Ulu97, WM96, ÇZUB99, DMV98, DFCR96, GP98, HMP99, HY95, yLcY98, RB99, SBB98, TL99, Ulu98]. Database-Oriented [NY84]. Databases [KW93, wLyLlH97, SW95b, SKS96, BG98, CKyL98, JNY84, KR98, yL98, VT98, VT99]. Dataflow [BS86, KD91, Hsi91b, TL89]. DATALOG [Hsi91b]. date [Gla97h, Gla98k]. DBMS [Gor91]. Ddbms [DGS88]. Ddbx [FG93]. Ddbx-LPP [FG93]. Deadlock [Coo90]. death [Gil88]. Debate [Rei90b, Zuc90b, Zuc90a]. Debugger [Car83, ZENA93]. Debugging [BW83, BH83, FG93, Fri83, FAS94, GH83, HO96, PL83, STJ83, SKF95, OCCN89]. Decentralized [HJ91, Ken80]. Decision [DLS94, Mos84b, CTZ92, Mos84a, Pre90]. Decision-Based [Mos84b, Mos84a]. Declarative [FAS94, Kom88]. Decomposition [LL98, MPS86, Moy96, Mue86, Raj94, Kor99a]. Decomposition/generalization [Raj94]. Deductive [LL97a, Uck91]. Deeply [HC86]. Defect [KSH92, TT98]. Defects [Cai98, CW90, Eva97]. defense [Ano93f, Gla93b]. Defining [MV93]. Definition [HB89, JP94, Ros87, TK91]. Definition-Based [Ros87]. Definitions [Sku91]. degrees [Gla97e]. Deletion [Hab85]. Deliberations [CB91, Kun91b]. Delivery [Emd91]. Demand [HH87]. Denelcor [Hay86]. Densities [KSH92]. depend [VC97]. Dependability [CG94, Pow86]. Dependence [HOT97, HUMT92]. Dependencies [HB83]. Dependency [HR96, WH91b]. Dependent [KO95, Car99]. Deployment [ES97]. Derivatives [Sta90]. Derive [AQ90]. Derived [LV97]. Deriving [FM90a, Kuo94]. Describing [She89]. Description [MR84, Ayr98]. Descriptions [BYY87, Mar84, Mil96a]. descriptive [PL99]. Design [AH81, Amm91, BW96, BCD92, BYY87, Blu93, BM83, CL94, CH94, CDJ^{+}84, DG92, EHS93, Fic89, FK92, Gla90d, Gom89, Gom94, GRS92, GA95, Hać91, HG91, How80, Hur93, Joy87, KC96, Ken84, KW91, KW93, LZL97, LG97, Loh84, MM81, MJ89, Moh81, MB84, OC90, PW87, PdC94, Pha94, Row86, RT93, SGL93, Sak84, TOYI95, Tsu85, UW95, Var91, WNSC96, Whe81, WSR^{+}83, Won93, WFZ96, ZK85, ZCd96, Zho93, ZX94, CA88, CLLC96, CDDF99, Dav99, Dav95, DLT99, ES97, GJ88, Gla94g, HZ79, HLAB99, HR95, HHL^{+}97, Jef92, JS90, Kam89, KB98, KY92, KSFT89, MM93b, Mey88a, MR99, Nav92, Ost92, Phi98, PK89, Rey89, SFM99, TL99, UhCLS94, WCV^{+}98, ZLZ^{+}96]. Designing [BL95, Car92, NC88, San95, VPM93, HLC99, MM93b]. Designs [AC97, TZ81, WSN92, OSG98, RPL97]. Details [Hen88]. Detect [BAH96]. Detecting [Tri86b]. Detection [BFR96, Gla93i, Goe80, JM90, KL95, LHC96, Wha90, LJM96, MC98, WHMP99]. Determining [NDM80]. Develop [Amm91, PD98, TC93, SMCL96]. developer [Lin99]. Developers [Por93, LS98]. Developing [DK94, HH97, Kal92, MTON94, TM97]. Development [ANB93, BBØ96, CB89b, Coo81, Di 87, DS85, FWP93, Gas96, GK91b, GR97, HZ84, HL90, HHSR94, HS95, HH87, Jef87, Jos83, Joy94, KS96, KT85, Lan98a, LP95, Lee93, MM93a, MB84, NG91, Pan81, Phi81, Pla92, PL96, PZ94, PU84a, Ros87, Sah94, Sei89, SM92a, Sta83, SB93, Sub93, TC89a, TKS95, TDB97, TT93, WKM94, Zim84, vS96, vS83, vC80, Bar94, Dav88, FPW96, FAI97, Gla98d, Got93, HHB^{+}99, HBJ^{+}99, tLF89, MA89, NL99, PU84b, Sal80, TC89b, Tha80, Wei79, Zel88]. Diagnosis [RB93a, Hat99]. diagram [Kuo94, LJ99]. Diagrams [BTT84, DS85, JN84, TK91, WSN92]. Dictionaries [Cha91]. Dictionary [Mar84, Owo96, RF84]. Differences [OS87, EL88]. Different [SKS96, MBL^{+}99]. Differential [Rod86]. Difficulties [Jef96]. Difficulty [Sch97]. diffusion [BM89]. digital [Sny79]. Dimensional [MPS86, ZMAER99]. Dimensionality [SB93]. Dimensions [LO92]. Directed [BDM^{+}93, Kor83, PU84a, RG79, PU84b, vEHvV89]. directions [FN99, Sai98]. Directory [LS97]. dirty [Gla94d]. Disabbreviation [LTHR97]. disaster [Gil88]. Discipline [Gla94c, Ano94e, Gla94d]. disciplines [GAK92]. discontinuities [Jav88]. Discrete [Cla86, MS97]. discussion [SW88]. Disk [Hać91, TC93, CB89a]. Disk-Buffer-Cache [Hać91]. disks [CLLC96]. displacement [WJ99]. Displaying [MS97]. Distance [Blu89]. Distributed [Ara95, BFR96, Bar86, BW96, BW83, Bha84, BP91, BW95, BM83, Car96, DS94, FG93, Gas96, Gom89, Hać86a, Hać86b, Haç89a, Hać89b, HJ90a, HJS91, Hać93, HL94a, Hać94, HW94, HMG96, Hsi91a, HFK92, IMM95, KN97, LM94, LK93, Mor86, NG91, Nit96, PM90a, PD98, Pow86, Rah92, RW97, RT93, SAASA94, Sch97, Sho91, SF92, TW95, Tsu85, Ulu95, Ulu97, Ura90, WTS95, WM96, YP94, ZENA93, Zho93, Zho94, ZR94, BdADH94, BS96, CZdV98, CDS99, Car94, CL99, DLT99, DFCR96, HMC98, ISS98, JM96, Jia99, KHL^{+}99, MC98, PM99, RC89, Rot89, SPK99, SMU98, SBB98, TW98, TM98, Ulu98, WM99, ZZ88, ZLZ^{+}96]. Distribution [BB81, Dye93, SL80]. Distribution-Based [Dye93]. Distributive [Ver89]. disturbing [Gla94h]. diversity [Rom99]. division [WC99]. Do [Ano87f, FN85, Hen88, OT92, FF89, Gla93h, Gla98h, LS98]. Document [BCD92]. Documentation [Emd91, SG91, Sch81, CSKB^{+}89, Gla93a]. Documenting [BAEH96]. DoD [Rav81, SG91, Wal91]. DoD-STD-2167A [Wal91]. Does [VC97]. doing [Gla88c, Gla98d]. Domain [Gla92f, Jar93, KO95, Lam97, Pas96, Pou95, TM97, Ano92g, GW95, JF99, Jen99]. Domain-Dependent [KO95]. Domain-Independent [KO95]. Domain-Specific [Lam97, Pou95, GW95]. Domains [GV92, MO84]. dominance [CV95]. Domino [DB95]. Done [Gla91h]. doubtful [Gla96g]. Down [MM81]. DPDP [ZENA93]. dramatic [Gla96d]. Drat [LDN87]. DRDB [SBB98]. Driven [Har81, Jar93, Por93, YY93, DY99, Phi98, Rey89]. During [KSH92, FAI97, Lut96, SFM99, Zel88]. dying [Gla97b]. Dynamic [BFR96, FG93, Gan91, HJ90a, HJ91, KL90, Leu97, OSG98, SF92, YRN80, EOM95, GJ88, vdBK94]. Dynamically [FS91]. Dynamics [AH93, PL99, PMB99, WC99].

Early [KL95, Lok96, Dav88]. Eastman [Sed93]. easy [Dav95]. Economics [Blu93]. Edge [Zuc90b]. Edges [Ber93]. EDICT [BLPB92]. Editor [Fai85b, Gla97e, Gla98i, Ree85, SM83, vEHvV89, Ano83, Ano86c, Ano86b, Ano87d, Ano90d, Ano91c, Ano91b, Ano92f, Ano92g, Ano93e, Ano93f, Ano93g, Ano94e, Ano94d, Ano94g, Ber94, BS93, Bol97a, Fai83a, Fai83b, Fai83c, Fai84, Gla86, Gla88a, Gla88b, Gla88c, Gla89a, Gla89b, Gla89c, Gla89h, Gla89d, Gla89g, Gla89f, Gla89e, Gla90b, Gla90a, Gla90g, Gla90c, Gla90d, Gla90e, Gla90f, Gla90h, Gla91a, Gla91h, Gla91c, Gla91d, Gla91b, Gla91e, Gla91i, Gla91g, Gla91f, Gla92d, Gla92c, Gla92b, Gla92e, Gla92f, Gla92g, Gla92h, Gla93a, Gla93b, Gla93f, Gla93g, Gla93c, Gla93e, Gla93d, Gla93h, Gla94c, Gla94d, Gla94g, Gla94e, Gla94f, Gla94h, Gla94b, Gla94i, Gla95c, Gla95d, Gla95a, Gla95f, Gla95e, Gla95b, Gla96b, Gla96c, Gla96d, Gla96e, Gla96f, Gla96i, Gla96g, Gla96h, Gla96j, Gla96k, Gla97d, Gla97f, Gla97c, Gla97i]. Editor [Gla97h, Gla97k, Gla97g, Gla97j, Gla98f, Gla98g, Gla98h, Gla98j, Gla98k, Gla98c, Gla98d, Got92a, Got93, Har90b, Har93, Har94, Har95b, Hoa94, yL98, Pla95, Rad84, Rid81, Sai98]. Editorial [Ano98d, Ano98e, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Gla98c, Gla98d, Gla98e]. Editors [BDM^{+}93, Ano94f, Ano95h, BS96, CU98, HY94, MS79a, MS79b, OP92, SM80, SM81a, SM81b, SM81c, SM81d, SW95a, ZS95, ZWM96]. EDS [Won93]. EDT [Lai97a]. Educate [Gla91e]. Education [BLPB92, CFSS98, Mat86, Rus90, Bra89, CP88, CR89, HHB^{+}99, Sai99, WR99]. Educational [KCK^{+}98, JS90]. Educator [Joy94]. Educators [Gla91e]. Effect [FAI94, GR97, Loh84, SW88]. Effective [Fen93, Shu99, Tre81, SZ98, Wey99]. Effectiveness [ARAS94, Emd91, FZ93, GC94, SYB97, CW89, FF96, FWH97, HS99, WHMP99]. Effects [DG80, OCCN89, Sch97, Gla99c]. efficacy [HBJ^{+}99]. Efficiency [vS83]. Efficient [Bel93, BDM^{+}93, Fra86, Har81, KH97, Owo96, Ram90, SAASA94, SD94, TW95, KA96, WK88]. Effort [Dol97, DG80, Eva95, FWD97, JB91, Lee93, NQ98, SB93, SB95, WSD81, MT98]. Efforts [HH97]. Eiffel [Mey88b]. either [Gla95g]. Electronic [JT97]. elements [SFM99]. Elicitation [Lan98a]. Elliptical [MPS86]. Else [Lak93]. Embedded [War89, Whe81, De 98, Mar81]. Embedding [PdC94]. Emphasis [Lit90]. Emphasizing [CH94]. Empirical [AS96, BGB90, BBP96, Emd91, Pas96, Por93, Sta93b, Sub93, SB95, SYB97, BB89, HS99, HBJ^{+}99, Rob98, Sol87, WM95, WDMR99]. Employing [MF90]. Emulation [YY93]. Encapsulation [Joy87]. ENCOMPASS [TC89a]. encryption [HY95]. End [ZK85, Gla99d]. end-of-century [Gla99d].

Enforcement [HB83]. enforcers [Ano87f]. engine [LS92]. Engineer [Bab91, Pla92]. Engineering [AJMP96, ACCD91, BF81, BCD92, Boe83, BW93, BHR89, Bux90, CB89b, Chr91, CL95, CDJ^{+}84, DR92, EHS93, Fen93, FG94, GHC91, Gla92a, Gla96a, Gla97a, Got90, Ham81, HD84, Jac98, Jef91, Jef96, KSS84, KL96, KL91, Lan90, LL85, Mai96, MA89, MR80, Mey88b, Mil89, O’N83, Rey80, Sag95, Sed93, Sne83, Sta93a, TR89, VM89, Zel96, ZC97, Ano96m, BM89, Ber95, BS96, Bra89, CR89, CU98, DS98, Eri92, Fug99, Gla89c, Gla94a, Gla95c, Gla96b, Gla98b, Gla99a, Gla99b, Gla00, Har88a, HHB^{+}99, JTW98, Pfl99, PC98b, Qui94, Rad84, Sai99, Tom89, TFLW99, VB99, WCV^{+}98, WR99, LAHS97]. Engineers [MP89, TB95]. English [CW97, GI95, Gla93a]. Enhanced [CL97, WSM^{+}95]. Enhancing [LTHR97, HY95, LHC95]. Enough [Gla97f]. Ensuring [HHSR94]. enterprise [Chu97]. Entity [BTT84, CH94, JN84, MR84, Sak84, San95, CPW98, JNY84, Kuo94]. Entity-Life [San95]. Entity-Relationship [JN84, MR84, Sak84, JNY84, Kuo94]. Entity/Class [CH94]. Entropy [Moh81]. Enumeration [Ni97]. Environment [AM85, BFG97, Blu86, Chr91, DS85, Fri83, Har88b, HL90, HS95, IKCN91, JL97, KZ91, Kom88, Kus90, KCK^{+}98, Law81, Mey88b, MMSH92, Ng93, OW84, Par86, TC89a, TDB97, TT93, UH86, WNSC96, WM90, Zel96, CDM98, CC99a, DB95, HHZ92, TL89]. Environments [ACCD91, BL95, FG94, GHC91, Jef87, KSS84, KW91, MMSH92, PT91, Sch97, ZC97, FPW96, Ni97]. Equate [Zei88]. Equations [Rod86]. Er-Data [Mar84]. Ergodic [FN86]. Erlang [Lai97b]. Error [BDM^{+}93, Dye87, Gla93i, Goe80, JM90, MM93c, OW84, Sel93, TVK95, Wei79]. Errors [DG92, HP92, TDB97, Gla89g, Lut96, OCCN89, SW88]. escape [Gla95a]. escrow [Nec96]. ESPRIT [WBR90]. Essential [Jef96]. Establishing [ANB93]. Estelle [HHL^{+}97, HL98, JL97, Lai97b, Lai97a, LL99]. Estelle-based [HHL^{+}97]. Estimate [SB95]. Estimates [LP95]. Estimating [Cai98, HH97, Ozk97]. Estimation [AH90, BB81, BF81, Cav84, FS88, FWD97, Gla93e, JB91, KT85, MT98, MTON94, SB93, vS83, HTO97, MHSM99, NQ98]. Estimators [HP90]. Ethical [Car99, Kal92, McF92, Spa92]. Ethics [BLPB92, CM92, Got92a, Got92b, LIC92, Lue92, SM92b, Got90]. European [AM94]. Evaluate [ARAS94, BP86]. Evaluating [CW89, FF96, LV97, MM92, MG81, Pan81, PS90, Wei79]. Evaluation [Bha84, Bol97b, CFK91, CG94, CZ91, CR85, DV94, Esk89, FLN91, Haç89a, HO97, Ham81, HLAB99, Het95, Hsi91a, IYKO95, Loh84, MPS86, Mil96b, Moh81, Pow86, Rey80, Rv93, SYB97, TLPH95, Ulu97, WNSC96, WH97, Wey99, Ano96m, CTZ92, CSKB^{+}89, HTO97, HRS95, KJB97, Kor99b, KKiMT96, MD89, PKK98, Rid81, Ano84c, Goe84]. evaluative [SC99]. EVEN [JL97]. Event [Chr86, LVB^{+}93, Sch91, BG98, Phi98]. event-driven [Phi98]. Events [KD91, KM89]. Evidence [Bro81]. Evolution [ES85, Leh80, NS87, Wic92, RR98, Sto92, Woo80, Har97]. Evolutionary [PL92, Poo93]. Evolving [Bas97, Lea95, Har99]. Examination [Sub93, LvSL81]. Examining [Gla99c]. Example [PU84a, She94, Gla94h, HB89, PU84b]. Example-Directed [PU84a, PU84b]. Examples [Eli92]. Exceptional [TB95]. Exchange [Tre81, Gla95g]. Exclusion [DHP86, MS90, TW95, WTS95, JM96]. Exclusions [DS94]. Executable [GMM90, JM90, Kun95, MGJT87, TKU93, TC89b]. Execution [AM85, Dil91, JO83, LK93, Rec93, TT93].

Execution-based [Dil91]. existence [Gla96h]. Existing [LTT92, His98]. Experience [Amb87, Arc81, Blu86, Joy94, Lai97a, LZL97, Sca99, Sei89, OCCN89]. experienced [LS98]. Experiences [HBCC94, Hay86, Iso95, Lak93, MMSH92, Rei87, WRW93, SSK98, TE99, VM89]. Experiment [BC91, MD81]. Experimental [CSKB^{+}89, FLN91, KKiMT96, Loh84, Mil96b, Moy96, NY84, TLPH95, WNSC96, FWH97, LJ99]. Experimentation [Mac91]. experimented [Vis99b]. experiments [Fle95, KSFT89, Vis99a]. Expert [Col92, Eli92, Gla98g, Ker92, LO92, MMSH92, OT92, Pla92, Pop92, SM92a, SYB97, Wic92, KJ99]. Explanation [Wic92]. Exploiting [Vla98]. Exploration [Dam96, SFM99]. exponentiation [LC98]. Expressing [Lak97, RB99]. Expressions [Bra96, BH83, Hee90]. Extended [Bra96]. Extending [MM92, ST89]. Extensions [CH83]. External [Arc81, Ver89, WLC95]. Extraction [AB90].

Facilitate [KK81]. Facility [Sho91, DG98, Wei79]. Fact [Gla95h, Ken84]. Fact-Based [Ken84]. Factor [CR90, MTG92]. Factors [DLT99, DG80, FWP93, KMO91, LL85, SYB97, MB97, RS98]. Fail [Par98]. Failed [Ker92, Gla93f]. Failure [Gla98g, She94, SM92a, Gla96d, Gla98c, Hat99, Lin99]. Failures [ASSA96, FN99, Lip79]. faking [Gla94g]. Families [Gom95, SD94]. Family [Zvi93]. Fan [RT86]. Fan-Out [RT86]. Fast [BS86, Kor99a, ZR94, vD93]. Fault [Ban86, BW95, CL94, CG94, DG92, Fri90, FAI94, HOT97, KN97, KP93, LH83, MS90, Mor86, Mue86, OK94, PdC94, Ram90, SAASA94, STJ83, She95, WTS95, WWF94, WFZ96, ZG97, ZX94, FAI97, JM96, LFY^{+}99, LM96, SMCL96, Shu99, WHMP99, Hoa94]. Fault-Tolerance [Ban86, KP93, ZX94]. Fault-Tolerant [BW95, CG94, DG92, MS90, Mor86, OK94, PdC94, Ram90, WTS95, WFZ96, ZG97, JM96, SMCL96]. Faults [CMP85, Eva95, VPM93, TVK95]. FEAST [WL99]. FEAST/1 [WL99]. feature [GJ88, LJM96, TFLW99]. FedEx [WC99]. Feedback [AHGS92, Por93, Hat99, LR99]. fewer [Gla97e]. Field [Gla97m, Gla97g, Vis99b]. Fifo [MR86]. Fifth [Ano84c, Goe84]. File [CM93, FC96, Hać86a, Hać86b, Haç89a, Hać89b, HJS91, HJ91, MIH92, CB89a, SMU98]. File-Usage [CM93]. Files [HL94a, CLLC96]. filter [CLL99]. Finally [Gla92b]. Finder [AB90]. Finding [CH94, MS97, TS89]. findings [Gla98i]. Finite [Har81, SP94, TS89]. finite-state [TS89]. Firms [RZ94]. First [RA96, vC80]. fit [DS98, Gla96f]. Fits [Gla92b]. Five [PT91]. Fixed [PNK96]. Fixed-priority [PNK96]. flash [CC99b]. Flavors [Gla93g]. Flexible [NG91, PW92]. Flow [FZ93, HUMT92, JO83, Las90, Liu93, MM93c, PBC93, TK91, WSN92, Çam99, Jen99, Kuo94, DS85]. Flowcharts [Sca88]. flowgraphs [RG79]. Fly [DV94, Mil98]. Fly-by-Wire [DV94]. FMS [DY99]. Follow [Sed93]. Follow-up [Sed93]. Foreword [FM90b, Har90a]. Form [MBCD86]. Formal [Arm98, Art87, Coo90, Dye87, EC98, Fur93, GK91b, Gla91c, Gla93d, Gla95d, Gla96d, GV99, Jac98, JTW98, KL91, L’E87, Liu93, LSD95, MGH97, MG81, MP95, Nit98, Ost92, Par98, TK91, TZ92, VP92, CLSC98, Gla94e, JMM99, LF98, MSHB98, Sai98]. Formalism [Kun95]. Formally [BG96, Rec93]. format [SW99]. formats [CSKB^{+}89]. Formatting [Fis91, L’E87]. Forms [GK91b, SKS96]. formulae [vEHvV89]. formulation [CJP98]. FORTRAN [AC97, Rey80]. Found [KSH92]. Foundation [NS87]. Foundational [ANB93]. foundations [Mat96]. Four [Blu86]. Fourth [Joy94, RA96]. Fourth-Generation [Joy94]. FP [BK92]. FP-S [BK92]. FP2 [Bel93]. fragmentation [DFCR96]. Frame [HFK92, SGL93]. Frame-Based [HFK92, SGL93]. Framework [ANB93, BFR96, Bhi90, BC94, BF90, EL94, HR96, Lak97, MWH97, MV93, Mos84b, MP90, NG91, NC96, PM90b, Pre95, Rah92, SW93, Sam93, BM89, CJP98, DB95, GJP96, Mos84a, PA99, SC99, SC88]. fraud [Gla95h]. Fred [Ano87d]. Free [HP90, HP92]. French [FM90b]. FRSM [Liu95]. FTAM [LL99]. Full [Gla88b, RUV92, Got93]. Function [AR94, BK92, Dol97, ES97, FWD97, Rei90a, TC93]. Functional [BM93a, Dye93, HZ83, How80, KP97a, Mil96b, Moy96, Nel81, SAA93, De 98, LF91]. Functional-Decomposition [Moy96]. Functionality [Moy96]. Functionally [Amb87]. Functions [FS88, Hsi91a, KA96]. Fundamental [EL88, Gla95j]. Further [CA89, VVS99]. Fusion [SW95b]. Future [Ano87e].

G [PdF97]. G-Nets [PdF97]. Game [MTW97]. Gandalf [AM85, ES85, Not85a, Not85b]. Gap [CFSS98]. Garbage [Yua90]. Gateway [Bar86]. gateways [HMP99]. General [BFC92, Sei89, Yua90]. General-Purpose [Yua90]. generalization [Raj94]. Generalized [Bhi90, BH83, CT97, KP97b, vD93, WHL89]. Generating [BDM^{+}93, OL99, PS90, JMM99]. Generation [APL95, AM85, Bel91, BCFG86, FAI94, Joy94, RA96, CLSC98, FAI97, Gla97i, JF99]. Generator [AF96, MM93a, NY84, YCGH92, KP97a]. Generators [AF96]. Generic [MM93a]. gets [Gla98f]. Given [Leu92]. Glass [Gla88b]. global [LR99]. Glotos [Son93]. go [FF96]. goal [Gla96i, PL99]. goal-oriented [PL99]. Goals [Pfl95]. gold [Gla93f]. Good [Gla97f, BB89]. Gorbachev [Ano90d, Gla90c]. GOTO [BGB90]. GQM [MB97]. Graduate [TR89, Bra89, TE99]. Graduate-Level [TR89]. Grammar [Ara95]. Grammars [HP90]. Graph [Ara95, Chr86, Fra86, HOT97, PBC93, PM99]. Graph-Based [PBC93]. Graphical [Arm98, DK97, HG91, LG97, Sny91, MD89]. Graphics [MC91, ZS88]. Graphs [Del92, HUMT92, HL94b]. GRASPIN [Chr91, GHC91, GSC91, IKCN91, Krä91a]. greatness [Gla95a]. GreatSPN [Lai97c]. Gross [LJM96]. Group [ARAS94, HR95, Sch81, Jia99]. Growth [DLG96, Tau92]. Guest [Ano93g, Ano94f, Ano94g, Ano95h, Ber94, BS96, Bol97a, CU98, Got93, Har90b, Har93, Har94, Har95b, Hoa94, HY94, yL98, OP92, Pla95, Rad84, Rid81, Sai98, SW95a, ZS95, ZWM96]. Guidance [HHB^{+}99]. Guidebook [NB93]. Guidelines [CTA94, Joy87, MMSH92, Phi98]. Gulezian [BT97].

Hacker [Spa92]. Half [RB93a]. Handling [BT97, Gul96, UH86]. happened [Gla96k, Gla97k]. Hard [Ham81, Kor99b, ZLZ^{+}96]. Hard-To-Use [Ham81]. Hardware [GH83, Mos84b, WWF94, Mos84a, Nav92, Ozk97]. HaRTS [ZLZ^{+}96]. Hashing [TC93, TL95]. Having [KL95]. HDLC [Cla86]. Hdm [MD81]. head [GAK92]. head-of-the-line [GAK92]. header [Çam99]. header-population [Çam99]. Heap [BKS85]. heighten [MBL^{+}99]. Help [BB81, Ano87f, Gla95g]. HEP [Hay86]. here [FF96]. Hermod [OHBR90]. Heterogeneous [BL95, KZ91, PD98, RR98, TW98]. Heuristic [Bow84, ZR87]. Heuristics [Fer93, Gla91c, CZdV98]. HIBOL [WM90]. HIBOL-2 [WM90]. Hide [VPM93]. Hiding [Hen88, RC94]. Hierarchical [Bla87, Cha91, Hać93, LF96, Pow86, CZdV98]. Hierarchies [MM81]. Hierarchy [FWP93, Lee93, TL89].

High [AA98, Amm91, BW83, BH83, BM93b, GH83, KL95, KP97b, KP91, MMSH92, PU84b, PU84a, She90, GJ88, GKP98, LP93, Nav92, TCMJ98]. high-integrity [TCMJ98]. High-Level [BW83, BH83, GH83, KP97b, KP91, MMSH92, PU84a, She90, PU84b, GJ88, GKP98, LP93, Nav92]. High-Performance [BM93b, AA98]. higher [RVM99]. Highly [LS97]. Historical [AH90]. History [FJ98, GV92, Gla97m, Ayr98, Sal80]. Homogeneous [BBG86]. honored [Gla97g]. HOS [LF96]. Hospital [KZ91]. hould [Ano87e]. House [RB93b]. Human [FK92, Har98, Jef91, LL85]. Hybrid [Fra90, GK91b, Gor91, GW95, WFZ96]. hype [Gla96b]. Hypercube [Fri90, KP93].

I/O [MD91]. IBIS [KSW93]. IC [JT97]. Id [RF84]. idea [Gla95e]. Identification [Joy87, BM98, HLC99]. Identifying [CDDF99]. If [OT92]. ignorance [Ber95]. II [Gla94f, Dol97]. imagery [LJM96]. Imbedded [MR86]. immense [GP98]. Impact [CS85, Hur93, LR99]. Impacts [Sta93a]. Imperative [BBC^{+}88]. Implementation [AHG93, BW96, Bel93, BKS85, Eng81, Har81, Ker92, KP91, MMSH92, RT93, SL96, WLC95, Zho93, CPW98, GJ88, Lai95, LL99, TVK95]. Implementations [Car96, YFY96, JCJ99]. Implemented [BW93, ZCd96]. Implementing [Blu86, FSA87, LS97, MA94, Poo93]. Implications [FK92]. Importance [Gla92e, Gla92f, Ano92f, Ano92g, Ber95]. Imprecise [CZ91, PZ94]. Improve [DB86, FC96, HS99]. Improved [HL83, Bak88]. Improvement [Bol97b, CBK96, CP97, DLS94, HBCC94, TTP97, Chr99, Ebe99, Gla88c, HRS95, MB97, SC99, Tia99, VVS99]. Improving [Mil96b, Pul90, SMU98, Pfl99, Wey99]. Inaccurate [LP95]. Incentives [Pou95].

Include [MvS95]. incorporating [XHW99]. Increasing [YN91]. Incremental [Fis91, FW90, Fri83, Hee90, KK85, TC89a, vAW93, HHL^{+}97, TC89b]. Independent [KO95, PT91]. Index [DGS88, Ano80a, Ano80d, Ano81a, Ano81d, Ano84a, Ano84d, Ano85a, Ano85c, Ano86a, Ano86e, Ano87a, Ano87g, Ano88a, Ano88e, Ano89a, Ano89b, Ano89g, Ano89h, Ano90a, Ano90e, Ano91a, Ano91d, Ano92a, Ano92j, Ano93a, Ano93h, Ano94a, Ano94h, Ano95a, Ano95i, Ano96a, Ano96n, Ano97a, Ano97l]. indexing [BF96, DF98]. Indices [Rv91, Rv92]. individual [WH99]. Industrial [AF96, Arm98, BFG97, BHR89, CB89b, CLO95, OW84, WR99]. Industrialization [Stu83]. industrially [Lai99]. Industry [DB86, GK91a, MBL^{+}99, FF89, Sny79, Tha80]. Industry/university [MBL^{+}99]. Inefficiency [BAH96]. Inexpensive [MPS86]. Inference [CL94, Sta85, LS92, VH89]. Influence [Sny79]. Influencing [SYB97]. influential [MB97]. INFORM [vEHvV89]. Informal [BYY87, LF98]. Information [ARAS94, Blu86, BY85, CFSS98, DR92, DLG96, DF99, FSA87, Gla92a, Hab85, Hen95, Hen88, HUMT92, KAL97, KJB97, MR83, PL96, RF84, SGL93, Tau92, TKS95, Tre81, WSN92, WNSC96, ZC97, Zho94, vS96, Bar94, Gla98h, HLAB99, HBJ^{+}99, Kam89, Ken80, LS99, LJ99, MD89, NDM80, RC94, SSvdW99, XHW99]. information-hiding [RC94]. information-systems [Kam89]. INGRES [HMC98]. Inheritance [AHGS92, AHG93, RMC93, LH98]. Initiated [HJ90a]. Initiative [DB86]. innovations [BM89]. Innovative [ACCD91]. input [SMU98]. input/output [SMU98]. ins [Spa92]. Insecurity [AP97]. Insight [Gla91h]. inspection [NL99]. Inspections [KSH92]. Instantiation [MM93a]. Institutions [Gla96a, Gla94a, Gla95c, Gla97a, Gla97e, Gla97j, Gla98b, Gla99a, Gla99b, Gla00]. Instrumentation [BAL81, KHL^{+}99]. Instruments [MP89]. integrate [ST89]. Integrated [Car94, CH94, FM93, Fri83, Mai96, TL99, Tia96, ZR87, Bhi90, CDM98, LL99]. Integrating [HL90, KZ91, LTT92, Mar84, Sed93, SW95b, CC94]. Integration [Arm98, EL94, HZ84, MR80, O’N83, PL99, RBM95, Sta99, VB99, Jen99, ZS88]. integrity [MA94, TCMJ98]. intelligence [PP94]. Intelligent [Dam96, KP97a, MWH97, Nit98, RF84, WM99, CJP98]. intended [Rom98]. Intensive [TL96, Sta99]. Interaction [KP97a, Nit98, Har98, Mur99]. Interactive [Amb87, BAL81, DK94, DK97, Fis91, MC91, Mer87, YNDS88, ZENA93, Bra89, DL99, ZS88, vEHvV89]. interchange [SMS94, SW99]. Interconnected [BFC92]. Interconnecting [ZEB88]. Interconnection [Arc81, PH93, PDN86, ZSGS93, Kor99a]. Interconnectivity [KH81, Sel93, RB89]. Interdisciplinary [Har98]. Interface [CB91, HHSR94, Hur93, JS90, Kun91a, LG97, WLC95, Bak88, Kun91b, MM93b, PL94]. Interfaces [GK91b]. Internal [Liu95]. International [Rus90, tLF89]. Interoperability [Tre81]. Interpreter [BS86]. Interrelationships [TD80]. Interrupts [Kri93]. Interweaving [PL96]. Introducing [Krä91a]. Introduction [Ano84c, Bas80, Bec86, Goe84, Har88a, Har90b, IYKO95, JNY84, Mar81, OP92, Ano83, Ano93g, Ano94g, CU98, Fai83a, Fai83b, Fai83c, Fai84, Fai85b, Har93, Har94, KY92, MS79a, MS79b, Rad84, Rid81, SM80, SM81a, SM81b, SM81c, SM81d, SM83, Wil89]. Intrusion [LHC96]. investigate [MB97]. Investigation [Emd91, Joy87, Loh84, RBM95, WH91a, HS99, LJ99, WDMR99]. investment [RS98]. Investments [RS98]. Invocation [Kor83].

IP [SSK98]. ISCC [LS99]. Ismos [NY84]. ISO [Lai95, Lai97a, LL97a, LL99]. Isolation [Dil91]. Issue [FM90b, Har90a, Ano84c, Bas80, Bec86, CU98, Goe84, Har88a, JNY84, Pla95]. Issues [FG94, Hać86b, CL99, De 97, Fic89, Gla96h]. Italian [RZ94]. Items [SG91]. Iteration [BBP96, Sta93b, PMB99]. ITOC [HLAB99].

JACK [BFG97]. Jackson [CC94]. JAD [Dav99]. Japan [AM94, Duv95]. Job [AP97, BBG86, VC97]. Join [AT97]. Joint [Dav99, Sch81]. Jointly [GAWC91]. Jointly-owned [GAWC91]. Journal [BT97, FM90b, Gla99b, Gla00, Li99, Gla89e]. JSS [Gla95j, Gla96k, Gla97k].

key [Nec96]. keys [WJ99]. kits [GW95]. Knowledge [AJMP96, Fra90, KB96, MW95, MP90, Mot96, Pla92, Pop92, Pre95, Sam93, STJ83, She90, Uck91, CU98, Eri92, HHZ92, Pla95]. Knowledge-Based [Fra90, KB96, MW95, MP90, Pre95, Sam93, STJ83, She90, HHZ92, Pla95]. Known [Hen88]. Kodak [Sed93]. Kung [CB91].

L [Gla88b]. Laboratories [HBCC94]. Laboratory [BF81, MA89, VM89]. Language [Arc81, BS86, Bel93, BYY87, BBC^{+}88, BK85, CL81, CT94, CH83, Coh81, GMM90, HG91, Joy94, KP91, Krä91a, MGH97, Maz81, Mey88b, MO90, PT91, PD98, Ros87, Sku91, TKU93, UW95, Whe81, ZSGS93, BK95, De 98, GJ88, Jav88, JMM99, Nav92, SW88, TCMJ98, UhCLS94]. language-supported [BK95]. Languages [AM81, CH83, CGD^{+}96, FM90b, Gan91, HP92, Kor83, MF90, NC96, PDN86, RMC93, SAA93, TK87, YFY96, PK89, Rom99, Sto92, War89]. LANs [BFC92]. Large [Ara95, Bla87, Di 87, Gom94, HL90, Leh80, MWH97, MWH98, Nit96, OKOM97, Rey80, Sel93, WWC97, WRW93, WWC98, Wey99, WCV^{+}98, WM99]. Large-Program [Leh80]. Large-Scale [Gom94, HL90, OKOM97, WWC97, WWC98]. Last [Gla90a, Gla98h]. later [Gla97k]. Laws [Leh80]. Layered [Amb87, HO96, YCGH92]. layering [WSM^{+}95]. Layout [BTT84, HHZ92, Son93]. lead [Gla93h]. leadership [HS99]. leads [Ano94d, Gla94b]. learn [Gla98d]. Learned [LV97, Rei87, BW80, Gla97j, Tom89]. Learner [Sca88]. Learning [OT92, SW88]. Legacy [AC97, BM98, CDDF99, TT98]. legend [Gla89h]. Less [HC86, OT92, VPM93, Gla93h]. Less-than-Successful [OT92]. lesson [Gla97j]. Lessons [Ker92, LV97, Rei87, Tom89, BW80, Qui94]. Letter [TL95, WJ99]. Letter-Oriented [TL95, WJ99]. Level [BW83, BH83, Fai85a, GH83, KP97b, KP91, MMSH92, MP90, PU84a, Rv92, She90, TR89, ZSGS93, GJ88, GKP98, LP93, Nav92, PU84b, Rv91, Rog89, TE99]. levels [RVM99]. librarians [Gla95g]. Life [AJMP96, Bas80, FF95, HZ83, Leh80, RUV92, San95, TD80, Dav88, Gla94d, Got93, OBS79, Tia99]. life-cycle [Tia99]. lifecycle [PMB99]. Lightning [Gla89f]. like [CT94]. Likely [VPM93]. Limited [RT86]. Limits [Tau92]. Linda [Has98]. line [GAK92]. Linear [RT86, YRN80]. Lines [Dol97]. link [Gla92g]. Linkage [ZS88]. linked [Kar94]. linking [FPW96]. LISP [Ng93, YY93]. List [SD94, YRN80]. Listings [LDN87]. Lists [DT90, GAW92]. Literature [LL85, Not85a, Ano91c, Gla91i]. Little [Gla90g]. Littlewood [Lit80]. Living [BR90]. Load [HJ90a, HJ91, HL94a, RCSD93, Sho91, DLT99, Hać92]. Load-Building [HL94a, Hać92]. Local [DT90, ZZ88]. Locality [TL89]. Localization [STJ83, Sel93]. Locally [CW97]. Location [CL94]. lock [CKyL98]. lock-based [CKyL98]. Logic [BCFG86, Fer93, GMM90, Jma96, Kom88, Kus90, UH96, De 97, She89]. Logical [PdF97, TT93]. Logistic [Sch81]. Long [BR90, Gla98h]. Long-Living [BR90]. look [Ano87d, Ano88d, Gla86, Gla89b, Gla91g]. lookup [WK88]. loop [Ozk97]. Loops [FN85]. Loosely [FG93, HJ90b]. Loosely-Coupled [HJ90b]. Losing [Zuc90b]. Lost [Gla92h, Hat99]. Lotus [Lit90]. Low [Fai85a, VH89]. low-cost [VH89]. Low-Level [Fai85a]. LPP [FG93]. LSS [Rob98].

M [MR86]. Machine

[Har81, LL97a, LTT92, TT93, YY93, VH89]. Machines [Yua90, ZCd96]. Macro [HC87, Lit90]. made [FF96]. Madness [Gla90d]. Mae [Har88b]. magnetic [CB89a]. Mainframe [Ano86d]. Maintainability [CLO95, LTHR97, LH93, OH94]. Maintaining [LDN87]. Maintenance [Gla90e, Gla91e, HS95, HR96, Jar93, KB96, KS96, KL95, LPLS87, PL92, RUV92, Sch97, SKV94, vAW93, DS98, Gla92g, Gla93a, Gla93h, PC98a, Vis99a, WT89]. majors [Gla93a]. makers [Pre90]. Making [DLS94, KHL^{+}99, Sku91]. Manage [Rv92]. manageable [KHL^{+}99]. Management [ARAS94, AM94, Ber81, Bla87, BR90, Bro87, Duv95, EHS93, Gla96h, HR96, JP94, KSW93, KK81, KW93, Lan90, LM94, LO92, MK90, PM90a, Pul90, RA91, RT93, SGL93, Sag95, Sch81, SW94a, SKV94, SB88, Tau80, AH88, Ano91b, CB89a, CC99a, DMV98, Gla91d, Gla91f, Ken80, KR98, Mur99, PL94, PM94, Rav81, RCL99, ZZ88, Bas80]. managerial [BM89]. Managing [GA95, Hać93, JL85, Rv91, ZM96]. Mandated [SG91]. Manipulation [MR83, SD94]. Manpower [BB81]. Manufacturing [TM97]. Many [BBG86, BM96, Gla93g]. mapping [CL99]. Mark [Dol97]. Marketing [FF87].

Massively [KM92]. Master [Mil89, TE99]. match [CL98, CLLC96]. Matching [Kor83, Cha93, HBJ^{+}99, PM99]. Material [TR89]. mathematical [Woo80]. Mathematics [Gla96e]. Matrix [CBK96, KL96, WWC98]. maturation [Gla89b]. Maturity [PM90b, Pfl95, Gla97j]. McCabe [Tak97]. Means [Esk89]. Measure [BE81, BAL81, HC87, MM92, Rey84, Tak97]. Measured [Gla90a]. Measurement [AR94, BBF^{+}90, BF81, BK95, BF90, Cav84, Cha95, DR92, Het95, Mac91, Mor99, MK93, Mus80, Por93, RUV92, Rus90, Sed93, TTP97, Tia99, vS83, PL99, TT98, VM89]. Measurement-Driven [Por93]. Measurements [Hon90]. Measures [BBC^{+}88, BK85, FM90a, LV97, MTG92, Moo98]. Measuring [AR90, CA88, Cio91, FRF98, Gon95, HL98, MA89, She94, TZ81]. Meaures [HB89]. Mechanism [BH83, CL98, FPW96]. mechanisms [JCJ99, KR98, SMCL96]. Medicine [Col92]. meeting [Gla96i, Bev99]. meetings [Jef92]. Member [Mot96]. memo [KA96]. memo-functions [KA96]. Memories [BW95, GP98]. Memory [Hać91, IMM95, SF92, CC99b, ISS98, LF91, SMU98, TL89]. MENDELS [UH96]. Mental [FAI94, LPLS87, SFM99]. Menu [Art87]. Menu-Based [Art87]. Merge [Yan94]. Mergesort [Ver89]. Message [Hać94, MF90]. Message-Passing [MF90]. Metadata [FC96, SGL93]. Metasystem [BST93]. Method [BAEH96, BYY87, BK92, CH94, CL97, Gla90d, Gom89, HL83, Hur93, KH96, LL98, LHC96, Liu93, MPS86, Pan81, Vel87, BM98, CC94, FJ98, Iso98, LC98, NDM80, WJ99, WK88, WDMR99]. Methodologies [Gla96f, TOYI95, HBJ^{+}99, KY92]. Methodology [BY85, Bro87, CS85, EHS93, Gas96, HBCC94, IYKO95, KZ91, KL96, Law81, LH90, Lee93, MB84, She94, ZCd96, Zvi93, Kuo94, Raj94, Rid81, RG79]. Methods [EC98, Esk89, Fen93, FF87, Fur93, Gla91c, Gla93d, Gla95d, Gom94, Hag91, Jac98, LH95, LSD95, Par98, WRW93, Zim84, DIP98, Gla93h, Gla94e, Gla96d, GV99, HLC99, JH99, JTW98, LF98, Ost92, Sai98, Wil89]. METKIT [WBR90]. Metric [Eva83, Gon95, Har88b, HS95, KAL97, MK90, NC96, OHK93, PS90, RY93, SKV94, vS83, CJP98, Li98, Li99, RC94, RB89]. Metric-based [PS90]. Metrics [AM94, Bhi90, BST93, BC91, BN90, BC94, CRV94, CL95, CBOR88, CR90, CMP85, DS92, Deh90, EL94, Gla89f, Har90a, HL90, HG91, KH81, KML94, LH93, Moh81, MV93, MM93c, Myr90, NB93, OKOM97, OC90, PM90b, Pfl95, RAC90, Rey84, SN91, Sed93, SI94, SS98, VM93, Wha90, WSD81, YNDS88, CDDF99, FN99, MD89, Rey89, RB89, TVK94, Har90b]. metrics-driven [Rey89]. Micro [HC87]. Micro/Macro [HC87]. Microarchitectures [OK94]. microcode [TVK94]. Microcomputers [MIH92]. Migration [MM95, DG98, ISS98, TW98]. Mikhail [Ano90d, Gla90c]. Mil [Coo81]. Mil-Std-1679 [Coo81]. millennium [Sai99]. MIMD [ZEB88]. mine [Gla93f]. minimization [WHMP99]. Minimum [Chr86]. Minimum-Time-Reachability [Chr86]. Minnowbrook [Ano84c, Goe84]. MIS [Jef87]. Misinterpretation [Cio91]. misleading [Gla86]. Mission [DB86]. Mission-Critical [DB86]. MKH [CLLC96]. mobile [CC99b, WM99]. Mod [DT90]. Model [AHGS92, AHG93, Ara95, Bel93, BW93, BY85, BFC92, EL94, FZHS95, FSA87, FAI94, Goe80, Hać89b, HZ83, HB83, HH97, HO96, HFK92, Jar93, Jef87, JB91, KP97a, KD91, KP97b, LM94, MGH97, MSI90, Phi81, PH86, PL92, Poo93, Pop92, PL83, SL96, Tak97, TZ92, UW95, Var91, VT87, WNHM86, ZK85, ZC97, CPW98, GAWC91, Hać88, HTO97, HMC98, JS99, KC98, Kuo94, LP93, LAHS97, Lit80, LM96, SZ98, SW99, SFM99, Vla98, WHL89, WD99, Woo80]. Model-Based [EL94]. Model-Driven [Jar93]. model-oriented [LM96]. Modeling [BT97, CB91, Eva95, FS88, FF95, FM87, Gul96, HOT97, Kun91a, Kun95, Mai96, Mer87, Mil96a, Mot96, Por93, RCSD93, Sak84, San95, SBM94, Sta83, SP94, WC99, WFZ96, Zel96, BW80, Car94, DY99, DL99, HR95, HCC91, KMR99, Kun91b, LP93, Nav92, Sca99]. Modelling [Cla86, CP97, Hać86a, KNT86, GV99, PL99, WL99]. Models [CLO95, Dha95, FWD97, FAS94, HS95, KMMG91, KL91, KW91, LV97, LL97b, LPLS87, MBCD86, MS97, MD91, MTON94, Pfl95, Sta85, Tau92, Tör90, ZEB88, MHSM99, MO84, SC99, TVK95]. Modern [Gla93h]. Modes [Sub93, Fug99]. Modest [Mat86]. Modifiability [Loh84]. Modified [PH86]. modular [EL88, FW90, HCC91]. Module [Arc81, Bow84, Loh84, OHK93, PDN86, ZSGS93, Leu97, TM98]. Modules [BMSB94, KL95, PBC93, SH98]. Monitor [TT93, Zho93]. Monitoring [HO96, YRN80, CZdV98]. Moral [Col92, Lue92]. Mortem [Hag91]. MOSIS [Ayr98]. MP [CM86]. MPEG [Ng99]. MSc [BHR89]. MSE [Mil89]. Multi [WDMR99, LJM96, PA99, WCV^{+}98]. multi-agent [PA99]. Multi-method [WDMR99]. multi-paradigm [LJM96]. multi-process [WCV^{+}98]. multicast [JCJ99]. Multicasting [Hać94]. Multiclass [MR86]. Multicomputer [Amb87, Çam99]. multicomputers [KH97]. multidisk [GAK92]. Multiflawed [Gla88b]. multijoin [vdBK94]. multimedia [CL99]. Multimicroprocessor [GDF86]. Multimode [KNT86]. Multiparadigm [HL93, Ng93]. Multiparty [Cho95]. Multipath [ZX94]. Multiple [CHB94, Del92, MBCD86, TKU93, CLLC96, OSG98, Rog89, ST89]. Multiple-Bus [MBCD86]. multiple-level [Rog89]. Multiple-Viewed [Del92]. Multiprocessor [MBCD86, PH86, SF92, ZCd96, CKyL98, Hać88, Kar98]. Multiprocessors [Rod86, SG89, WWC97, CTA94, LF91, WWC98]. Multiproject [AH93]. Multisequential [TK87]. multistage [PV94]. Multitasking [CM86]. Multithreaded [RW97]. Multiuser [GAW92, MIH92]. Multivariate [LHC95]. Multiviewpoint [MW95]. mundane [Ano88d]. Mutation [BM93b, FWH97, WM95]. Mutual [DS94, DHP86, MS90, TW95, WTS95, JM96, MBL^{+}99]. Mutual-Exclusion [DHP86]. My [Gla89d, Gla94e]. Mystery [Gla90g]. Myth [Gla90b]. mythology [Gla88a]. Myths [FF89, LF98]. MYTHSEEKER [Rog94, Hei95].

narrative [Ayr98]. natal [Gil88]. Natural [BYY87, Bra96, Maz81, JMM99]. Naturalness [Gla96g]. navigation [OSG98]. necessarily [LF98]. Needed [BM96]. Needs [AM81, Bev99, CR89]. Nested [FN85, HW94, HC86]. nesting [HB89]. Net [GMP94, KP93, KP97b, LM94, LL97b, PdF97, Var91, JS99, LP93, SC88]. Nets [Chr86, Coo90, FN86, KH96, Sch91, SBM94, vD93, GKP98, HCC91, Kor99b, PdF97]. Network [BP86, Bla87, Hać93, KL95, LS97, LG97, MR86, SW93, Ulu97]. Networks [BBG86, DHP86, FWD97, Hać86a, MWH97, MWH98, Nit98, PH93, Tsu85, ZK85, Çam99, Kor99a, PV94, Rog89, VVS99]. Neural [FWD97, KL95]. next [Gla89b, Sai99]. No [Ano87d]. Node [IMM95]. Non [Sch91, Gla96i, KWME99]. non-goal-oriented [Gla96i]. non-repudiation [KWME99]. Non-Sequential [Sch91]. Nonblocking [WM96]. Nonchange [Gla96k, Gla95j, Gla97k]. Nondeterminism [DS92]. Nondominated [Nei97]. Nonexceptional [TB95]. Nonprogrammer [OS87]. Nonuniform [PH93]. Normal [SKS96]. Normalization [Mil98]. Notification [HR96]. Notion [Tai93]. Notions [Mot96]. Novices [YN91]. NPath [MM92]. NUMA [WWC97, WWC98]. Number [Cai98]. numbers [Gla95f].

O [KL96, MD91]. O/A [KL96]. OASys [Vla98]. Object [AC97, AHG93, Bar94, BW93, BC94, CH94, Dav95, EHS93, ES97, Gla96g, JB91, JH99, KO95, KS96, KSW93, KN97, Kun95, KGH^{+}96, LH93, MWH98, MSI90, Mil96a, MO90, MD89, Moy96, NC96, Ng93, PM90a, PBC93, PD98, RA96, RMC93, SW93, SCG^{+}93, Sei89, SW94a, Sta93a, Tai93, TL96, UW95, WRW93, ZZ88, BK95, BF96, Car94, ÇZUB99, CPW98, CLSC98, CC94, DIP98, FPW96, GV99, GJP96, Har97, HL94b, Jia99, KC98, LS92, LP93, Li98, Li99, LFY^{+}99, Mat96, MT98, NQ98, PL94, PM94, Raj94, RR98, Rom99, SPK99, SW96, SS98, SMCL96, UhCLS94, WK88, WDMR99, Chu97, Gla93c, Gla94f, Got93]. object-based [BK95]. object-linking [FPW96]. Object-Orientation [Moy96, Gla96g, Gla93c, Gla94f]. Object-Oriented [AC97, AHG93, BC94, CH94, EHS93, JB91, KO95, KSW93, Kun95, KGH^{+}96, LH93, MSI90, Mil96a, NC96, Ng93, PM90a, PBC93, PD98, RA96, RMC93, SW93, SCG^{+}93, Sei89, SW94a, Sta93a, TL96, UW95, WRW93, Bar94, Dav95, ES97, JH99, MO90, MD89, BK95, BF96, Car94, ÇZUB99, CPW98, CLSC98, CC94, DIP98, GJP96, Har97, HL94b, LS92, LP93, Li98, Li99, Mat96, MT98, NQ98, PL94, Raj94, Rom99, SW96, SS98, SMCL96, WK88, WDMR99, Chu97, Got93]. Objective [ÇZUB99]. Objectives [ANB93]. Objects [MS97, PL96, WM90, CDDF99, GAWC91, ZMAER99]. Obsolete [Hab85, Gla92c]. obvious [Gla95e]. OCCAM [BdADH94]. odd [Ano94e, Gla94c]. Offs [GA95]. old [Gla95j]. OMT [HK98]. once [CB89a]. One [CL97, FN86, Rei87]. One-Place [FN86]. only [Gla98k]. OO [JMM99]. Open [FG94]. Operating [SCK86, TT93, ST89, GAWC91]. Operating-system [GAWC91]. Operation [CH94]. Operational [ANB93, FAS94, RBM95]. Operations [Hać91, NM93, RA91, HL94b]. OPNets [LP93]. opportunities [MBL^{+}99, TE99]. optical [CB89a]. Optimal [CL97, PM99, Pha94, UH86, CZdV98, WHL89]. Optimistic [BKS85, wLyLlH97, CKyL98]. optimization [vdBK94]. Optimum [Leu92, OG80]. Order [BP86, KML94]. Ordered [KD91]. Orderings [LVB^{+}93]. orders [CTA94]. Oregon [Har90b]. Organization [BY85, Car99, MP94]. Organizational [AP97, Lan98b, Law81, Mat96, Tha80, Thi94]. Organizations [Owo96, Sny79]. organized [RB89]. Organizing [BBØ96, Tan96]. Orientation [Moy96, Gla94f, Gla96g, Gla93c]. Oriented [AC97, AHG93, BC94, CFK91, CH94, CG94, EHS93, JO83, JB91, KO95, KSW93, Kun95, KGH^{+}96, LH93, MWH98, MSI90, Mil96a, NY84, NC96, Ng93, PM90a, PBC93, PD98, RA96, RMC93, SW93, SCG^{+}93, Sei89, SW94a, Sta93a, TOYI95, TL96, TL95, TDB97, UW95, WRW93, Bar94, BK95, BF96, Car94, ÇZUB99, CPW98, CLSC98, CC94, Chu97, Dav95, DIP98, ES97, Gla96i, GV99, Got93, GJP96, Har97, HL94b, JH99, KC98, LS92, LP93, Li98, Li99, LM96, Mat96, MT98, MO90, MD89, NQ98, PL94, PL99, Raj94, RR98, Rv91, Rom99, SW96, SS98, SMCL96, TM98, WJ99, WK88, WDMR99]. ORL [UhCLS94]. Other [MS97]. Our [Gla92h]. outlook [Rav81]. output [SMU98]. overheads [WWC98]. overloads [Rot89]. Overview [AF96, Ber91, CBOR88, IKCN91, Kam95, PK89]. owned [GAWC91]. Owns [Har95a].

Package [DK94, SL96]. Packages [LTT92]. Packaging [Bas97]. Pager [WLC95]. papers [Ano92d, Ano93b, Ano93c, Ano93d, Ano96m, Mar81, Gla86]. Paradigm [GHC91, Sah94, LJM96, MB97, SPK99]. Paradigms [Moy96]. Parallel [AT97, BP86, Bel93, BAH96, FG93, Fra86, Hać86a, HL94a, Hay86, IMM95, KM92, LZL97, MPS86, MIH92, RT86, Sho91, SP94, Tan96, Won93, WNHM86, ZENA93, De 97, Has98, HCC91, LF91, SMCL96, SMU98]. Parallel-Processing [Hay86]. Parallelism [Ban86, FN85, Tri86b, Vla98]. Parallels [HD84]. Paramedic [CM92]. Parameterized [NM93]. Parametric [JB91, Krä91b]. Paramita [JC98]. Parnas [Gla86]. Parr [BB81]. Parser [KK85]. Parsing [Hee90, KK85, MGH97, vAW93]. Part [MvS95, Mot96]. Partial [EC98, LVB^{+}93, Rey84, Rod86, CLLC96, Rey89]. Partially [KD91]. Partition [CLL99]. Partitioning [BE81, Gie79, KC96]. partly [Gla91g]. Parts [BDM^{+}93]. Party [Gla90b]. partying [Gla99d]. Pascal [L’E87, Lok96]. PASS [MIH92]. Passing [MF90]. password [JC98]. past [SW88]. Patchwork [BS86]. Path [BH83, CL97, UH86, VVS99, WHL89]. path-selection [WHL89]. Paths [BM96]. Pattern [Kor83, Cha93, FPW96, War89]. pattern-based [FPW96]. Pattern-Directed [Kor83]. Patterns [CM93, LIC92, MS97, PH93, TL99]. Payoff [Bro81, Gla91a]. PC [HHZ92]. PDL [OC90]. Pencil [Gla96j]. People [OS87]. perception [KJ99]. perceptions [Lin99]. Performance [Amm91, Ano84c, Bha84, BAL81, BM93b, CZ91, FC96, Goe84, Gor91, GDF86, Hać86b, Haç89a, Hać89b, HJ90b, Hać92, IMM95, KP97b, KNT86, Lai97c, LZL97, MPS86, Par86, PH93, RA91, Rv93, RCSD93, SAA93, SKS96, Ver89, WNHM86, AA98, CLL99, GAK92, KR98, Kor99b, Row86, SMU98, SW99, ZLZ^{+}96]. Performance-based [ZLZ^{+}96]. Performance-Reliability [Hać86b]. Periodic [ML95]. Person [KCK^{+}98]. personal [ST89, VH89]. personnel [FF89]. Perspective [AH93, Deh90, Hon90, Joy94, O’N83, Pla92, RAC90, RA91, Hal92, JS90, tLF89]. perspectives [MBL^{+}99]. Petri [Chr86, Coo90, FN86, GKP98, HCC91, JS99, KH96, KP93, KP97b, LP93, LM94, LL97b, PdF97, SC88, SBM94, Var91, vD93]. Ph.D. [Gla97e]. Phase [KL95, OKOM97, TD80, HY95]. phased [WD99]. Phases [Zim84]. Philosophy [BBF^{+}90]. Phrase [AB90]. pictures [CL98]. Pilot [Gla97l, MM95]. Pioneer [Gla97m]. piracy [Thi94]. pitfalls [Gla95f]. PL [Bel91]. PL/1 [Bel91]. Place [FN86]. Plagiarism [Wha90]. Plan [Pfl95, Sam93, MD89]. Planning [DLG96, Sam93, WC99]. PLEASE [TC89b]. Plenty [Gla95f]. Point [BK92, Rei90a]. Pointer [BL98]. Points [AR94, Dol97, FWD97, FTAM99, Cha95]. Points-to [FTAM99]. Poisson [Eva97]. Policies [CH83, HH87, RCSD93, WKM94, CC99b, DIP98]. Policy [Sch81, Gla95j, Gla97k, Qui94]. Polymorphism [Krä91b]. Polynomials [OH94]. Populating [Pou95]. population [Çam99]. Portability [HO97]. Portrait [Bab91]. Post [Hag91]. Post-Mortem [Hag91]. Postgraduate [BHR89]. Potholes [Gla88b]. Practical [DA86, VP92, WH97, TCMJ98, ZMAER99]. Practice [AM94, Duv95, Gla90h, Hag91, Het95, LSD95, RBM95, Ano94d, Dav99, Gla89i, Gla93c, Gla94h, Gla94b, Gla95b, Gla98j, yL98, Qui94, Gla94f]. Practices [GK91a, PW87, FF89, JH99, PC98a, PC98b]. Practicing [MP89]. Practitioner [MRW^{+}94]. Pragmatic [Bar92, Jef92, NS83, GP98]. pre [Gil88]. pre-natal [Gil88]. Predicate [Sch91]. Predicate-Event [Sch91]. Predict [LH93, LS98]. Predicting [Hur93, OH94]. Prediction [Cav84, CW90, Lee93, Lip79, Lok96, She95, NQ98, TVK95, XHW99]. Predictive [LV97]. Preference [Sca88]. prefixes [WH99]. Prepaging [HH87, WKM94]. Prepare [Ano87e, Sam93]. presence [PV94]. presentations [Jef92]. primary [HMC98]. prime [Gla96h]. Principle [ZX94]. Principles [Boe83, Loh84, PW87, FJ98]. Prior [SL80]. priorities [Hać88, Liu98]. prioritization [RST98]. Prioritizing [FWP93]. Priority [RCSD93, GAK92, Hać92, PNK96]. Probabilistic [FZHS95]. Probability [HP90]. Problem [Chr86, Gla90e, Nit98, Ano91c, Gla89d, Gla91i, Gla97h, HR95, PA99, PV94]. Problems [BB81, MPS86, WB89]. Procedure [SAASA94, ZM96, KKiMT96]. Procedures [KK81, OS87]. Process [APL95, ANB93, Bhi90, BBØ96, BW93, CT94, CB91, CP97, De 97, DLS94, FWP93, FG94, Gla88b, Gla93e, HBCC94, HHSR94, Kun91a, Lai97a, Lan90, Lee93, LAHS97, PM90b, Phi81, RY93, SL96, TM97, CC99a, Chr99, Ebe99, HRS95, KMR99, KSFT89, Kun91b, LR99, MB97, PL99, RVM99, RCL99, SC99, Sca99, Vis99a, Vis99b, WL99, WCV^{+}98]. Process-based [De 97]. Process-Centered [FG94]. Process-integrated [Bhi90]. Processes [AR94, AS96, BCD92, KD91, KL91, Let87, TK87, GAW92, Hać88, Mor99]. Processing [Amb87, Hay86, Lai97a, PD98, Rah92, RW97, Sho91, Tsu85, Ulu95, ZENA93, vS83, De 98, LJ99, Ulu98]. Processor [Par86, RT93, MJ89]. Processors [FG93, ML95, Kar94]. Produce [SG91]. Producibility [Car92]. Producing [VL94]. Product [Esk89, Lan90, MBCD86].

Product-Form [MBCD86]. Production [BCD92, DK97, HBCC94, HP90, Ker92, Rv93, Gla97e]. Production-Based [Rv93]. Productivity [Blu89, Cha95, DB86, FWP93, Gla88b, Gla90c, Gon95, GR97, JL85, Jef87, KMO91, Law81, Tau92, Ano90d, Ano91b, Gla88c, Gla91f]. Professional [Got92a, Mat86, TKS95]. Professionals [CM92, Lue92, RZ94, FF89]. Program [AS96, BYY87, BL98, CS85, CH83, Eva83, FS91, GA95, HOT97, HL83, HB89, HUMT92, HU96, JO83, KL95, LDN87, Leh80, Let87, MS81, Mar84, Mil89, MD91, NB93, PW92, RBCM91, Sed93, SKV94, TZ92, WBR90, Yan94, Zho93, Bra89, FTAM99, Kam95, RB89, TL89, WHL89, Qui94]. Programmer [KMO91, OS87]. Programmer-Nonprogrammer [OS87]. Programmers [AP97, Gla97c]. Programming [AHG93, BF81, Bla87, BCFG86, BN90, CS85, CH83, Coh81, DG80, FM90b, Gan91, Gla90b, HL93, JL85, JB91, Kom88, Kor83, Kus90, Law81, Lit90, MO90, Nel81, OC91, PT91, SCG^{+}93, She90, TK87, WM90, WSD81, BB89, CC94, CP88, De 97, EL88, LF91, Li98, Li99, Mat96, OCCN89, Phi98, Raj94, SMCL96, Sol87, SW88]. Programs [AR90, BAH96, Ber93, BBC^{+}88, BK85, BP91, Car96, Dil91, Fer93, Har95a, KM92, KML94, KL90, KGH^{+}96, L’E87, LTHR97, LZL97, LK93, Lok96, MGJT87, Rey84, SBM94, TL96, UH96, VPM93, WNHM86, BdADH94, BB89, CLSC98, EOM95, HHB^{+}99, HCC91, Moo98, Rey89, Rot89, VB99]. progress [WT89]. Project [AH90, AH93, Ber81, Iso95, IKCN91, KT85, KK81, LM94, MM95, MK90, Pul90, RB93a, Tau80, WRW93, AH88, Gil88, Gla89h, Lin99, Mur99, Rog94, XHW99, Hei95, Not85b]. Projection [Sta83]. Projects [Bla87, Eli92, MRW^{+}94, OT92, SM92a, DL99, SS98]. PROLOG [BP91, LS92, Moo98, Ura90, Vla98].

Promise [SCG^{+}93]. Promising [KCK^{+}98]. Promoting [DIP98]. PROMPT [Lai95]. Prone [Coo90]. Proofreader [Gla96j]. Proper [RB93a]. Properties [BFR96, Nei97, PdC94]. Proposal [AR90, Cha95, Mat86, vC80]. Proprietary [Hen88]. Pros [Gla90f]. ProSet [Has98]. protection [GAWC91]. Protocol [BMSB94, JT97, Lai97c, LL97a, wLyLlH97, WM96, ZK85, HHL^{+}97, JS99, KR98, Lai95, Lai97d, LL99, NS92]. Protocols [ASSA96, CG94, Jma96, LL98, LL97b, SW93, Ulu97, CKyL98, Jia99, Rog89]. Prototype [FS91]. Prototypes [Bel91, BM93a, IKCN91, CTZ92]. Prototyping [BG96, LVB^{+}93, Sch91, SSK98, VC97, Zho94, CDS99, HK98, Has98, LCZ98]. Provide [Por93]. Providing [Hen88, TE99]. Pseudocode [Sca88, Rey89]. public [Nec96]. public-key-based [Nec96]. publish [Gla89e]. pull [DF99]. Purpose [Yua90, Ozk97]. push [DF99]. push/pull [DF99]. PVM [PD98].

QA [Fis81]. QoS [CL99]. Quality [Bev99, Bol97a, Bol97b, DR92, DB86, Esk89, Gla92b, Gla92e, Gon95, HG91, Hon90, KH81, LV97, Pre95, RB93b, Rv91, Rv92, Tak97, TZ81, Ano87f, Ano92f, Ber94, Gla92g, Liu98, Pla95, RST98, SBA97, Wey99, ES97, Gla91d]. Quantifier [Bra96]. quantify [KB98]. Quantifying [KAL97, KSFT89]. Quantitative [DLS94, Dha95, DL99, HRS95, Lan90, TLPH95, Liu98]. queries [CJP98, VL94, vdBK94]. Query [RT93, vdBK94]. Quest [SW94b]. question [LH98]. Queue [MR86]. Queueing [BBG86, FMP86, Hać86a, MR86, Mue86, RCSD93]. Queues [Cla86, Hać92]. Quick [KK81]. Quo [MWH97]. Quorum [NM93]. Quorum-Based [NM93].

R [Rei87, Rei90a]. Race [Bri92]. radical [Gla95e]. Raised [Mat86]. Random [MTW97]. Rank [TC93]. Rapid [CDS99, Zho94]. Rational [Gla93e]. re [Gla97j]. re-learned [Gla97j]. Reachability [Chr86, NS92]. reachable [TS89]. reaching [Gla97j]. reactions [DF99]. Reactive [Fur93, JVP^{+}98, Sah94]. Readable [HC86]. Reader [Ano92h, Ano92i, Fle95, Gla95g, Gul92, Len95, Pau92]. readers [Gla94e]. Reading [Bas97]. Real [BG98, CL94, CRV94, GMM90, Gom89, Gom94, GRS92, HW94, Hal92, HFK92, KY92, wLyLlH97, yL98, LH95, ML95, NC96, OK94, PZ94, Rei90a, Ulu95, Ulu97, WM96, Yua90, ZCd96, CKyL98, DMV98, DY99, Hoa94, KR98, Kor99b, Lai97d, yLcY98, LP93, Ost92, Ozk97, PNK96, RG79, SSvdW99, SBB98, Sto92, Ulu98, VT98, VT99, ZLZ^{+}96]. Real-Time [CL94, CRV94, GMM90, Gom89, Gom94, GRS92, HW94, HFK92, wLyLlH97, LH95, ML95, NC96, OK94, PZ94, Rei90a, Ulu95, Ulu97, WM96, Yua90, ZCd96, BG98, Hal92, KY92, yL98, CKyL98, DMV98, DY99, Hoa94, KR98, Kor99b, yLcY98, LP93, Ost92, Ozk97, PNK96, RG79, SBB98, Sto92, Ulu98, VT98, VT99, ZLZ^{+}96]. real-world [SSvdW99]. Reality [SCG^{+}93]. realization [Rog94]. Really [Gla97c, Har95a, Rei90b, Zuc90a, Ano87f, Ano87h, FF89]. Reasoning [FWD97]. REBOOT [SCK95]. Reborn [CHB94]. recapture [Iso98]. recognition [AA98]. Recognize [Hen88]. Recombining [Ber98]. Reconciliation [Lan90]. Recovering [DG87]. Recovery [ASSA96, BDM^{+}93, SAASA94, Won93, YP94, HLAB99, WCV^{+}98]. Recursion [BBP96]. Recursive [JO83]. Redesign [BBØ96]. Reducing [WM95]. Reduction [Bra96, Hag91, LHC96, SB93]. redundancy [EL88]. Redundant [CLLC96]. Reengineered [SW95b]. Reengineering [APL95, AS96, Jar93, MM95, Sag95, SW95a, WSM^{+}95, WLPL95, Ano96m, BM98, CDM98]. References [Gla90f]. Refinement [Raj85, Var91]. refinements [BdADH94]. Reflections [Gla97m]. Reformulating [Gul91]. Regression [BT97, FWD97, Gul96, KGH^{+}96, MTON94]. Reifer [Rei90b, Zuc90b, Zuc90a]. related [JNY84, Lut96]. Relational [Bra96, JN84, Pop92, SKS96, Uck91, HMP99]. Relations [MSI90]. Relationship [BTT84, CH94, JN84, JP94, MR84, Sak84, CPW98, Eri92, Gla89i, HZ79, JNY84, Kuo94]. Relationships [Dol97, HB83, RB99]. Relative [HS95, MK90]. Relatively [Sca88]. Release [Leu92, OG80, XH98]. relevance [WR99]. relevant [Lai99]. Reliability [Bha84, Cav84, DV94, FS88, Hać86b, HCC91, KK81, KNT86, Mus80, OG80, Rot89, SL80, SW94b, ZEB88, CW89, EL88, Iso98, Kor99a, Lit80, MO84, RCL99, Tia99, TM98, XHW99]. reliability-assurance [CW89]. reliability-oriented [TM98]. Reliable [Di 87, Jos83]. Remaining [Cai98]. remarks [CA89]. Remote [ZM96]. remotely [LJM96]. Removal [Dye87]. Rendering [SF92]. Rendezvous [DS92, WNHM86]. renewal [Vis99b]. Repeated [AB90]. Repetitive [Hat99]. replicated [KM89, Vis99a]. Replication [HJS91]. report [Gla91b, Got90, Sch81]. Reports [AH81]. Repositories [Pou95]. Representation [BBC^{+}88, MR83, Pop92, Uck91]. representations [KC98]. repudiation [KWME99]. Required [HH97]. Requirements [AM81, AB90, ANB93, CL95, Del92, DF84, GMP94, Gom95, HHSR94, JP94, Lam97, Lan98a, Liu93, MvS95, Sam93, Wal91, Ber95, Dav95, FF89, JTW98, Liu98, Lut96, NDM80, Rav81]. Research [KSW93, MRW^{+}94, RA91, SB88, Wei79, Ano87d, Fug99, Gla86, Gla91g, Gla95i, RST98, Sai98, WDMR99, MD89]. Researchers [Hen88, Gla95g]. researches [Lai99]. resequencing [Kar98]. reserved [Ng99]. Resilience [MvS95]. resolution [Zwe90]. Resolving [CA87b, CA87a, KMM89]. Resource [BB81, Cho95, Coo90, FMP86, Sch81, SG89, Zel88, ZR87, Leu97]. resource-allocation [Leu97]. Resource-Deadlock [Coo90]. Resources [Hać86a, Hać93]. Responding [DG92]. Response [BP86, BT97, KMM89, Kun91b, Zuc90a, Mur99]. Response-Time [BP86]. Responses [LIC92]. Responsibility [Col92, HHSR94]. Responsible [FK92]. restoration [VVS99]. Restricted [BS86]. restructure [KB98]. Restructuring [HL83, TL89]. Results [AH90, AM94, CBOR88, Gla90a, Lai97c]. Rethinking [Fug99]. Retrieval [Owo96, CL98, CLLC96, UhCLS94]. Retrieve [GI95]. Retrospect [Wic92]. Reusable [DJL93, Gom95, WH91b, BM98, RS98]. Reuse [DJL93, FF95, Hen95, Iso95, Lam97, MRW^{+}94, PP94, SCK95, TL96, TDB97, WRW93, WLPL95, ZSGS93, Ano95h, BK95, CDM98, Gla98e, His98, LH98, RS98, WD99, ZS95]. Reuse-Oriented [TDB97]. Reusing [BMSB94]. Reverse [BCD92, WCV^{+}98, Ano96m]. Review [KCK^{+}98, LL85]. Reviews [Gla93i, PW87, KKiMT96]. Revisited [Ebe94, Gla90h, Raj85, Hei95]. Revolution [Fis81, Gla90a, Gla99c, Har97]. RHODOS [DG98, ZG97]. Rigor [Mac91]. Rigorous [HB89]. ring [KH97]. Risk [KL95, She94]. rivals [Lit80]. Road [Gla88b]. Robert [Gla88b]. Robust [MTON94]. Rods [Gla89f]. Role [AP97, FM90b, RZ94, Ano91b, Fle95, Gla91d, Gla91f, HS99]. Rollback [YP94]. Rollback-Recovery [YP94]. Ronald [BT97]. roots [Har98]. Routing [Hać94, MWH97, Kar98]. Rule [MP95, UW95, Fic89, Moo98]. Rule-Based [MP95, Fic89, Moo98]. Rules [Eva83, L’E87, PL96, Hsi91b]. Run [BFR96, LF91, Bak88]. Run-time [BFR96, LF91, Bak88]. runaways [Gla98i]. s [Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, BK92]. Safety [CFK91, FM93, GC94, LSD95, MvS95, LM96, Lut96, Ost92]. Safety-Critical [GC94, LSD95]. Safety-Oriented [CFK91]. safety-related [Lut96]. salvaging [CV95]. same [Gla95j]. Sampling [Dye93]. satellite [Rog89]. satisfaction [Gla94i]. Satisfy [UH96]. Satisfying [VT98, VT99]. Scale [AR94, Gom94, HH97, HL90, OKOM97, WWC97, WWC98]. scenario [CLSC98]. Schedule [AH90]. Scheduling [CZ91, DK97, Ker92, MC91, WWC97, ZR87, CTA94, PNK96]. Schedulings [BAH96]. Schema [Sak84]. schemas [RB99]. Scheme [CT97, TC93, Won93, HY95, JC98, SH98]. Schemes [TL95, Gla99d, Rom98]. Scholars [Gla96a, Gla94a, Gla95c, Gla97a, Gla97j, Gla98b, Gla99a, Gla99b, Gla00]. Science [CA87b, FM90b, Gla92a, KMMG91, LIC92, TLPH95, CA87a, CA89, CA90, Fle95, Gla89c, Gla97e, KMM89, LvSL81, VB99, Zwe90]. Scientific [KSW93, Rei90a]. scientist [Gla96e]. Scientists [LIC92]. Scoring [RPL97]. scripts [Chu97]. seamlessness [Gla96g]. search [CLL99, Gla95e, PM99, SBA97]. searches [Ano91c, Gla91i]. Searching [Tan96]. Seattle [Mil89]. Secondary [Kus90, WK88]. Secret [CT97, HLC99]. section [Sol87]. Secure [JT97, SZ98]. Security [BM83, MvS95, HY95, LHC95, VB99]. Seeding [HOT97]. seem [Gla96g]. SEGRAS [Krä91a]. SEL [RUV92]. Selecting [MS97, Ozk97]. Selection [CL97, DA86, Fra90, LH90, MMSH92, Pas96, Vel87, Zvi93, NDM80, WHL89]. self [CTA94]. self-scheduling [CTA94]. Semantic [MR84, MV93, MM93c, VM93, OCCN89, She89]. Semantics [HMG96, MP95]. Semaphore [NM93]. Seminar [FM90b]. Sender [HJ90a]. Sender-Initiated [HJ90a]. sensed [LJM96]. Sensitivity [Eva83, XH98]. SEPS [LAHS97]. Sequences [MTW97]. Sequencing [HL83]. Sequential [AQ90, Sch91, VH89]. Server [Won93, WNHM86, Gla97d, SMS94]. servers [ZG97]. Service [LS97, Nit98, Rv91, Rv92, LFY^{+}99]. Service-Level [Rv92, Rv91]. Services [Gas96, Rv91, Rv92, SBB98]. Set [CL97, FM93, ML95, SKV94, VL94, WHMP99]. set- [VL94]. Sets [BCFG86, LVB^{+}93]. Setting [Ano86d]. Seven [Boe83, Sta93a]. Several [WSD81]. SGML [MGH97]. Shared [BW95, Hać86a, GAW92, ISS98, LF91, SMU98]. shared-memory [LF91]. Sharing [CT97, FMP86, Sho91, HLC99]. Short [Sca88]. Should [Gla89e]. showing [RB89]. signaled [SSK98]. silver [Ano87d]. SIMD [AT97]. similar [XHW99]. Similarity [Owo96, yLcY98]. SIMPARC [BAH96]. Simple [KK81, ZR87, vD93, Kor99a]. Simplified [BK92, MR83]. Simulated [Ree85]. Simulating [MWH98]. Simulation [AH90, BP86, Chr99, Kar94, LG97, Mer87, Rey80, SW93, WSN92, WNSC96, AH88, Chu97, DB95, DL99, HMC98, KMR99, RVM99, RCL99, Sca99, LAHS97]. Simulation-based [AH90, AH88]. Simulators [BAH96]. Singapore [PC98a, PC98b]. Single [AH93, VL94]. Single-Project [AH93]. single-valued [VL94]. Size [Bow84, Lok96, HTO97, WHMP99]. Sizing [BC91, Rei90a, VT87, Ber88]. SL-trees [BF96]. Slice [Hsi91b, HU96]. Slicing [BL98, KL90, Kam95]. Small [DLG96, Eva97, HH97, RZ94]. Small-Scale [HH97]. Smdc [YNDS88]. Smerfs [FS88]. SMILI [KM92]. smoothing [Ng99]. Snooping [BW95]. SOFL [OL99]. Sofspec [NS83]. SoftClass [MRW^{+}94]. Softcost [Rei87]. Softcost-R [Rei87]. Softening [Sne83]. Softest [MS81]. Software [AM81, APL95, AJMP96, ACCD91, Amm91, Ano84c, Ano86d, Ara95, ANB93, AS96, AM94, Bab91, BBF^{+}90, BF81, BL95, Ber81, Ber91, Bhi90, BBC^{+}88, BS96, BW80, Boe83, Bol97a, BST93, BC91, BN90, BW93, BT97, BC94, Bro81, BHR89, BF90, Bux90, Cai98, CA87b, CB89b, Car92, Cav84, CL81, Cha95, CT94, Chr91, CV95, Cio91, CLO95, CR89, CW90, CRV94, CGD^{+}96, Coo81, CBOR88, CMP85, CDJ^{+}84, DS92, DGM93, DG92, DLS94, Deh90, Del92, DJL93, Dha95, Di 87, DB86, DS85, Duv95, Dye87, Dye93, Emd91, Esk89, EL94, Eva97, Eva83, Fai85a, FS88, FM93, FM90a, Fen93, FN99, FG93, FM90b, FWP93, FWD97, Fis81, FF95, FF87, FG94, GHC91, GI95, Gla88b, Gla89f, Gla90c, Gla90d, Gla90e, Gla91e, Gla92a, Gla92b, Gla92e, GV92, Gla93e]. Software [Gla95h, Gla96a, Gla97a, Gla97f, Gla97m, Goe80, Goe84, Gom89, Gom94, Gom95, Gon95, GR97, GC94, Gul96, HL94a, Hag91, HO97, HBCC94, Ham81, Har95a, HC87, Har90a, Har90b, Hen95, HL90, HG91, Het95, HD84, Hon90, HS95, Hur93, Iso95, JVP^{+}98, Jar93, Jef87, Jef91, Jef96, JL97, Jos83, Joy87, Joy94, KH81, KC96, KB96, KSS84, KMMG91, KMR99, KSH92, KS96, KAL97, KN97, KT85, KK81, KL91, KCK^{+}98, Lak97, Lan90, LV97, LH98, LL85, Lea95, LP95, Lee93, LM94, Leu92, LH83, Li99, LTT92, Lin99, LPLS87, Loh84, tLF89, LF96, Mac91, MM95, MTG92, MM92, MR80, Mey88b, MRW^{+}94, Mil89, MTON94, Moh81, MP89, MB84, MP90, MK90, Mus80, Myr90, NS87, NG91, OKOM97, OHK93, OG80, OH94, OW84, Pan81, PBC93, PdC94]. Software [PdF97, PM90b, Pha94, Phi98, Phi81, PL92, Poo93, PC98b, Por93, Pou95, PU84a, Pul90, RZ94, RVM99, RST98, RAC90, Rey80, RB93b, RCL99, Rus90, Sah94, Sai99, SN91, SL80, Sch97, Sch81, Sed93, Sei89, She94, She95, SL96, SCK95, Sne83, Sta93a, SKV94, Sta83, Sta85, Sta90, Sta93b, SP94, Stu83, SB93, Sub93, SB95, SB88, Tak97, Tar92, Tau80, Tau92, TC89a, TTP97, TR89, Tör90, TVK94, Tri86b, TKU93, TB95, VT87, VM93, Wal91, WH97, WL99, Wha90, WH91b, WSR^{+}83, WLPL95, WWF94, WFZ96, XHW99, YN91, YNDS88, ZS95, Zel96, ZC97, Zuc90b, vS83, vAW93, vC80, AH88, Ano87d, Ano87f, Ano88d, Ano90d, Ano92f, Ano94d, Ayr98, BM89, Ber94, Ber98, BK95, Bra89, CA87a, CA88, CA89, CA90, CTZ92, Car99]. software [CU98, CSKB^{+}89, Dav88, DF98, DS98, DL99, Ebe99, ES97, Eri92, FPW96, Fug99, FAI97, Gil88, Gla86, Gla89c, Gla89h, Gla89g, Gla91d, Gla92d, Gla92g, Gla93h, Gla94a, Gla94d, Gla94g, Gla94b, Gla95c, Gla95b, Gla96b, Gla96h, Gla97g, Gla98b, Gla99a, Gla99b, Gla00, Got93, HTO97, Har88a, Har99, HRS95, HHB^{+}99, His98, Iso98, JNY84, Kam89, KB98, KMM89, LHC95, LvSL81, LR99, Leu97, LS98, Lip79, Lit80, Liu98, LJM96, LCZ98, Lut96, MB97, MA89, Mey88a, Mor99, Mur99, MO84, NC88, OSG98, Ozk97, Pfl99, PK89, Pla95, PC98a, PU84b, Rad84, RPL97, RS98, SC99, Sal80, Sca99, SS98, SMCL96, SH98, Sta99, SFM99, TC89b, Tha80, TT98, Tia99, Tom89, VM89, VB99, Wei79, Wey99, WCV^{+}98, Wil89, WR99, Woo80, XH98, Zel88, ZLZ^{+}96, Zwe90, Ano91b]. Software [Ano95h, Bas80, DB86, Gla88c, Gla91f, Gla98i, Got90, LAHS97, MA89, Qui94, VM89]. Software- [MP90]. software-dependent [Car99]. Software-Engineering [LAHS97]. software-module [Leu97]. Solution [BBG86, Chr86, Gla90e, MBCD86, RT86, Thi94]. Solutions [FN86]. solved [Ano91c, Gla91i, Gla98f]. Solving [Rod86, Gla89d, PA99]. Some [AM94, Bro81, Gla89g, Gla90a, Gla90f, Gla91h, Gla97i, Gla97h, HL94b, Sah94, ZX94, Lit80, PK89, BW80, Gla94e, Gla98i, LF98]. Someone [Gla92b, Lak93]. Sorry [Het95]. Sort [Krä91b]. Sorting [Ver89, Amm89]. Source [CR90, LTHR97]. Space [KA96, PM99, PA99, WHMP99]. Space-efficient [KA96]. Spare [VVS99].

Sparse [CBK96]. spatial [CL98]. Spc [DB86]. Special [CL81, CU98, Har90a, Sol87, Ano84c, Bas80, Bec86, Goe84, Har88a, JNY84, Pla95, FM90b]. Specific [DK94, Lam97, Pou95, TM97, GW95]. Specification [Ara95, Art87, BFR96, BMSB94, BBC^{+}88, BS93, BST93, CL81, CGD^{+}96, DA86, DR92, Fur93, HL98, JVP^{+}98, JL97, Jma96, KD91, Krä91a, Kri93, Lai97a, LL97b, Liu93, LF96, Mil96b, MvS95, NC96, NS83, TKU93, VP92, Wal91, WSR^{+}83, Ano93e, CLSC98, CL99, Jav88, LL99, Ost92, Rob98, RG79]. Specification-based [JVP^{+}98]. Specification-in-Large [Ara95]. Specifications [AM81, Arm98, Bel91, BM93a, BCFG86, Coo90, DGM93, EC98, GMM90, GMP94, JvB83, Krä91b, LF98, Liu95, LCZ98, MG81, PU84a, Ura90, Ber98, FRF98, JMM99, MSHB98, OSG98, OL99, PU84b, TC89b]. Specified [BG96]. Specifying [CH83, Del92, OS87, Rec93, Sny91]. Speedup [BP86]. splitting [VVS99]. Spreadsheet [DK94, Lit90, NB93, ZR94]. SPRINT [FJ98]. Stack [Amm89, BKS85, CHB94]. Stack-based [Amm89]. Stack-Heap [BKS85]. STAD [Las90]. Stages [DLG96, Dav88]. stand [ST89]. stand-alone [ST89]. Standard [BBC^{+}88, Bow84, Sch81, Rom98]. Standardization [Coh81]. Standards [Ano86d, Ano87f, Eng81, Fis81, Tre81, Mar81]. Start [RB93a]. State [BL98, Duv95, FN86, FG94, Har81, Het95, RBM95, PM99, Sto92, TS89]. state-of-the-art [Sto92]. Stateful [HMP99]. Statement [BGB90]. States [Chr86, TS89, Duv95]. Static [BL98, CMP85, SC88, TVK95]. Stationary [Mue86]. Statistical [Bro81, Dye93, FS88, KMO91]. Statistics [AH90]. Status [Mat86, Rav81, Ano95h, ZS95]. STD [Wal91, Coo81]. Steady [FN86]. Steady-State [FN86]. Stennis [SGL93]. Step [RA96, vC80]. Stepwise [Raj85]. Stochastic [FN86, KMMG91, KP97b, LM94, WSN92, vD93, HCC91]. Storage [Kus90, Maz81, ZK85, CB89a, MP94, WK88]. story [Gil88, Gla94d, Gla96d, Gla98c, Lai97d]. Strange [Gla96j]. STRAPS [Fai85a]. strategic [WC99]. Strategies [Eli92, FZ93, PMB99, Tar92, WR99]. Strategy [CW97, UH86, Zei88, HMC98, NDM80]. Stream [JO83]. Stream-Oriented [JO83]. street [Gla95j]. Stress [FAI94, FAI97]. String [Maz81, Cha93]. Strings [MS97]. Strongly [Gan91]. Structural [AR90, MP90, PL92, Poo93, WHL89]. Structurally [FM90a]. Structure [Arc81, BCD92, BY85, CG94, Gla95i, GR97, HU96, MK93, Tau80, BF96, HCC91]. Structure-based [Gla95i]. Structure-Oriented [CG94]. Structured [BC91, Fra90, Gla90a, IYKO95, Lee93, MGJT87, Sca88, TOYI95, TZ81, CC94, Gla91g]. Structures [JN84, YRN80, Tha80]. Structuring [Eva83]. students [Gla97e, SSvdW99]. Studies [PW92, Gla97l, Sol87]. Study [AH90, AR94, BGB90, BBP96, BMP97, DGM93, DJL93, Dol97, Duv95, EC98, FZ93, GK91a, Gla96h, Gor91, HO97, JVP^{+}98, KMO91, MRW^{+}94, PT91, Rv92, SN91, SAA93, Sed93, SW94b, Sta93b, SB88, TOYI95, TL95, TLPH95, Ulu95, Wic92, WSD81, AH88, BP80, BB89, DFCR96, GKP98, Gla89b, Gla97j, HLAB99, HBJ^{+}99, Kar94, KR98, Lin99, MT98, RR98, RS98, Rob98, Rom98, SSvdW99, WM95, WHMP99, XH98]. Style [OC91, BB89]. Subdomain [MPS86]. Subgraph [BL98]. Subgroup [Sch81]. Subject [Ano80d, Ano81d, Ano84d, Ano85c, Ano86e, Ano87g, Ano88e, Ano89g, Ano89h, Ano90e, Ano91d, Ano92j, Ano93h, Ano94h, Ano95i, Ano96n, Ano97l, Pha94]. Subjective [SL80]. Subsets [BT97, Gul96]. Subsystem [Lak97]. Subway [DGM93]. Success [SM92a, Gla96d, Gla98g, Gla98c, Lai97d, RS98]. success/failure [Gla98c]. successes [FN99]. Successful [OT92]. successive [BdADH94]. Sufficient [Hen88]. suitable [DF98]. Suite [YFY96, Li98, Li99]. Summary [Sca88, VM89]. supercomputing [GJP96]. Support [ARAS94, DR84, KB96, MP90, NS87, SW95b, TTP97, Chr99, DB95, Gla96c, GAWC91, LF91, LM96, MSHB98, Rey89, Rom99, SFM99]. supported [Bar94, BK95]. Supporting [DS98, TT93]. supports [Gla96h]. Surface [SF92]. surprising [Gla98i]. Survey [AM81, AM94, Rus90, CL99, De 97, Eri92, KY92, RST98]. Switched [PH93]. Symbol [Maz81]. Symbolic [CR85, Dil91, Fri83, CL98]. Synchronizing [KM89]. Synchronous [PH86]. syndrome [AH88]. synergies [JTW98]. Syntactic [Har88b]. Syntax [BDM^{+}93, vEHvV89]. Syntax-Directed [BDM^{+}93, vEHvV89]. Synthesis [CDJ^{+}84, JS99, OK94, LJ99, OHBR90]. synthetic [Kam89]. System [Amm91, ARAS94, Bar92, BW96, BE81, BG96, Ber88, Bol97b, BAL81, BBØ96, CHB94, Coo81, DR84, Dam96, DK94, DF84, ES85, Fai85a, FC96, FK92, Hać86a, Haç89a, HJ90a, HJS91, Joy94, KS96, LO92, Loh84, Mai96, MS81, MBCD86, MG81, Mer87, MIH92, Moy96, NS87, OHK93, OT92, Pha94, Pla92, Pow86, PW92, Rec93, RB93b, RT93, RA96, RF84, SGL93, Sam93, SW94a, Sku91, Sne83, Stu83, SCK86, TC93, TKS95, TW95, Var91, Whe81, Wic92, WSR^{+}83, WTS95, YNDS88, YCGH92, Zho94, ZM96, Zim84, vS83, vC80, Bak88, CB89a, DFCR96, Fic89, GAWC91, GAK92, HLAB99, HHL^{+}97, JM96, Kar94, Kar98, Ken80, KJB97, KJ99, LHC95, LP93, LM96, Nec96, OHBR90, OBS79, PL99, PM94, PMB99, PP94, Rey89, Sal80, SW96, WC99, ZG97]. Systems [Art87, Bar86, BW83, Bha84, Blu86, BAL81, BT97, BM83, CL94, CZ91, CLO95, Col92, DS94, DR92, DLG96, DV94, Eli92, Emd91, FMP86, FSA87, FM90b, FM87, Fur93, GMM90, Gla92a, Gla96a, Gla97a, Gom94, Gom95, GC94, GDF86, Hać86b, Hać89b, HFK92, Jef91, Jos83, KO95, KB96, Ker92, KP93, Kor83, Kri93, KNT86, Lan98a, Lea95, Li99, LSD95, LVB^{+}93, MW95, MR83, MO90, Mor86, MMSH92, MP90, Mue86, MP95, NC96, Nit96, OG80, PdC94, PdF97, PH86, PL96, Pop92, PZ94, Pre95, Rah92, RW97, Rei90a, RT86, Sag95, Sah94, SAASA94, San95, Sch91, Sel93, SKF95, She90, SM92a, Sta85, Sta90, SP94, SYB97, TT93, Ulu95, Ulu97, Ura90, WSN92, WNSC96, WM96, YP94, ZEB88, ZCd96, vS96, Bar94, CZdV98, Car94, ÇZUB99, CL99, Chu97]. systems [CDDF99, DMV98, DY99, DLT99, FW90, Gie79, Gla94a, Gla95c, Gla98b, Gla98g, Gla98h, Gla98d, Gla99a, Gla99b, Gla00, GP98, Hal92, Has98, Hoa94, HBJ^{+}99, ISS98, Jia99, Kam89, KY92, Kor99b, KM89, Lai99, yLcY98, LS99, Mar81, MR99, MD89, NL99, NQ98, Ost92, PM99, PNK96, Pla95, RC89, RG79, SW95a, SZ98, ST89, SMCL96, SMU98, Sta99, TT98, TW98, TS89, TM98, TVK94, Ulu98, WSM^{+}95, Wey99, WK88, WCV^{+}98, WM99, ZS88, ZMAER99, ZZ88]. systems-centric [LS99].

Table [Har81]. Table-Driven [Har81].

tables [JC98]. tabulation

[Ano94d, Gla94b]. TACFIRE [Sal80]. Tale [SM92a]. tandem [Kar94]. Targeting [AP97, Lut96]. Tarilan [De 98]. TarTAn [PL96]. Task [Kar98, DS98, PM99, TW98]. Tasking [Dil91, SC88]. Tasks [ML95, ZR87]. Taxonomy [BC94, GV92, OC91, TK87, Dav88, MC98, War89]. Teaching [Mur99, Tom89]. team [HS99]. Technical [ANB93, Ebe99, Sku91, KKiMT96]. Technique [DG92, Hen88, KC96, KL96, LTT92, MvS95, Nit96, WK88, WCV^{+}98, vdBK94]. Techniques [BS93, FWD97, Lak97, MS97, Mil96b, Pan81, PS90, Rey80, Sel93, YRN80, Ano93e, CW89, Eri92, Kam95, Kor99a, Wey99]. Technologies [Bas97, Gla98j, Sta93a, Ano96m]. Technology [APL95, Bro81, CFSS98, DA86, Gla88b, JVP^{+}98, KS96, MR80, Par98, Rv93, ZC97, Zuc90b, DS98, DF99, Gla88a, Gla89b, Har97, Pfl99, Sai98, SSvdW99, WDMR99]. telecom [VVS99]. Telecommunications [Gas96]. television [Bra89]. Template [ZSGS93]. Template/Module [ZSGS93]. Temporal [Jma96, UH96, Gla89i, MP94, PM94, VT99]. Tension [Gla89f]. Tenth [FM90b]. Tenure [AP97]. termination [MC98]. Test [BCFG86, Dye93, FLN91, LH90, MS81, OKOM97, Pas96, Sam93, Sed93, Tia96, Vel87, WHMP99, CLSC98, JF99, NS92, OL99]. Testability [VM93]. Testable [BL95]. Testing [Ber91, Ber93, BM96, BMP97, BM93b, Car96, CLSC98, DGM93, FZ93, Gla93g, Gla93i, HZ84, Har99, How80, JVP^{+}98, JvB83, KGH^{+}96, Las90, MG81, Mil96b, OH94, PBC93, Pla92, PU84a, SCK86, Zei88, FWH97, Jen99, Lai99, Leu97, PU84b, WHL89, WM95]. Texas [CR89, MP89]. Text [Fis91, Ree85, TOYI95]. Text-Oriented [TOYI95]. Texts [Yan94]. Textual [HG91, Sny91]. Their [AR94, Ber93, Car96, LVB^{+}93, MD91, Eri92, SW88]. theoretic [MJ89]. Theory [Gla90h, KAL97, KP93, yL98, Rv91, Ano94d, BM89, CTZ92, Gla89i, Gla94b, Gla95b, LJ99, Gla93c, Gla94f, Gla94h]. Theory-Based [KAL97]. There [Gla90d, Gla91h, Gla95f, Gla97g, Gla95f]. things [Gla90f]. think [Gla93h]. thoughts [Gla89g, Gla94e, Gla97i, Gla97h]. Thread [ISS98]. threads [WCV^{+}98]. Threat [Rei90b, Zuc90a]. Three [MPS86, SI94, ZMAER99]. Three-Dimensional [MPS86, ZMAER99]. thriving [Gla97b]. throughout [Tia99]. Time [AQ90, BP86, CL94, Chr86, Cla86, CRV94, GMM90, GMP94, Gla91e, Gom89, Gom94, GRS92, HW94, HFK92, wLyLlH97, LM94, Leu92, LH95, ML95, NC96, OG80, OK94, PZ94, Rei90a, Ulu95, Ulu97, WM96, Yua90, ZCd96, ZR87, BFR96, BG98, Bak88, CKyL98, DMV98, DY99, Gla97g, Hal92, Hoa94, KY92, KR98, Kor99b, KKiMT96, yL98, yLcY98, LF91, LP93, MO84, Ost92, Ozk97, PNK96, RG79, SBB98, Sto92, Ulu98, VT98, VT99, XH98, ZLZ^{+}96]. time-critical [Ozk97, SBB98]. time-honored [Gla97g]. Time/Cost [LM94]. Timed [Chr86, FZHS95, LVB^{+}93, WM96, JS99]. Timed-Event [Chr86]. Timed-Probabilistic [FZHS95]. Timing [GMP94, PdF97, Sah94, VT98]. TIMS [SGL93]. Token [TW95]. Token-Based [TW95]. Tolerance [Ban86, Fri90, KN97, KP93, SAASA94, WWF94, ZX94, Hoa94, Shu99]. Tolerant [BW95, CG94, DG92, MS90, Mor86, Mue86, OK94, PdC94, Ram90, WTS95, WFZ96, JM96, LFY^{+}99, SMCL96, ZG97]. tongue [Gla91g]. tongue-in-cheek [Gla91g]. too [Mor99]. Tool [BN90, Bro87, FS88, FM93, FG93, GA95, IYKO95, LZL97, ML95, NY84, NB93, OC90, Rei90a, Rid81, TTP97, HLAB99, Rey89, Son93]. toolkit [Rob98, TCMJ98]. Tools [AM85, BYY87, Hen95, HO96, JP94, KP91, TKS95, TM97, Zim84, vAW93, Ano88d, DS98, Eri92, Ni97, PKK98, TCMJ98]. Toolset [WH91b, MSHB98]. Top [MM81, Gla97j]. Top-Down [MM81]. Topic [Gla92a]. topics [Ano94d, Gla94b]. Traceability [DF84, OC90, Ni97]. Tracing [LK93]. Tracking [Tia96]. Tractable [Nit96]. Trade [GA95, Rec93]. Trade-Offs [GA95]. Traditional [GC94, Jac98]. Traffic [PH93, PV94, VVS99]. Training [AP97, BN90, MP89, MMSH92, Rus90, Sai99]. Transaction [KR98, KW91, KW93, Lai97a, Rah92, RW97, SW94a, Ulu95, Ulu98]. Transactions [BR90, HW94, ZM96, KWME99]. Transfer [Gla88b, HJ91, JVP^{+}98, Par98, Gla88a, Pfl99, Sai98]. Transferring [HBCC94]. Transformation [GA95, Lan98b]. Transformational [MB84]. Transformations [AR94]. transition [Dav95]. Transitive [Fra86]. Translation [JN84, UhCLS94]. Translator [HL93]. Translator-Based [HL93]. transmission [Ng99]. transparency [DFCR96]. Transport [Fai85a]. Transporting [BP80]. Transposition [HP92]. Transputer [Row86, YY93]. Transputer-Based [YY93]. trash [Gla98f]. Tree [LH83, CV95]. Trees [PS90, BF96, LM96, WH99]. trends [Ano95h, Har98, ZS95]. Triad [Zim84]. Triangular [RT86]. TRIO [GMM90]. trivial [Gla89d]. trivial/brilliant [Gla89d]. truly [Gla89c]. Trusted [RT93]. Tuning [LZL97]. Tunisian [FM90b]. Tunisian-French [FM90b]. tuple [PA99]. tuple-space [PA99]. Turning [Gla95j]. TutORing [KP97a]. Two [CFK91, Chr86, DHP86, Gla97k, KCK^{+}98, MF90, MRW^{+}94, Sah94, SM92a, TC93, Ulu95, ZMAER99, HY95, Kar94]. Two-Axis [Sah94]. Two-Disk [TC93]. Two-Person [KCK^{+}98]. two-phase [HY95]. Two-Version [CFK91]. Type [Bel91]. Typed [Gan91]. typology [KJB97].

U.S. [Sny79, Tha80, Zuc90b]. UK [TE99]. unbalanced [PV94]. Unbounded [FN86]. Unconstrained [Ber93]. undergraduate [SSvdW99]. Understanding [AH88, Gla93d, GA95, Lak93, Leh80, LF96, Pfl99, BM89, FTAM99]. Understands [Gla92b]. unethical [FF89]. Unified [BFR96, Gon95, Lak97, Mai96]. Uniform [WWF94]. Unique [Gla97f]. United [Duv95]. Units [Joy87, BM98]. university [MBL^{+}99, Bra89, CR89, Mil89]. UNIX [WLC95, Bar86]. Unix-Based [Bar86]. UNIX^{T M} [Ni97]. unsolved [Ano91c, Gla91i]. Updating [FS91]. Usability [HHSR94, NL99, SKF95]. Usage [BAH96, CM93, WH91a]. Use [AB90, ARAS94, BGB90, Got92b, HZ83, Ham81, JM90, Kal92, KML94, MGJT87, RBM95, SL80, SB88, YN91, Bev99, GP98, JNY84, Rob98]. used [CB89a, Tha80]. User [BAL81, CM93, Deh90, GK91b, HHSR94, Hur93, KJ99, KC98, LG97, RAC90, Rei87, Rv91, Bev99, PL94, Gla94i]. User-Computer [GK91b]. User-oriented [Rv91]. Users [AH81, Moy96]. uses [FWH97]. Using [BST93, BCFG86, CL81, Cha91, DGM93, DJL93, DS85, FC96, FWP93, FWD97, FAS94, GI95, Gor91, Hać86a, HJ90a, Har90a, HOT97, HG91, HS99, HUMT92, Jma96, Joy87, KB98, KD91, KP93, Lai97a, Lai97c, LL97a, Lee93, LTT92, Mar84, MB97, MTW97, OC90, PD98, Por93, Rv91, Rv92, Sca88, SG91, Sch91, Sta90, TC93, Uck91, WH99, WNHM86, CC99b, CDDF99, HCC91, HB89, JS99, Lai95, yLcY98, LH98, LM96, LJM96, NS92, PNK96, VVS99]. Utility [AH90, Rv91]. utilization [PNK96, Zel88]. utilizing [ES97].

Vadis [MWH97]. Validated [Hać89b]. Validating [GMP94, LH95]. Validation [BS93, EC98, Pas96, Ano93e]. Value [Gon95]. valued [VL94]. VDM [BM93a]. Verification [BS93, Dil91, EC98, GC94, JL97, KO95, LL97a, LF96, NS92, Nit96, Nit98, NS83, TK91, Ano93e, HZ79, HHZ92, JC98, OBS79]. Verify [LL98]. verifying [Lai97d]. Verrall [Lit80]. Version [CFK91, Pha94, EL88]. versus [Gla90h, ZS88]. Very [ZR94, Wey99]. via [Ano87d, BB89, Gla86, SMU98, TVK95]. video [Ng99]. View [Gla97m, LIC92, HR95, OSG98, SBA97, VC97]. Viewed [Del92]. Views [Lan90, TKU93, Uck91]. Violence [SM92b]. Virtual [LTT92]. Virus [DG87, Gla89e]. viruses [Thi94]. visibility [OBS79]. Visual [Kun95, Ng93, WM90, DB95]. Visualization [KM92, MTW97]. visualize [KB98]. vital [Ano88d]. VLSI [CDJ^{+}84, HHZ92, HD84, KSS84, MB84, Rad84, RT86]. Volume [Ano97m, Ano97n, Ano97o, Ano98f, Blu89, Ano85a, Ano85c]. Voting [JT97]. Vs [Sca88, BBP96, FWH97, Gla91c, Rei90b, Zuc90b, Zuc90a].

warning [Gla98e]. Way [Gla92h, Wey99]. Weapon [Coo81, Stu83, Gie79, Sal80]. Web [HMP99, KWME99]. Webwork [Gla98d]. We’d [OT92]. Well [Hen88, LS98]. Well-Known [Hen88]. WEP [CP88]. We’re [Zuc90b]. we’ve [Gla93f]. Where [Gla92b, Ano94d, FF96, Gla94b]. Which [Gla93i, Gla93h]. While [Hen88]. White [HZ84, WL99]. Who [Gla92b, Gla98j, Har95a]. Who’s [Gla88c]. Wide [HMP99]. Will [Par98, Gla93h]. Willingness [YN91]. WINDOWS^{T M} [Ni97]. Wire [DV94]. wisdom [JC98]. Within [SF92, BC94, Sed93]. without [JC98, KK85]. Women [CFSS98]. word [Gla94d, Gla98e]. Work [Tau80, WT89]. Workload [HFK92, Rah92]. Workshop [Ano84c, Goe84, Har90b, Got90, Sch81]. workspace [GAW92]. workstation [CP88]. Workstations [Sho91]. world [SSvdW99, HMP99]. worst [Gla92d]. Write [Hać91, CB89a, DF98, Gla93a, Gla94e]. write-once [CB89a]. Writes [FC96]. Written [WNHM86]. wrong [Ano87h]. WWW [CJP98]. WWW-based [CJP98].

X [CM86]. X-MP [CM86].

Y2K [Gla98f, Gla98k]. year [Gla89b]. Years [Blu86, Bux90, Sta93a, Gla97k].

Zero [CL97]. Zero-One [CL97]. Ziegler [CBK96]. ZONE [UH96]. Zucconi [Rei90b, Zuc90b, Zuc90a].

References

Alherbish:1998:HPA

[AA98] Jasir Alherbish and Reda Ammar. High-performance Arabic character recognition. The Journal of Systems and Software, 44(1):53–71, December 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                                   URL http://www.elsevier.
                                                   com/cas/tree/store/jss/sub/
                                                   1998/44/1/6071.pdf.

Aguilera:1990:URP

[AB90] Christine Aguilera and Daniel M. Berry. The use of a repeated phrase finder in requirements extraction. The Journal of Systems and Software, 13(3):209– 230, November 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Achee:1997:COO

[AC97] B. L. Achee and Doris L. Carver. Creating object-oriented designs from legacy FORTRAN code. The Journal of Systems and Software, 39(2):179–194, November 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ambriola:1991:TIS

[ACCD91] Vincenzo Ambriola, Paolo Ciancarini, Andrea Corradini, and Nicoletta DeFrancesco. Towards innovative software engineering environments. The Journal of Systems and Software, 14(1):17– 29, January 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Andersson:1996:OIA

[AF96] Niclas Andersson and Peter Fritzson. Overview and industrial application of code generator generators. The Journal of Systems and Software, 32 (3):185–214, March 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahituv:1981:DRC

[AH81] Niv Ahituv and Jonathan Halpern. Data and reports — contents design and users allocation. The Journal of Systems and Software, 2(3):193–199, September 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abdel-Hamid:1988:USS

[AH88] Tarek K. Abdel-Hamid. Understanding the “90% syndrome” in software project management: a simulation-based case study. The Journal of Systems and Software, 8(4):319–330, September 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abdel-Hamid:1990:UHP

[AH90] Tarek K. Abdel-Hamid. On the utility of historical project statistics for cost and schedule estimation: Results from a simulationbased case study. The Journal of Systems and Software, 13(1):71– 82, September 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abdel-Hamid:1993:MPS

[AH93] Tarek K. Abdel-Hamid. A multiproject perspective of singleproject dynamics. The Journal of Systems and Software, 22(3): 151–165, September 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Al-Haddad:1993:IIM

[AHG93] H. M. Al-Haddad and K. M. George. An implementation inheritance model for objectoriented programming. The Journal of Systems and Software, 20(1):3–17, January 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Al-Haddad:1992:FIM

[AHGS92] H. M. Al-Haddad, K. M. George, and M. H. Samadzadeh. A feedback inheritance model. The Journal of Systems and Software, 18(3):219–234, July 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alonso:1996:SEK

[AJMP96] F. Alonso, N. Juristo, J. L. Maté, and J. Pazos. Software engineering and knowledge engineering: Towards a common life cycle. The Journal of Systems and Software, 33(1):65–79, April 1996.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abbott:1981:SRS

[AM81] Russell J. Abbott and D. K.

Moorhead. Software requirements and specifications: a survey of needs and languages. The Journal of Systems and Software, 2(4):297–316, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ambriola:1985:AGE

[AM85] Vincenzo Ambriola and Carlo Montangero. Automatic generation of execution tools in a Gandalf environment. The Journal of Systems and Software, 5(2): 155–171, May 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Azuma:1994:SMP

[AM94] Motoei Azuma and David Mole.

Software management practice and metrics in the European community and Japan: Some results of a survey. The Journal of Systems and Software, 26(1): 5–18, July 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ambler:1987:EFL

[Amb87] Allen L. Ambler. Experience with a functionally layered multicomputer architecture for interactive processing. The Journal of Systems and Software, 7(4):

267–277, December 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ammar:1989:SBS

[Amm89] Reda A. Ammar. Stack-based sorting algorithms. The Journal of Systems and Software, 9 (3):225–239, March 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ammar:1991:CAD

[Amm91] Reda A. Ammar. A computer aided design system to develop high performance software. The Journal of Systems and Software, 15(2):139–147, May 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Arthur:1993:ESD

[ANB93] James D. Arthur, Richard E. Nance, and Osman Balci. Establishing software development process control: Technical objectives, operational requirements, and the foundational framework. The Journal of Systems and Software, 22(2):117–128, August 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1979:B

[Ano79] Anonymous. Biography. The Journal of Systems and Software, 1(1):103–104, ???? 1979. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1980:AI

[Ano80a] Anonymous. Author index. The Journal of Systems and Software, 1(4):347–348, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1980:Ba

[Ano80b] Anonymous. Biography. The Journal of Systems and Software, 1(2):177–178, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1980:Bb

[Ano80c] Anonymous. Biography. The Journal of Systems and Software, 1(3):267–269, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1980:SI

[Ano80d] Anonymous. Subject index. The Journal of Systems and Software, 1(4):349, ???? 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1981:AI

[Ano81a] Anonymous. Author index. The Journal of Systems and Software, 2(4):381, December 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1981:Ba

[Ano81b] Anonymous. Biographies. The Journal of Systems and Software, 2(1):81–82, February 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1981:Bb

[Ano81c] Anonymous. Biographies. The Journal of Systems and Software, 2(4):379–380, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1981:SI

[Ano81d] Anonymous. Subject index. The Journal of Systems and Software, 2(4):383, December 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1983:EI

[Ano83] Anonymous. Editor’s introduction. The Journal of Systems and Software, 3(4):iii–iv, December 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1984:AI

[Ano84a] Anonymous. Author index. The Journal of Systems and Software, 4(4):333, November 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1984:B

[Ano84b] Anonymous. Biographies. The Journal of Systems and Software, 4(2–3):265, July 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1984:ISI

[Ano84c] Anonymous. Introduction to the special issue on the Fifth Minnowbrook Workshop on Software Performance Evaluation. The Journal of Systems and Software, 4(4):iii–v, November 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1984:SI

[Ano84d] Anonymous. Subject index. The Journal of Systems and Software, 4(4):335, November 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1985:AIV

[Ano85a] Anonymous. Author index to volume 5. The Journal of Systems and Software, 5(4):305, November 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1985:B

[Ano85b] Anonymous. Biographies. The Journal of Systems and Software, 5(4):303, November 1985. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1985:SIV

[Ano85c] Anonymous. Subject index to volume 5. The Journal of Systems and Software, 5(4):307, November 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1986:AI

[Ano86a] Anonymous. Author index. The Journal of Systems and Software, 6(4):393, November 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1986:EC

[Ano86b] Anonymous. Editor’s corner. The Journal of Systems and Software, 6(1–2):3, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1986:E

[Ano86c] Anonymous. From the Editor. The Journal of Systems and Software, 6(1–2):1, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1986:SSM

[Ano86d] Anonymous. Setting standards in mainframe applications software. The Journal of Systems and Software, 6(3):295–305, August 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1986:SI

[Ano86e] Anonymous. Subject index. The Journal of Systems and Software, 6(4):391, November 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1987:AI

[Ano87a] Anonymous. Author index. The Journal of Systems and Software, 7(4):371, December 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1987:Ba

[Ano87b] Anonymous. Biographies. The Journal of Systems and Software, 7(1):85, March 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1987:Bb

[Ano87c] Anonymous. Biographies. The Journal of Systems and Software, 7(4):370, December 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1987:ECN

[Ano87d] Anonymous. Editor’s corner: No silver bullet: a look at software research via the Fred Brooks article. The Journal of Systems and Software, 7(3):181–182, September 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1987:HWP

[Ano87e] Anonymous. hould we prepare for a 4GL future? The Journal of Systems and Software, 7(1): 1–2, March 1987. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1987:SED

[Ano87f] Anonymous. Standards and enforcers: Do they really help achieve software quality? The Journal of Systems and Software, 7(2):87–88, June 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1987:SI

[Ano87g] Anonymous. Subject index. The Journal of Systems and Software, 7(4):372, December 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1987:WRW

[Ano87h] Anonymous. What’s really wrong with COBOL? The Journal of Systems and Software, 7(4):251–252, December 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1988:AI

[Ano88a] Anonymous. Author index. The Journal of Systems and Software, 8(5):434, December 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1988:Ba

[Ano88b] Anonymous. Biographies. The Journal of Systems and Software, 8(3):255, June 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1988:Bb

[Ano88c] Anonymous. Biographies. The Journal of Systems and Software, 8(5):431, December 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1988:MVL

[Ano88d] Anonymous. A mundane but vital look at software tools. The Journal of Systems and Software, 8(2):75–76, March 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1988:SI

[Ano88e] Anonymous. Subject index. The Journal of Systems and Software, 8(5):433, December 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1989:AIa

[Ano89a] Anonymous. Author index. The Journal of Systems and Software, 9(4):314, May 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1989:AIb

[Ano89b] Anonymous. Author index. The Journal of Systems and Software, 10(4):289, November 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1989:Ba

[Ano89c] Anonymous. Biographies. The Journal of Systems and Software, 9(4):311, May 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1989:Bb

[Ano89d] Anonymous. Biographies. The Journal of Systems and Software, 10(2):155–156, September 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1989:Bc

[Ano89e] Anonymous. Biographies. The Journal of Systems and Software, 10(3):227, October 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1989:Bd

[Ano89f] Anonymous. Biographies. The Journal of Systems and Software, 10(4):285–286, November 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1989:SIa

[Ano89g] Anonymous. Subject index. The Journal of Systems and Software, 9(4):313, May 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1989:SIb

[Ano89h] Anonymous. Subject index. The Journal of Systems and Software, 10(4):287–288, November 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1990:AI

[Ano90a] Anonymous. Author index. The Journal of Systems and Software, 13(3):239–241, November 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1990:Ba

[Ano90b] Anonymous. Biographies. The Journal of Systems and Software, 11(1):75, January 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1990:Bb

[Ano90c] Anonymous. Biographies. The Journal of Systems and Software, 11(3):225–226, March 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1990:ECM

[Ano90d] Anonymous. Editor’s corner:

Mikhail Gorbachev and software productivity (!?). The Journal of Systems and Software, 13(3): iii–viii, November 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1990:SI

[Ano90e] Anonymous. Subject index. The Journal of Systems and Software, 13(3):237–238, November 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1991:AI

[Ano91a] Anonymous. Author index. The Journal of Systems and Software, 16(3):265–266, November 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1991:ECSb

[Ano91b] Anonymous. Editor’s corner:

Software productivity: What is management’s role? The Journal of Systems and Software, 16(3):i– iv, November 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1991:ECSa

[Ano91c] Anonymous. Editor’s corner:

The (solved, unsolved) problem of literature searches. The Journal of Systems and Software, 15 (3):iii–v, July 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1991:SI

[Ano91d] Anonymous. Subject index. The Journal of Systems and Software, 16(3):267–268, November 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:AI

[Ano92a] Anonymous. Author index. The Journal of Systems and Software, 19(3):291–292, November 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:Ba

[Ano92b] Anonymous. Biographies. The Journal of Systems and Software, 18(1):109, April 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:Bb

[Ano92c] Anonymous. Biographies. The Journal of Systems and Software, 18(2):201–202, May 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:CPA

[Ano92d] Anonymous. Calls for papers and announcements. The Journal of Systems and Software, 18 (3):287–288, July 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:CC

[Ano92e] Anonymous. Controversy corner. The Journal of Systems and Software, 18(1):101, April 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:ECIa

[Ano92f] Anonymous. Editor’s corner:

The importance of software quality in the 1990s. The Journal of Systems and Software, 17(3): ii–iv, March 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:ECIb

[Ano92g] Anonymous. Editor’s corner:

The importance of the application domain cluster. The Journal of Systems and Software, 18 (3):iii–iv, July 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:RCa

[Ano92h] Anonymous. Readers’ corner.

The Journal of Systems and Software, 17(2):107, February 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:RCb

[Ano92i] Anonymous. Readers’ corner.

The Journal of Systems and Software, 18(1):1–2, April 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1992:SI

[Ano92j] Anonymous. Subject index. The Journal of Systems and Software, 19(3):293–294, November 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1993:AI

[Ano93a] Anonymous. Author index. The Journal of Systems and Software, 23(3):276, December 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1993:CPa

[Ano93b] Anonymous. Call for papers. The Journal of Systems and Software, 20(1):101–102, January 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1993:CPb

[Ano93c] Anonymous. Call for papers. The Journal of Systems and Software, 22(2):147–148, August 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1993:CPc

[Ano93d] Anonymous. Call for papers. The Journal of Systems and Software, 22(3):225, September 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1993:ECA

[Ano93e] Anonymous. Editor’s corner: Applying specification, verification, and validation techniques. The Journal of Systems and Software, 21(3):i–iii, June 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1993:ECD

[Ano93f] Anonymous. Editor’s corner: In defense of adhocracy. The Journal of Systems and Software, 22 (3):i–ii, September 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1993:GEI

[Ano93g] Anonymous. Guest Editor’s introduction. The Journal of Systems and Software, 20(3):i–iii, March 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1993:SI

[Ano93h] Anonymous. Subject index. The Journal of Systems and Software, 23(3):277–278, December 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1994:AI

[Ano94a] Anonymous. Author index. The Journal of Systems and Software, 27(3):253–254, December 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1994:Ba

[Ano94b] Anonymous. Biographies. The Journal of Systems and Software, 27(1):69–70, October 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1994:Bb

[Ano94c] Anonymous. Biographies. The Journal of Systems and Software, 27(2):155–156, November 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1994:ECT

[Ano94d] Anonymous. Editor’s corner: a tabulation of topics where software practice leads software theory. The Journal of Systems and Software, 25(3):i–ii, June 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1994:ECD

[Ano94e] Anonymous. Editor’s corner: discipline and creativity: an odd couple. The Journal of Systems and Software, 26(3):i–ii, September 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1994:GEC

[Ano94f] Anonymous. Guest Editors’ corner. The Journal of Systems and Software, 27(3):i–ii, December 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1994:GEI

[Ano94g] Anonymous. Guest Editor’s introduction. The Journal of Systems and Software, 24(3):i–ii, March 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1994:SI

[Ano94h] Anonymous. Subject index. The Journal of Systems and Software, 27(3):255–256, December 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1995:AI

[Ano95a] Anonymous. Author index. The Journal of Systems and Software, 31(3):277–278, December 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1995:Ba

[Ano95b] Anonymous. Biographies. The Journal of Systems and Software, 29(2):195–196, May 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1995:Bb

[Ano95c] Anonymous. Biographies. The Journal of Systems and Software, 29(3):271–272, June 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1995:Bc

[Ano95d] Anonymous. Biographies. The Journal of Systems and Software, 30(1–2):164–166, July/ August 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1995:Bd

[Ano95e] Anonymous. Biographies. The Journal of Systems and Software, 30(3):303–304, September 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1995:Be

[Ano95f] Anonymous. Biographies. The Journal of Systems and Software, 31(1):93, October 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1995:Bf

[Ano95g] Anonymous. Biographies. The Journal of Systems and Software, 31(2):181, November 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1995:GEC

[Ano95h] Anonymous. Guest Editors’ corner: Software reuse: Current status and trends. The Journal of Systems and Software, 30(3):i–ii, September 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1995:SI

[Ano95i] Anonymous. Subject index. The Journal of Systems and Software, 31(3):279–280, December 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:AI

[Ano96a] Anonymous. Author index. The Journal of Systems and Software, 35(3):223–224, December 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Ba

[Ano96b] Anonymous. Biographies. The Journal of Systems and Software, 32(1):97–98, January 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Bb

[Ano96c] Anonymous. Biographies. The Journal of Systems and Software, 32(2):181, February 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Bc

[Ano96d] Anonymous. Biographies. The Journal of Systems and Software, 32(3):309–310, March 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Bd

[Ano96e] Anonymous. Biographies. The Journal of Systems and Software, 33(1):109–110, April 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Be

[Ano96f] Anonymous. Biographies. The Journal of Systems and Software, 33(2):203–204, May 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Bf

[Ano96g] Anonymous. Biographies. The Journal of Systems and Software, 33(3):329–330, June 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Bg

[Ano96h] Anonymous. Biographies. The Journal of Systems and Software, 34(1):83–84, July 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Bh

[Ano96i] Anonymous. Biographies. The Journal of Systems and Software, 34(2):171–172, August 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Bi

[Ano96j] Anonymous. Biographies. The Journal of Systems and Software, 34(3):255–256, September 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Bj

[Ano96k] Anonymous. Biographies. The Journal of Systems and Software, 35(1):91–92, October 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:Bk

[Ano96l] Anonymous. Biographies. The Journal of Systems and Software, 35(2):165–166, November 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:CPE

[Ano96m] Anonymous. Call for papers evaluation of reverse engineering and reengineering technologies. The Journal of Systems and Software, 33(3):331, June 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1996:SI

[Ano96n] Anonymous. Subject index. The Journal of Systems and Software, 35(3):225–226, December 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:AI

[Ano97a] Anonymous. Author index. The Journal of Systems and Software, 39(3):293–294, December 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Ba

[Ano97b] Anonymous. Biographies. The Journal of Systems and Software, 36(2):201–202, February 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Bb

[Ano97c] Anonymous. Biographies. The Journal of Systems and Software, 36(3):315–316, March 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Bc

[Ano97d] Anonymous. Biographies. The Journal of Systems and Software, 37(1):97–98, April 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Bd

[Ano97e] Anonymous. Biographies. The Journal of Systems and Software, 37(2):175, May 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Be

[Ano97f] Anonymous. Biographies. The Journal of Systems and Software, 38(1):103–104, July 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Bf

[Ano97g] Anonymous. Biographies. The Journal of Systems and Software, 38(2):193–194, August 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Bg

[Ano97h] Anonymous. Biographies. The Journal of Systems and Software, 38(3):279–280, September 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Bh

[Ano97i] Anonymous. Biographies. The Journal of Systems and Software, 39(1):89–90, October 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Bi

[Ano97j] Anonymous. Biographies. The Journal of Systems and Software, 39(2):195–196, November 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:Bj

[Ano97k] Anonymous. Biographies. The Journal of Systems and Software, 39(3):291–292, December 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:SI

[Ano97l] Anonymous. Subject index. The Journal of Systems and Software, 39(3):295–299, December 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:VCa

[Ano97m] Anonymous. Volume contents.

The Journal of Systems and Software, 37(3):265–266, June 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:VCb

[Ano97n] Anonymous. Volume contents.

The Journal of Systems and Software, 38(3):281–283, September 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1997:VCc

[Ano97o] Anonymous. Volume contents.

The Journal of Systems and Software, 39(3):301–302, December 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1998:Ba

[Ano98a] Anonymous. Biographies. The Journal of Systems and Software, 40(1):89–90, January 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1998:Bb

[Ano98b] Anonymous. Biographies. The Journal of Systems and Software, 40(2):185–186, February 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1998:Bc

[Ano98c] Anonymous. Biographies. The Journal of Systems and Software, 40(3):297–299, March 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1998:Ec

[Ano98d] Anonymous. Editorial. The Journal of Systems and Software, 42(1):1–2, July 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1998:Ej

[Ano98e] Anonymous. Editorial. The Journal of Systems and Software, 44(2):85–86, December 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1998:VC

[Ano98f] Anonymous. Volume contents. The Journal of Systems and Software, 40(3):300–302, March 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1999:Ea

[Ano99a] Anonymous. Editorial. The Journal of Systems and Software, 44(3):169–170, January 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1999:Eb

[Ano99b] Anonymous. Editorial. The Journal of Systems and Software, 45(1):1–2, February 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1999:Ec

[Ano99c] Anonymous. Editorial. The Journal of Systems and Software, 45(2):79–80, March 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1999:Ed

[Ano99d] Anonymous. Editorial. The Journal of Systems and Software, 45(3):173–175, March 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1999:Ee

[Ano99e] Anonymous. Editorial. The Journal of Systems and Software, 46(2–3):89, April 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1999:Ef

[Ano99f] Anonymous. Editorial. The Journal of Systems and Software, 47(2–3):71–72, July 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1999:Eg

[Ano99g] Anonymous. Editorial. The Journal of Systems and Software, 47(2–3):73, July 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:1999:Eh

[Ano99h] Anonymous. Editorial(s). The Journal of Systems and Software, 48(1):1–3, August 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.elsevier.nl/gej-ng/10/
29/11/37/17/17/;   http://
www.elsevier.nl/gej-ng/10/
29/11/37/17/17/article.pdf.

Anonymous:1999:El

[Ano99i] Anonymous. Editorial(s). The Journal of Systems and Software, 48(2):77–78, October 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.elsevier.nl/gej-ng/10/
29/11/37/23/17/;   http://
www.elsevier.nl/gej-ng/10/
29/11/37/23/17/article.pdf.

Anonymous:1999:En

[Ano99j] Anonymous. Editorial(s). The Journal of Systems and Software, 48(3):163–164, November 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/

29/11/37/24/;   http://www.
elsevier.nl/gej-ng/10/29/
11/37/24/17/article.pdf.

Anonymous:1999:Ep

[Ano99k] Anonymous. Editorial(s). The Journal of Systems and Software, 49(1):1–2, December 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/25/25/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/25/abstract.
html.

Anonymous:1999:Er

[Ano99l] Anonymous. Editorial(s). The Journal of Systems and Software, 49(1):87, December 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/25/33/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/25/abstract.
html.

Agarwal:1997:TCP

[AP97] Ritu Agarwal and Jayesh Prasad.

Targeting COBOL programmers for C training: The role of job insecurity and organizational tenure. The Journal of Systems and Software, 37(1):5–17, April 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahrens:1995:SPR

[APL95] Judith D. Ahrens, Noah Prywes, and Evan Lock. Software process reengineering: Toward a new generation of CASE technology. The Journal of Systems and Software, 30(1–2):71–84, July/ August 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ammar:1990:ADT

[AQ90] Reda A. Ammar and Bin Qin. An approach to derive time costs of sequential computations. The Journal of Systems and Software, 11(3):173–180, March 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Adamov:1990:PMS

[AR90] Rade Adamov and Lutz Richter. A proposal for measuring the structural complexity of programs. The Journal of Systems and Software, 12(1):55–70, April 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abran:1994:FPS

[AR94] Alain Abran and Pierre N. Robillard. Function points: a study of their measurement processes and scale transformations. The Journal of Systems and Software, 25(2):171–184, May 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Arafeh:1995:GGM

[Ara95] Bassel R. Arafeh. A graph grammar model for concurrent and distributed software specification-in-large. The Journal of Systems and Software, 31 (1):7–32, October 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ashqar:1994:UGS

[ARAS94] Abdelhaleem Ashqar, Brian J.

Reithel, Milam W. Aiken, and Ashraf Shirani. Use of a group support system to evaluate management information system effectiveness. The Journal of Systems and Software, 24(3):267– 275, March 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Archibald:1981:ESE

[Arc81] Jerry L. Archibald. External structure: Experience with an automated module interconnection language. The Journal of Systems and Software, 2(2):147– 157, June 1981. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Armstrong:1998:IIG

[Arm98] Jim Armstrong. Industrial integration of graphical and formal specifications. The Journal of Systems and Software, 40 (3):211–225, March 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Arthur:1987:TFS

[Art87] James D. Arthur. Toward a formal specification of menu-based systems. The Journal of Systems and Software, 7(1):73–82, March 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Arunachalam:1996:CPP

[AS96] Vairam Arunachalam and William Sasso. Cognitive processes in program comprehension: An empirical analysis in the context of software reengineering. The Journal of Systems and Software, 34(3):177–189, September 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Al-Saqabi:1996:RCF

[ASSA96] Khaled Al-Saqabi, Kassem Saleh, and Imtiaz Ahmad. Recovery from concurrent failures in communication protocols. The Journal of Systems and Software, 35(1):55–65, October 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Azadegan:1997:PJA

[AT97] Shiva Azadegan and Anand Tripathi. A parallel join algorithm for SIMD architectures. The Journal of Systems and Software, 39(3):265–280, December 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ayres:1998:NHD

[Ayr98] Ronald Ayres. A narrative history and description of MO-SIS software. The Journal of Systems and Software, 43(1): 67–82, October 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/1/6050.pdf.

Baber:1991:CCP

[Bab91] Robert L. Baber. Controversy corner: Portrait of a (software) engineer. The Journal of Systems and Software, 15(1):91–100, April 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Basili:1996:MDC

[BAEH96] Victor R. Basili and Salwa K.

Abd-El-Hafiz. A method for documenting code components. The Journal of Systems and Software, 34(2):89–104, August 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ben-Asher:1996:USD

[BAH96] Yosi Ben-Asher and Gadi Haber.

On the usage of simulators to detect inefficiency of parallel programs caused by “bad” schedulings: The SIMPARC approach. The Journal of Systems and Software, 33(3):313–327, June 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baker:1988:IAR

[Bak88] T. P. Baker. An improved Ada run-time system interface. The Journal of Systems and Software, 8(5):373–393, December 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Booth:1981:ISM

[BAL81] Taylor L. Booth, Reda Ammar, and Robert Lenk. Instrumentation system to measure user performance in interactive systems. The Journal of Systems and Software, 2(2):139–146, June 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Banino:1986:PFC

[Ban86] J. S. Banino. Parallelism and fault-tolerance in the chorus. The Journal of Systems and Software, 6(1–2):205–211, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barr:1986:UBG

[Bar86] W. J. Barr. Unix-based gateway to distributed database systems. The Journal of Systems and Software, 6(3):225–235, August 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barros:1992:PAC

[Bar92] Oscar Barros. A pragmatic approach to computer-assisted system building. The Journal of Systems and Software, 18(3): 235–244, July 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barros:1994:OOC

[Bar94] Oscar Barros. Object-oriented case-supported development of information systems. The Journal of Systems and Software, 24 (2):95–113, February 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Basili:1980:ISI

[Bas80] Victor R. Basili. Introduction to the special issue on Software Life Cycle Management. The Journal of Systems and Software, 1 (3):179, ???? 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Basili:1997:EPR

[Bas97] Victor R. Basili. Evolving and packaging reading technologies. The Journal of Systems and Software, 38(1):3–12, July 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Basili:1981:CPC

[BB81] Victor R. Basili and John Beane.

Can the Parr curve help with manpower distribution and resource estimation problems? The Journal of Systems and Software, 2(1):59–69, February 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Benander:1989:ESC

[BB89] Alan C. Benander and Barbara A. Benander. An empirical study of COBOL programs via a style analyzer: The benefits of good programming style. The Journal of Systems and Software, 10(4):271–279, November 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bieman:1988:SRI

[BBC^{+}88] James M. Bieman, Albert L.

Baker, Paul N. Clites, David A.

Gustafson, and Austin C. Melton. Standard representation of imperative language programs for data collection and software measures specification. The Journal of Systems and Software, 8(1):13–37, January 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baker:1990:PSM

[BBF^{+}90] Albert L. Baker, James M. Bieman, Norman Fenton, David A. Gustafson, Austin Melton, and Robin Whitty. A philosophy for software measurement. The Journal of Systems and Software, 12(3):277–281, July 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Balbo:1986:SHQ

[BBG86] G. Balbo, S. C. Bruell, and S. Ghanta. Solution of homogeneous queueing networks with many job classes. The Journal of Systems and Software, 6(1–2): 41–53, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Braa:1996:ORP

[BBØ96] Kristin Braa, Tone Bratteteig, and Leikny Øgrim. Organizing the redesign process in system development. The Journal of Systems and Software, 33(2): 133–140, May 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Benander:1996:RVI

[BBP96] Alan C. Benander, Barbara A.

Benander, and Howard Pu. Recursion vs. iteration: An empirical study of comprehension. The Journal of Systems and Software, 32(1):73–82, January 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bourque:1991:ESS

[BC91] Pierre Bourque and Vianney Côté. An experiment in software sizing with structured analysis metrics. The Journal of Systems and Software, 15(2):159– 172, May 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

BritoeAbreu:1994:CMO

[BC94] Fernando Brito e Abreu and Rogério Carapuça. Candidate metrics for object-oriented software within a taxonomy framework. The Journal of Systems and Software, 26(1):87–96, July 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Benedusi:1992:REP

[BCD92] P. Benedusi, A. Cimitile, and U. De Carlini. Reverse engineering processes, design document production, and structure charts. The Journal of Systems and Software, 19(3):225–245, November 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bouge:1986:TSG

[BCFG86] L. Bouge, N. Choquet, L. Fribourg, and M.-C. Gaudel. Test sets generation from algebraic specifications using logic programming. The Journal of Systems and Software, 6(4):343–360, November 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barbosa:1994:DAO

[BdADH94] Valmir C. Barbosa, Lúcia Maria de A. Drummond, and Astrid Luise H. Hellmuth. From distributed algorithms to OCCAM programs by successive refinements. The Journal of Systems and Software, 26(3):257– 272, September 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bianchi:1993:GAC

[BDM^{+}93] U. Bianchi, P. Degano, S. Mannucci, S. Martini, B. Mojana, C. Priami, and E. Salvatori. Generating the analytic component parts of syntax-directed editors with efficient error recovery. The Journal of Systems and Software, 23(1):65–79, October 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Belady:1981:SPM

[BE81] L. A. Belady and C. J. Evangelisti. System partitioning and its measure. The Journal of Systems and Software, 2(1):23– 29, February 1981. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Becker:1986:ISI

[Bec86] Monique Becker. Introduction to the special issue. The Journal of Systems and Software, 6 (1–2):5, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Belkhouche:1991:GAP

[Bel91] Boumediene Belkhouche. Generation of ADA and PL/1 prototypes from abstract data type specifications. The Journal of Systems and Software, 16(3): 255–264, November 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Belmesk:1993:EIM

[Bel93] Zoubir Belmesk. An efficient implementation model for the FP2 parallel language. The Journal of Systems and Software, 21(1): 89–101, April 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bernstein:1981:SPM

[Ber81] Lawrence Bernstein. Software project management audits. The Journal of Systems and Software, 2(4):281–287, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bernstein:1988:SS

[Ber88] Lawrence Bernstein. System sizing. The Journal of Systems and Software, 8(3):253, June 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bertolino:1991:OAS

[Ber91] Antonia Bertolino. An overview of automated software testing. The Journal of Systems and Software, 15(2):133–138, May 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bertolino:1993:UET

[Ber93]  Antonia Bertolino.  Uncon-

strained edges and their application to branch analysis and testing of programs. The Journal of Systems and Software, 20(2): 125–133, February 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bertolino:1994:GEC

[Ber94] Antonia Bertolino. Guest Editor’s corner: Achieving quality in software. The Journal of Systems and Software, 26(1): 1–3, July 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Berry:1995:IIR

[Ber95] Daniel M. Berry. The importance of ignorance in requirements engineering. The Journal of Systems and Software, 28(2):179– 184, February 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Berzins:1998:RCS

[Ber98]  V. Berzins.  Recombining

changes to software specifications. The Journal of Systems and Software, 42(2):165– 174, August 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/2/6035.pdf.

Bevan:1999:QUM

[Bev99] Nigel Bevan. Quality in use:

Meeting user needs for quality. The Journal of Systems and Software, 49(1):89–96, December 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

elsevier.nl/gej-ng/10/29/
11/46/25/34/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/25/abstract.
html.

Basili:1981:PME

[BF81] Victor R. Basili and Karl Freburger. Programming measurement and estimation in the software engineering laboratory. The Journal of Systems and Software, 2(1):47–57, February 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bush:1990:SMC

[BF90] Martin E. Bush and Norman E.

Fenton. Software measurement: a conceptual framework. The Journal of Systems and Software, 12(3):223–231, July 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Buehrer:1996:STI

[BF96] Daniel J. Buehrer and Yang-Wen Fan. SL-trees: an indexing structure for object-oriented data bases. The Journal of Systems and Software, 32(3):237– 249, March 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bulgren:1992:MIL

[BFC92] William Bulgren, Victor Frost, and Chi-Shih Chang. A model for interconnected LANs with general arrivals. The Journal of Systems and Software, 17(2): 145–153, February 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bernardeschi:1997:IAJ

[BFG97] C. Bernardeschi, A. Fantechi, and S. Gnesi. An industrial application for the JACK environment. The Journal of Systems and Software, 39(3):249–264, December 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Babaoglu:1996:UFS

[BFR96] Özalp Babaoğlu, Eddy Fromentin, and Michel Raynal. A unified framework for the specification and run-time detection of dynamic properties in distributed computations. The Journal of Systems and Software, 33(3):287–298, June 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Belkhouche:1996:FSP

[BG96] Boumediene Belkhouche and Bart J. Geraci. A formally specified prototyping system. The Journal of Systems and Software, 34(1):67–81, July 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baekgaard:1998:RTE

[BG98] Lars Bækgaard and Jens Christian Godskesen. Real-time event control in active databases. The Journal of Systems and Software, 42(3):263–271, September 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/3/6042.pdf.

Benander:1990:ESU

[BGB90] Barbara A. Benander, Narasimhaiah Gorla, and Alan C. Benander. An empirical study of the use of the GOTO statement. The Journal of Systems and Software, 11(3):217–223, March 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bruegge:1983:GPE

[BH83] Bernd Bruegge and Peter Hibbard. Generalized path expressions: a high-level debugging mechanism. The Journal of Systems and Software, 3(4):265–276, December 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bhargava:1984:PER

[Bha84] Bharat Bhargava. Performance evaluation of reliability control algorithms for distributed database systems. The Journal of Systems and Software, 4(2–3):

239–264, July 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bhide:1990:GSP

[Bhi90] Sandhiprakash Bhide. Generalized software process-integrated metrics framework. The Journal of Systems and Software, 12(3): 249–254, July 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Budgen:1989:AIC

[BHR89] David Budgen, Peter Henderson, and Chic Rattray. Academic/industrial collaboration in a postgraduate MSc course in software engineering. The Journal of Systems and Software, 10(4):261–266, November 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Blaine:1985:CMA

[BK85] J. David Blaine and Richard A. Kemmerer. Complexity measures for assembly language programs. The Journal of Systems and Software, 5(3):229–245, August 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bock:1992:FSF

[BK92] Douglas B. Bock and Robert Klepper. FP-S: a simplified function point counting method. The Journal of Systems and Software, 18(3):245–254, July 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bieman:1995:MLS

[BK95] James M. Bieman and Santhi Karunanithi. Measurement of language-supported reuse in object-oriented and object-based software. The Journal of Systems and Software, 30(3):271– 293, September 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Blair:1985:OIS

[BKS85] Jean R. Stephenson Blair, Phil Kearns, and Mary Lou Soffa. Optimistic implementation of the stack-heap. The Journal of Systems and Software, 5(3):193–202, August 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bellman:1995:DTH

[BL95] Kirstie L. Bellman and Christopher Landauer. Designing testable, heterogeneous software environments. The Journal of Systems and Software, 29(3): 199–217, June 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Binkley:1998:APS

[BL98] David W. Binkley and James R.

Lyle. Application of the pointer state subgraph to static program slicing. The Journal of Systems and Software, 40(1):17–27, January 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Blatt:1987:CNH

[Bla87] David W. E. Blatt. Comparison of network and hierarchical data management in large programming projects. The Journal of Systems and Software, 7(4): 297–309, December 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bickel:1992:ECE

[BLPB92] Riva Wenig Bickel, Maria M. Lorrondo-Petrie, and David F. Bush. EDICT for computer ethics education. The Journal of Systems and Software, 17(1):81– 89, January 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Blum:1986:FYE

[Blu86] Bruce I. Blum. Four years of experience with an environment for implementing information systems. The Journal of Systems and Software, 6(3):261– 271, August 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Blum:1989:CCV

[Blu89] Bruce I. Blum. Controversy corner: Volume, distance and productivity. The Journal of Systems and Software, 10(3):217– 226, October 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Blum:1993:EAD

[Blu93] Bruce I. Blum. The economics of adaptive design. The Journal of Systems and Software, 21(2): 117–128, May 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bussolati:1983:SDD

[BM83] U. Bussolati and G. Martella.

Security design in distributed database systems. The Journal of Systems and Software, 3(3): 219–229, September 1983. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bayer:1989:CDT

[BM89] Judy Bayer and Nancy Melone.

A critique of diffusion theory as a managerial framework for understanding adoption of software engineering innovations. The Journal of Systems and Software, 9 (2):161–166, February 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Borba:1993:VSF

[BM93a] Paulo Borba and Silvio Meira.

From VDM specifications to functional prototypes. The Journal of Systems and Software, 21 (3):267–278, June 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Byoungju:1993:HPM

[BM93b] Choi Byoungju and Aditya P.

Mathur. High-performance mutation testing. The Journal of Systems and Software, 20(2): 135–152, February 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bertolino:1996:HMP

[BM96] Antonia Bertolino and Martina Marré. How many paths are needed for branch testing? The Journal of Systems and Software, 35(2):95–106, November 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Burd:1998:MIR

[BM98] Elizabeth Burd and Malcolm Munro. A method for the identification of reusable units through the reengineering of legacy code. The Journal of Systems and Software, 44(2):121– 134, December 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/2/6077.pdf.

Bertolino:1997:CSB

[BMP97] Antonia Bertolino, Raffaela Mirandola, and Emilia Peciola. A case study in branch testing automation. The Journal of Systems and Software, 38(1):47–59, July 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bettaz:1994:RAM

[BMSB94] Mohamed Bettaz, Mourad Maouche, Moussa Soualmi, and Madani Boukebeche. On reusing ATNet modules in protocol specification. The Journal of Systems and Software, 27(2):119–128, November 1994. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Bowman:1990:SMP

[BN90] Brent J. Bowman and William A.

Newman. Software metrics as a programming training tool. The Journal of Systems and Software, 13(2):139–147, October 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Boehm:1983:SBP

[Boe83] Barry W. Boehm. Seven basic principles of software engineering. The Journal of Systems and Software, 3(1):3–24, March 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bologna:1997:GEC

[Bol97a] Sandro Bologna. Guest Editor’s corner: Achieving quality in software. The Journal of Systems and Software, 38(1):1–2, July 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Boloix:1997:SEQ

[Bol97b] Germinal Boloix. System evaluation and quality improvement. The Journal of Systems and Software, 36(3):297– 311, March 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bowen:1984:MSS

[Bow84] John B. Bowen. Module size:

a standard or heuristic? The Journal of Systems and Software, 4(4):327–332, November 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Basili:1980:TCS

[BP80] Victor R. Basili and John G. Perry, Jr. Transporting up: a case study. The Journal of Systems and Software, 1(2):123–129, February 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Becker:1986:SAN

[BP86] M. Becker and G. Pion. Simulation of an Apollo network in order to evaluate response-time speedup of parallel algorithms. The Journal of Systems and Software, 6(1–2):81–91, May 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bobbie:1991:CPP

[BP91] Patrick O. Bobbie and Mike Papazoglou. Clustering PROLOG programs for distributed computations. The Journal of Systems and Software, 16(3):205– 218, November 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brahmadathan:1990:MLL

[BR90] K. Brahmadathan and K. V. S. Ramarao. On the management of long-living transactions. The Journal of Systems and Software, 11(1):45–52, January 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brackett:1989:BUS

[Bra89] John W. Brackett. The Boston University software engineering graduate program: Continuing education through interactive television. The Journal of Systems and Software, 10(4): 267–269, November 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bradley:1996:ERA

[Bra96] James Bradley. Extended relational algebra for reduction of natural quantifier COOL expressions. The Journal of Systems and Software, 33(1):87–100, April 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Britcher:1992:CCC

[Bri92] Robert N. Britcher. Controversy corner: Cards, couriers, and the race to correctness. The Journal of Systems and Software, 17 (3):281–284, March 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brooks:1981:STP

[Bro81] W. Douglas Brooks. Software technology payoff: Some statistical evidence. The Journal of Systems and Software, 2(1):3– 9, February 1981. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brown:1987:CMC

[Bro87] Bradley J. Brown. Checksum methodology as a configuration management tool. The Journal of Systems and Software, 7(2): 141–143, June 1987. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barzel:1986:PFI

[BS86] Ronen Barzel and David Salesin. Patchwork: a fast interpreter for a restricted dataflow language. The Journal of Systems and Software, 6(3):251–259, August 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bieman:1993:ECA

[BS93] James M. Bieman and Pradip K. Srimani. Editor’s corner: Applying specification, verification, and validation techniques. The Journal of Systems and Software, 21(3):199–200, June 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bieman:1996:GEC

[BS96] James M. Bieman and Pradip K. Srimani. Guest Editors’ corner: Software engineering for distributed computing. The Journal of Systems and Software, 33(3): 205–206, June 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Boloix:1993:SMU

[BST93] Germinal Boloix, Paul G. Sorenson, and J. Paul Tremblay. Software metrics using a metasystem approach to software specification. The Journal of Systems and Software, 20(3):273– 294, March 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Briand:1997:RHR

[BT97] Lionel C. Briand and William M.

Thomas. Response to “Handling Regression Subsets in Software Modeling” by Ronald Gulezian, The Journal of Systems and Software, April 1996. The Journal of Systems and Software, 36 (3):313–314, March 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [Gul96].

Batini:1984:CAL

[BTT84] C. Batini, M. Talamo, and R. Tamassia. Computer aided layout of entity relationship diagrams. The Journal of Systems and Software, 4(2–3):163– 173, July 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Buxton:1990:SEY

[Bux90] J. N. Buxton. Software engineering — 20 years on and 20 years back. The Journal of Systems and Software, 13(3): 153–155, November 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Boehm:1980:SCM

[BW80] Barry W. Boehm and R. W.

Wolverton. Software cost modeling: Some lessons learned. The Journal of Systems and Software, 1(3):195–201, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bates:1983:HLD

[BW83] Peter C. Bates and Jack C. Wileden. High-level debugging of distributed systems: the behavioral abstraction approach. The Journal of Systems and Software, 3(4):255–264, December 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brandl:1993:IOM

[BW93] Dennis L. Brandl and James H. Worley. An implemented object model of the software engineering process. The Journal of Systems and Software, 23(2): 171–181, November 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brown:1995:SFT

[BW95] Larry Brown and Jie Wu. Snooping fault-tolerant distributed shared memories. The Journal of Systems and Software, 29(2): 149–165, May 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Basumallick:1996:DID

[BW96] Swagato Basumallick and Johnny S. K. Wong. Design and implementation of a distributed database system. The Journal of Systems and Software, 34(1):

21–29, July 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bravoco:1985:MMI

[BY85] Ralph R. Bravoco and Surya B.

Yadav. Methodology to model the information structure of an organization. The Journal of Systems and Software, 5(1):59– 71, February 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Berry:1987:APD

[BYY87] Daniel M. Berry, Nancy Yavne, and Moshe Yavne. Application of program design language tools to Abbott’s method of program design by informal natural language descriptions. The Journal of Systems and Software, 7(3): 221–247, September 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Card:1987:CRS

[CA87a] D. N. Card and W. W. Agresti.

Comments on resolving the software science anomaly. The Journal of Systems and Software, 7 (1):83–84, March 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Card:1987:RSS

[CA87b] D. N. Card and W. W. Agresti.

Resolving the software science anomaly. The Journal of Systems and Software, 7(1):29–35, March 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Card:1988:MSD

[CA88] David N. Card and William W. Agresti. Measuring software design complexity. The Journal of Systems and Software, 8(??): 185–197, June 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Card:1989:FRS

[CA89] D. N. Card and W. W. Agresti. Further remarks on software science. The Journal of Systems and Software, 10(1):71–72, July 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Card:1990:MSS

[CA90] David N. Card and William W. Agresti. More on software science. The Journal of Systems and Software, 12(2):170– 171, May 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cai:1998:END

[Cai98] Kai-Yuan Cai. On estimating the number of defects remaining in software. The Journal of Systems and Software, 40(2):93– 114, February 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cam:1999:HPB

[Çam99]  Hasan Çam.  A header-

population based flow control for multicomputer networks. The Journal of Systems and Software, 47(1):19–26, May 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/47/1/6135.pdf.

Cargill:1983:BDb

[Car83] Thomas A. Cargill. Blit debugger. The Journal of Systems and Software, 3(4):277–284, December 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Card:1992:DSP

[Car92] David N. Card. Designing software for producibility. The Journal of Systems and Software, 17 (3):219–225, March 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Carver:1994:IMD

[Car94] Doris L. Carver. Integrated modeling of distributed objectoriented systems. The Journal of Systems and Software, 26(3): 233–244, September 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Carver:1996:TAD

[Car96] Richard H. Carver. Testing abstract distributed programs and their implementations: a constraint-based approach. The Journal of Systems and Software, 33(3):223–237, June 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Carlisle:1999:ECS

[Car99] Judith Pinn Carlisle. Ethical considerations of the softwaredependent organization. The Journal of Systems and Software, 44(3):251–255, January 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/44/3/6089.
pdf.

Cavano:1984:SRM

[Cav84] Joseph P. Cavano. Software reliability measurement: Prediction, estimation, and assessment. The Journal of Systems and Software, 4(4):269–275, November 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Canas:1989:FMS

[CB89a] Daniel A. Cañas and William G. Bulgren. A file management system for a magnetic disk used as a buffer to write-once optical storage. The Journal of Systems and Software, 10(1):15–21, July 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Card:1989:IEA

[CB89b] D. N. Card and R. A. Berg. An industrial engineering approach to software development. The Journal of Systems and Software, 10(3):159–168, October 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Coleman:1991:DKP

[CB91] David L. Coleman and Albert L.

Baker. Deliberations on Kung’s process interface modeling. The Journal of Systems and Software, 15(2):193–198, May 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See response [Kun91b].

Chang:1996:IZS

[CBK96] C. C. Chang, D. J. Buehrer, and H. C. Kowng. An improvement to Ziegler’s sparse matrix compression algorithm. The Journal of Systems and Software, 35 (1):67–71, October 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cote:1988:SMO

[CBOR88] V. Cote, P. Bourque, S. Oligny, and N. Rivard. Software metrics: an overview of recent results. The Journal of Systems and Software, 8(2):121–131, March 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:1994:NOO

[CC94] Jen-Yen Chen and Yao-Ting Tim Chen. A new object-oriented method integrating Jackson structured programming method. The Journal of Systems and Software, 24(2):125–137, February 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:1999:CMP

[CC99a] J.-Y. J. Chen and S.-C. Chou.

Consistency management in a process environment. The Journal of Systems and Software, 47 (2–3):105–110, July 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/47/2-3/6145.pdf.

Chiang:1999:CPM

[CC99b] M.-L. Chiang and R.-C. Chang. Cleaning policies in mobile computers using flash memory. The Journal of Systems and Software, 48(3):213–231, November 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

elsevier.nl/gej-ng/10/29/
11/37/24/22/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/24/abstract.
html.

Cimitile:1999:IOL

[CDDF99] Aniello Cimitile, Andrea De Lucia, Giuseppe Antonio Di Lucca, and Anna Rita Fasolino. Identifying objects in legacy systems using design metrics. The Journal of Systems and Software, 44(3):199– 211, January 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/44/3/6085.pdf.

Cuykendall:1984:DSV

[CDJ^{+}84] Robert Cuykendall, Anton Domic, William H. Joyner, Steve C. Johnson, Steve Kelem, Dennis McBride, Jack Mostow, John E. Savage, and Gabriele Saucier. Design synthesis in VLSI and software engineering. The Journal of Systems and Software, 4(1):7–12, April 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Canfora:1998:IER

[CDM98] Gerardo Canfora, Andrea De Lucia, and Malcolm Munro. An integrated environment for reuse reengineering C code. The Journal of Systems and Software, 42(2):153–164, August 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/2/6034.pdf.

Cao:1999:RPD

[CDS99] Jiannong Cao, Olivier De Vel, and Ling Shi. Rapid prototyping of distributed algorithms. The Journal of Systems and Software, 45(2):141– 154, March 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/2/6101.pdf.

Carrasco:1991:ESO

[CFK91] Juan A. Carrasco, Joan Figueras, and Annie Kuntzmann. Evaluation of safety-oriented twoversion architectures. The Journal of Systems and Software, 14 (3):155–162, March 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Craig:1998:CGW

[CFSS98] Annemieke Craig, Julie Fisher, Angela Scollary, and Mohini Singh. Closing the gap: Women, education, and information technology courses in Australia. The Journal of Systems and Software, 40(1):7–15, January 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:1994:SOD

[CG94] Yinong Chen and Winfried Görke. Structure-oriented dependability evaluation of faulttolerant protocols. The Journal of Systems and Software, 25(1): 85–99, April 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cooke:1996:LSS

[CGD^{+}96] Daniel Cooke, Ann Gates, Elif Demirors, Onur Demirors, Murat M. Tanik, and Bernd Kramer. Languages for the specification of software. The Journal of Systems and Software, 32 (3):269–308, March 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Claybrook:1983:LES

[CH83] Billy G. Claybrook and H. Rex Hartson. Language extensions for specifying program access control policies in programming languages. The Journal of Systems and Software, 3(2):155–162, June 1983. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:1994:IOO

[CH94] Jen-Yen Chen and Yu-Shiang Hung. An integrated objectoriented analysis and design method emphasizing entity/class relationship and operation finding. The Journal of Systems and Software, 24(1):31–47, January 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1991:DCU

[Cha91] Daniel K. Chang. Data compression using hierarchical dictionaries. The Journal of Systems and Software, 15(3):233– 238, July 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1993:SPM

[Cha93] Daniel K. Chang. A string pattern—matching algorithm. The Journal of Systems and Software, 22(3):207–216, September 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chatman:1995:CPP

[Cha95] Vernon V. Chatman, III. CHANGE-POINTs: a proposal for software productivity measurement. The Journal of Systems and Software, 31(1):71–91, October 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1994:NCM

[CHB94] Chin-Chen Chang, Shin-Jia Hwang, and Daniel J. Buehrer. New concepts for multiple stack system: Born and reborn. The Journal of Systems and Software, 27(1):45–54, October 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chow:1995:RAM

[Cho95] Ching-Hua Edward Chow. Resource allocation for multiparty connections. The Journal of Systems and Software, 28(3):253– 266, March 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chretienne:1986:TPN

[Chr86] Philippe Chretienne. Timed Petri nets: a solution to the minimum-time-reachability problem between two states of a timed-event graph. The Journal of Systems and Software, 6(1–2): 95–101, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Christodoulakis:1991:GSE

[Chr91] Dimitris Christodoulakis. The GRASPIN software engineering environment. The Journal of Systems and Software, 15(1):3, April 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Christie:1999:SSC

[Chr99] Alan M. Christie. Simulation in support of CMM-based process improvement. The Journal of Systems and Software, 46(2–3): 107–112, April 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/2-3/6119.pdf.

Chu:1997:ASO

[Chu97] Pai-Cheng Chu. Actors and scripts: Object-oriented simulation of enterprise systems. The Journal of Systems and Software, 37(3):187–199, June 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cioch:1991:MSM

[Cio91] Frank A. Cioch. Measuring software misinterpretation. The Journal of Systems and Software, 14(2):85–95, February 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chee:1998:MWB

[CJP98] C. L. Chee, S. Jarzabek, and R. Paul. F -metric: a WWWbased framework for intelligent formulation and analysis of metric queries. The Journal of Systems and Software, 43(2):119– 132, November 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/2/6055.pdf.

Chiu:1998:ALB

[CKyL98] Anthony Chiu, Ben Kao, and Kam yiu Lam. An analysis of lock-based and optimistic concurrency control protocols in multiprocessor real-time databases. The Journal of Systems and Software, 42(3):273– 286, September 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/42/3/6043.pdf.

Chandersekaran:1981:SSU

[CL81] C. S. Chandersekaran and R. C. Linger. Software specification using the special language. The Journal of Systems and Software, 2(1):31–38, February 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1994:IDF

[CL94] Yu Lo Cyrus Chang and Leslie C. Lander. An inference design for fault location in real-time control systems. The Journal of Systems and Software, 25(1):3– 21, April 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Costello:1995:MRE

[CL95] Rita J. Costello and Dar-Biau Liu. Metrics for requirements engineering. The Journal of Systems and Software, 29(1):39–63, April 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chung:1997:EZO

[CL97] Chyan-Goei Chung and Jen-Gaw Lee. An enhanced zero-one optimal path set selection method. The Journal of Systems and Software, 39(2):145–164, November 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1998:SMR

[CL98] Chin-Chen Chang and Chin-Feng Lee. A spatial match retrieval mechanism for symbolic pictures. The Journal of Systems and Software, 44(1):73– 83, December 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/1/6072.pdf.

Cheong:1999:QSM

[CL99] F. Cheong and R. Lai. QoS specification and mapping for distributed multimedia systems: a survey of issues. The Journal of Systems and Software, 45(2): 127–139, March 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/2/6100.pdf.

Claude:1986:DTQ

[Cla86] J. P. Claude. Discrete time queues for modelling an HDLC coupler. The Journal of Systems and Software, 6(1–2):55–59, May 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1999:PSF

[CLL99] Chin-Chen Chang, Tian-Fu Lee, and Jyh-Jong Leu. Partition search filter and its performance analysis. The Journal of Systems and Software, 47 (1):35–43, May 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1999/47/1/6137.pdf.

Chen:1996:RMF

[CLLC96] C. Y. Chen, H. F. Lin, R. C. T. Lee, and C. C. Chang. Redundant MKH files design among multiple disks for concurrent partial match retrieval. The Journal of Systems and Software, 35(3): 199–207, December 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Coleman:1995:ASM

[CLO95] Don Coleman, Bruce Lowther, and Paul Oman. The application of software maintainability models in industrial software systems. The Journal of Systems and Software, 29(1):3–16, April 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1998:TOO

[CLSC98] Kai H. Chang, Shih-Sung Liao, Stephen B. Seidman, and Richard Chapman. Testing object-oriented programs: From formal specification to test scenario generation. The Journal of Systems and Software, 42(2): 141–151, August 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/2/6033.pdf.

Chauvet:1986:MCX

[CM86] Yves Chauvet and Gerard Meurant. Multitasking on the Cray X-MP. The Journal of Systems and Software, 6(1–2):17–20, May 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Collins:1992:PEC

[CM92] W. Robert Collins and Keith W.

Miller. Paramedic ethics for computer professionals. The Journal of Systems and Software, 17 (1):23–38, January 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chiang:1993:CUF

[CM93] Chi-Ming Chiang and Matt W.

Mutka. Characteristics of user file-usage patterns. The Journal of Systems and Software, 23(3): 257–268, December 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Crawford:1985:ASM

[CMP85] Stewart G. Crawford, Allen A.

McIntosh, and Daryl Pregibon. Analysis of static metrics and faults in C software. The Journal of Systems and Software, 5 (1):37–48, February 1985. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cohen:1981:APL

[Coh81] Paul M. Cohen. Ada programming language standardization. The Journal of Systems and Software, 2(4):351–355, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Collste:1992:ESM

[Col92] Göran Collste. Expert systems in medicine and moral responsibility. The Journal of Systems and Software, 17(1):15–22, January 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cooper:1981:MWS

[Coo81]  Jack Cooper.  Mil-Std-1679,

weapon system software development. The Journal of Systems and Software, 2(4):319–327, December 1981. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cooke:1990:FSR

[Coo90] D. E. Cooke. Formal specifications of resource-deadlock prone Petri nets. The Journal of Systems and Software, 11(1):53–69, January 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Christodoulakis:1988:WWE

[CP88] D. Christodoulakis and T. Pantziou. WEP — a workstation for education in programming. The Journal of Systems and Software, 8(5):401–408, December 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Coppola:1997:PIT

[CP97] P. Coppola and P. Panaroni.

Process improvement through process modelling. The Journal of Systems and Software, 38(1): 95–101, July 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:1998:OOI

[CPW98] Hock Chuan Chan, Danny C. C.

Poo, and Cheng Peng Woon. An object-oriented implementation of an entity relationship model. The Journal of Systems and Software, 41(2): 117–125, May 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/2/6012.pdf.

Clarke:1985:ASE

[CR85] Lori A. Clarke and Debra J.

Richardson. Applications of symbolic evaluation. The Journal of Systems and Software, 5(1):15– 35, February 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Comer:1989:SEE

[CR89] James R. Comer and David J.

Rodjak. Software engineering education at Texas Christian University: Adapting a curriculum to changing needs. The Journal of Systems and Software, 10(4): 235–244, November 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Coupal:1990:FAS

[CR90] Daniel Coupal and Pierre N. Robillard. Factor analysis of source code metrics. The Journal of Systems and Software, 12(3): 263–269, July 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cook:1994:RTS

[CRV94] Curtis R. Cook, Andreas Roesch, and C. W. Vowell. Real-time software metrics. The Journal of Systems and Software, 24 (3):223–237, March 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Carver:1985:IPM

[CS85] Doris L. Carver and D. B. Simmons. Impact of programming methodology on program complexity. The Journal of Systems and Software, 5(4):279–289, November 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Curtis:1989:EES

[CSKB^{+}89] Bill Curtis, Sylvia B. Sheppard, Elizabeth Kruesi-Bailey, John Bailey, and Deborah A. Boehm-Davis. Experimental evaluation of software documentation formats. The Journal of Systems and Software, 9(2):167–207, February 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:1994:ALS

[CT94] Jen-Yen Chen and Chia-Ming Tu. An Ada-like software process language. The Journal of Systems and Software, 27(1):17– 25, October 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1997:GSS

[CT97] Chin-Chen Chang and Hui-Min Tsai. A generalized secret sharing scheme. The Journal of Systems and Software, 36(3):267– 272, March 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chou:1994:GCS

[CTA94] Yu-Hur Chou, Cheng-Hsien Tung, and Reda Ammar. Guidelines for choosing self-scheduling orders on multiprocessors. The Journal of Systems and Software, 26(3):245–255, September 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cardenas:1992:ADT

[CTZ92] Sergio R. Cárdenas, Jianhui Tian, and Marvin V. Zelkowitz. An application of decision theory for the evaluation of software prototypes. The Journal of Systems and Software, 19(1):27–39, September 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cooke:1998:GEI

[CU98] D. E. Cooke and J. E. Urban. Guest Editors’ introduction: Special issue on software engineering and knowledge engineering. The Journal of Systems and Software, 42(2):103– 104, August 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cimitile:1995:SSC

[CV95] A. Cimitile and G. Visaggio. Software salvaging and the call dominance tree. The Journal of Systems and Software, 28(2): 117–127, February 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Collofello:1989:EER

[CW89] James S. Collofello and Scott N. Woodfield. Evaluating the effectiveness of reliability-assurance techniques. The Journal of Systems and Software, 9(3):191–195, March 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Compton:1990:PCA

[CW90] B. Terry Compton and Carol Withrow. Prediction and control of ADA software defects. The Journal of Systems and Software, 12(3):199–207, July 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1997:LAD

[CW97] Chin-Chen Chang and Chih-Hung Wang. A locally adaptive data compression strategy for Chinese–English characters. The Journal of Systems and Software, 36(2):167–179, February 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chong:1991:PES

[CZ91] Edwin K. P. Chong and Wei Zhao. Performance evaluation of scheduling algorithms for imprecise computer systems. The Journal of Systems and Software, 15(3):261–277, July 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cao:1998:HOC

[CZdV98] Jiannong Cao, Kang Zhang, and Olivier de Vel. On heuristics for optimal configuration of hierarchical distributed monitoring systems. The Journal of Systems and Software, 43(3):197– 206, November 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/3/6061.pdf.

Cetintemel:1999:OBO

[ÇZUB99] Ugũr Çetintemel, Jürgen Zimmermann, Özgür Ulusoy, and Alejandro Buchmann. Objective: a benchmark for objectoriented active database systems. The Journal of Systems and Software, 45(1):31– 43, February 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1999/45/1/6093.pdf.

Davis:1986:PAS

[DA86] Margaret J. Davis and David R. Addleman. A practical approach to specification technology selection. The Journal of Systems and Software, 6(3):285–294, August 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Damiani:1996:ISC

[Dam96] Maria Luisa Damiani. An intelligent system for complex data exploration. The Journal of Systems and Software, 33(2):123– 132, May 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Davis:1988:TES

[Dav88] Alan M. Davis. A taxonomy for the early stages of the software development life cycle. The Journal of Systems and Software, 8(4):297–311, September 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Davis:1995:OOR

[Dav95] Alan M. Davis. Object-oriented requirements to object-oriented design: an easy transition? The Journal of Systems and Software, 30(1–2):151–159, July/ August 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Davidson:1999:JAD

[Dav99] E. J. Davidson. Joint application design (JAD) in practice. The Journal of Systems and Software, 45(3):215– 223, March 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/3/6107.pdf.

Doe:1986:SPC

[DB86] Dennis D. Doe and Edward H.

Bersoff. Software Productivity Consortium (Spc): an industry initiative to improve the productivity and quality of missioncritical software. The Journal of Systems and Software, 6(4): 367–378, November 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Derrick:1995:VSS

[DB95] E. Joseph Derrick and Osman Balci. A visual simulation support environment based on the Domino conceptual framework. The Journal of Systems and Software, 31(3):215–237, December 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DeBosschere:1997:PBP

[De 97] Koen De Bosschere. Processbased parallel logic programming: a survey of the basic issues. The Journal of Systems and Software, 39(1):71–82, October 1997. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

DeBosschere:1998:TEF

[De 98] Koen De Bosschere. Tarilan: An embedded functional data processing language. The Journal of Systems and Software, 43(2): 93–102, November 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/43/2/6053.pdf.

Dehnad:1990:SMU

[Deh90] Khosrow Dehnad. Software metrics from a user’s perspective. The Journal of Systems and Software, 13(2):111–115, October 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Delugach:1992:SMV

[Del92]  Harry S. Delugach.  Speci-

fying multiple-viewed software requirements with conceptual graphs. The Journal of Systems and Software, 19(3):207– 224, November 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dorfman:1984:AAR

[DF84] Merlin Dorfman and Richard F. Flynn. ARTS — an automated requirements traceability system. The Journal of Systems and Software, 4(1):63–74, April 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DiFelice:1998:HWC

[DF98] Paolino Di Felice and Goffredo Fonzi. How to write comments suitable for automatic software indexing. The Journal of Systems and Software, 42 (1):17–28, July 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/1/6024.pdf.

Drury:1999:ITP

[DF99] D. H. Drury and A. Farhoomand.

Information technology push/ pull reactions. The Journal of Systems and Software, 47 (1):3–10, May 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/47/1/6133.pdf.

Dietrich:1996:AFT

[DFCR96] Suzanne Wagner Dietrich, Changguan Fan, and Enrique Cortes-Rello. An application of fragmentation transparency in a distributed database system: a case study. The Journal of Systems and Software, 35(3):185–197, December 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dunsmore:1980:AEP

[DG80] H. E. Dunsmore and J. D. Gannon. Analysis of the effects of programming factors on programming effort. The Journal of Systems and Software, 1(2):141– 153, February 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Davis:1987:RCV

[DG87] Frank G. F. Davis and Rex E. Gantenbein. Recovering from a computer virus attack. The Journal of Systems and Software, 7(4):253–258, December 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Davis:1992:RCE

[DG92] Frank G. F. Davis and Rex E. Gantenbein. Responding to catastrophic errors: a design technique for fault-tolerant software. The Journal of Systems and Software, 17(3):243– 251, March 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DePaoli:1998:RMF

[DG98] Damien De Paoli and Andrzej Goscinski. The RHODOS migration facility. The Journal of Systems and Software, 40(1):51– 65, January 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dauchy:1993:UAS

[DGM93] P. Dauchy, M.-C. Gaudel, and B. Marre. Using algebraic specifications in software testing: a case study on the software of an automatic subway. The Journal of Systems and Software, 21(3):

229–244, June 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Desai:1988:CID

[DGS88] B. C. Desai, P. Goyal, and F. Sadri. Composite index in Ddbms. The Journal of Systems and Software, 8(2):105–119, March 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dhama:1995:QMC

[Dha95] Harpal Dhama. Quantitative models of cohesion and coupling in software. The Journal of Systems and Software, 29(1):65–74, April 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dupuis:1986:CTM

[DHP86] Alain Dupuis, Gerard Hebuterne, and Jean-Marc Pitie. Comparison of two mutual-exclusion algorithms for computer networks. The Journal of Systems and Software, 6(1–2):137–145, May 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DiFelice:1987:DRL

[Di 87] Paolino Di Felice. On the development of reliable large software. The Journal of Systems and Software, 7(2):99–107, June 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dillon:1991:IAS

[Dil91] Laura K. Dillon. An isolation approach to symbolic executionbased verification of Ada tasking programs. The Journal of Systems and Software, 14(3):183– 198, March 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Diaz:1998:PBP

[DIP98] Oscar Diaz, Jon Iturrioz, and Mario G. Piattini. Promoting business policies in objectoriented methods. The Journal of Systems and Software, 41 (2):105–115, May 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/2/6011.pdf.

Deng-Jyi:1993:SSR

[DJL93] Chen Deng-Jyi and P. J. Lee. On the study of software reuse using reusable C++ components. The Journal of Systems and Software, 20(1):19–36, January 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Davis:1994:DAS

[DK94] J. Steve Davis and J. J. Kanet. Developing an application-specific interactive system based on a spreadsheet package. The Journal of Systems and Software, 25 (2):201–206, May 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Davis:1997:PSI

[DK97] J. Steve Davis and J. J. Kanet.

Production scheduling: An interactive graphical approach. The Journal of Systems and Software, 38(2):155–163, August 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Drappa:1999:QMI

[DL99] A. Drappa and J. Ludewig.

Quantitative modeling for the interactive simulation of software projects. The Journal of Systems and Software, 46(2–3): 113–122, April 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/2-3/6120.pdf.

Doukidis:1996:ISP

[DLG96] Georgios I. Doukidis, Panagiotis Lybereas, and Robert D. Galliers. Information systems planning in small business: a stages of growth analysis. The Journal of Systems and Software, 33(2): 189–201, May 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Debou:1994:DMS

[DLS94] Christophe Debou, Jaroslav Lipták, and Herbert Schippers. Decision making for software process improvement: a quantitative approach. The Journal of Systems and Software, 26(1): 43–52, July 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DiStefano:1999:FAD

[DLT99] A. Di Stefano, L. Lo Bello, and E. Tramontana. Factors affecting the design of load balancing algorithms in distributed systems. The Journal of Systems and Software, 48(2):105–117, October 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

elsevier.nl/gej-ng/10/29/
11/37/23/20/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/23/abstract.
html.

Datta:1998:BMR

[DMV98] Anindya Datta, Sarit Mukherjee, and Igor R. Viguier. Buffer management in real-time active database systems. The Journal of Systems and Software, 42(3):227–246, September 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/3/6040.pdf.

Dolado:1997:SRA

[Dol97] J. J. Dolado. A study of the relationships among Albrecht and Mark II function points, lines of code 4GL and effort. The Journal of Systems and Software, 37(2):161–173, May 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DCunha:1984:DDA

[DR84] Adolph D’Cunha and T. Radhakrishnan. Dass: a data administration support system. The Journal of Systems and Software, 4(2–3):175–184, July 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Delen:1992:SEM

[DR92] G. P. A. J. Delen and D. B. B. Rijsenbrij. The specification, engineering, and measurement of information systems quality. The Journal of Systems and Software, 17(3):205–217, March 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dooley:1985:FSD

[DS85] Jeffrey W. M. Dooley and Stephen R. Schach. Flow: a software development environment using diagrams. The Journal of Systems and Software, 5(3):203– 219, August 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Damerla:1992:SCA

[DS92] Srinivasarao Damerla and Sol M.

Shatz. Software complexity and Ada rendezvous: Metrics based on nondeterminism. The Journal of Systems and Software, 17 (2):119–127, February 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DeMent:1994:NAM

[DS94] Naomi S. DeMent and Pradip K.

Srimani. A new algorithm for k mutual exclusions in distributed systems. The Journal of Systems and Software, 26(2):179– 191, August 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dishaw:1998:SSM

[DS98] Mark T. Dishaw and Diane M. Strong. Supporting software maintenance with software engineering tools: a Computed task–technology fit analysis. The Journal of Systems and Software, 44(2):107–120, December 1, 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/2/6076.pdf.

Davis:1990:LCM

[DT90] I. J. Davis and D. J. Taylor. Local correction of mod(k ) lists. The Journal of Systems and Software, 11(3):205– 214, March 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Duvall:1995:SSM

[Duv95] Lorraine M. Duvall. A study of software management: The state of the practice in the United States and Japan. The Journal of Systems and Software, 31(2): 109–124, November 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dugan:1994:REF

[DV94] Joanne Bechta Dugan and Randy Van Buren. Reliability evaluation of fly-by-wire computer systems. The Journal of Systems and Software, 25 (1):109–120, April 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Deng:1999:ADM

[DY99] Yi Deng and Chia-Rung Yang.

Architecture-driven modeling of real-time concurrent systems with applications in FMS. The Journal of Systems and Software, 45(1):61–78, February 1, 1999. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/1/6095.pdf.

Dyer:1987:FAS

[Dye87] Michael Dyer. Formal approach to software error removal. The Journal of Systems and Software, 7(2):109–114, June 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dyer:1993:DBS

[Dye93] Michael Dyer. Distributionbased statistical sampling: An approach to software functional test. The Journal of Systems and Software, 20(2):107–114, February 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ebert:1994:CR

[Ebe94] Christof Ebert. Controversy revisited. The Journal of Systems and Software, 27(2):154, November 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ebert:1999:TCS

[Ebe99] Christof Ebert. Technical controlling and software process improvement. The Journal of Systems and Software, 46(1): 25–39, April 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1999/46/1/6113.pdf.

Easterbrook:1998:FMV

[EC98] Steve Easterbrook and John Callahan. Formal methods for verification and validation of partial specifications: a case study. The Journal of Systems and Software, 40(3):199– 210, March 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Edwards:1993:AOO

[EHS93] J. M. Edwards and B. Henderson-Sellers. Application of an objectoriented analysis and design methodology to engineering cost management. The Journal of Systems and Software, 23(2): 123–138, November 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Eckhardt:1988:FDR

[EL88] Dave E. Eckhardt and Larry D.

Lee. Fundamental differences in the reliability of N-modular redundancy and N-version programming. The Journal of Systems and Software, 8(4):313–318, September 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Evanco:1994:MBF

[EL94] William M. Evanco and Robert Lacovara. A model-based framework for the integration of software metrics. The Journal of Systems and Software, 26(1):77– 86, July 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Eliot:1992:CAE

[Eli92] Lance B. Eliot. Case analysis of expert systems projects: Strategies and examples. The Journal of Systems and Software, 19 (2):153–157, October 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Emdad:1991:EIE

[Emd91] Ali Emdad. An empirical investigation of the effectiveness of software documentation delivery systems. The Journal of Systems and Software, 16(2):129– 134, October 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

England:1981:ICS

[Eng81] Gordon R. England. Implementation of computing standards. The Journal of Systems and Software, 2(4):371–377, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Eassa:1995:ADA

[EOM95] F. E. Eassa, L. J. Osterweil, and M. Z. Abdel Mageed. AIDA: a dynamic analyzer for Ada programs. The Journal of Systems and Software, 31(3):239–255, December 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Eriksson:1992:SKA

[Eri92] Henrik Eriksson. A survey of knowledge acquisition techniques and tools and their relationship to software engineering. The Journal of Systems and Software, 19(1):97–107, September 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ellison:1985:EGS

[ES85] Robert J. Ellison and Barbara J. Staudt. Evolution of the Gandalf system. The Journal of Systems and Software, 5(2):107–119, May 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Elboushi:1997:OOS

[ES97] Mekki I. Elboushi and Joseph S. Sherif. Object-oriented software design utilizing Quality Function Deployment. The Journal of Systems and Software, 38(2):133– 143, August 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Eskenasi:1989:ESP

[Esk89] Avram Eskenasi. Evaluation of software product quality by means of classification methods. The Journal of Systems and Software, 10(3):213–216, October 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Evangelist:1983:SCM

[Eva83] W. M. Evangelist. Software complexity metric sensitivity to program structuring rules. The Journal of Systems and Software, 3(3):231–243, September 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Evanco:1995:MEC

[Eva95] William M. Evanco. Modeling the effort to correct faults. The Journal of Systems and Software, 29(1):75–84, April 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Evanco:1997:PAD

[Eva97] William M. Evanco. Poisson analyses of defects for small software components. The Journal of Systems and Software, 38(1): 27–35, July 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fairley:1983:EIa

[Fai83a] Richard E. Fairley. Editor’s introduction. The Journal of Systems and Software, 3(2):105, June 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fairley:1983:EIb

[Fai83b] Richard E. Fairley. Editor’s introduction. The Journal of Systems and Software, 3(3):185, September 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fairley:1983:EIc

[Fai83c] Richard E. Fairley. Editor’s introduction. The Journal of Systems and Software, 3(4):253–254, December 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fairley:1984:EI

[Fai84] Richard E. Fairley. Editor’s introduction. The Journal of Systems and Software, 4(1):1, April 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fairfield:1985:SST

[Fai85a] P. Fairfield. STRAPS — a software transport system for lowlevel software. The Journal of Systems and Software, 5(4): 291–302, November 1985. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fairley:1985:EI

[Fai85b] Richard E. Fairley. Editor’s introduction. The Journal of Systems and Software, 5(1):1, February 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Furuyama:1994:FGM

[FAI94] Tsuneo Furuyama, Yoshio Arai, and Kazuhiko Iio. Fault generation model and mental stress effect analysis. The Journal of Systems and Software, 26(1):31–42, July 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Furuyama:1997:AFG

[FAI97] Tsuneo Furuyama, Yoshio Arai, and Kazuhiko Iio. Analysis of fault generation caused by stress during software development. The Journal of Systems and Software, 38(1):13–25, July 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fritzson:1994:UAD

[FAS94] Peter Fritzson, Mikhail Auguston, and Nahid Shahmehri. Using assertions in declarative and operational models for automated debugging. The Journal of Systems and Software, 25(3): 223–239, June 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Feng:1996:UAW

[FC96] Li-Chi Feng and Ruei-Chuan Chang. Using asynchronous writes on metadata to improve file system performance. The Journal of Systems and Software, 35(1):43–54, October 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fenton:1993:HES

[Fen93] Norman Fenton. How effective are software engineering methods? The Journal of Systems and Software, 22(2):141– 146, August 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ferchichi:1993:HCL

[Fer93] Ahmed Ferchichi. Heuristics for constructing logic programs. The Journal of Systems and Software, 23(3):225–233, December 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Friedman:1987:MMS

[FF87] Hershey H. Friedman and Linda Weiser Friedman. Marketing methods for software. The Journal of Systems and Software, 7(3):207–212, September 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Friedman:1989:MUP

[FF89] Hershey H. Friedman and Linda Weiser Friedman. Myths, unethical practices, personnel requirements: What do computer industry professionals really believe? The Journal of Systems and Software, 10(2): 151–153, September 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frakes:1995:MRA

[FF95] William B. Frakes and Christopher J. Fox. Modeling reuse across the software life cycle. The Journal of Systems and Software, 30(3):295–301, September 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Finney:1996:EEZ

[FF96] Kate Finney and Norman Fenton. Evaluating the effectiveness of Z: The claims made about CICS and where we go from here. The Journal of Systems and Software, 35(3):209–216, December 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fernandez:1993:DLD

[FG93] Mariano G. Fernandez and Sumit Ghosh. Ddbx-LPP: a dynamic software tool for debugging asynchronous distributed algorithms on loosely coupled parallel processors. The Journal of Systems and Software, 22(1): 27–43, July 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fuggetta:1994:SAO

[FG94] Alfonso Fuggetta and Carlo Ghezzi. State of the art and open issues in process-centered software engineering environments. The Journal of Systems and Software, 26(1):53–60, July 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fickas:1989:DIR

[Fic89] Stephen Fickas. Design issues in a rule-based system. The Journal of Systems and Software, 10(2): 113–123, September 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fisher:1981:SQS

[Fis81] Matthew J. Fisher. Software QA standards — the coming revolution. The Journal of Systems and Software, 2(4):357–362, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fisher:1991:IAI

[Fis91] A. J. Fisher. Incremental algorithms for interactive text formatting. The Journal of Systems and Software, 16(1):3–16, September 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Feijs:1998:HPA

[FJ98] L. M. G. Feijs and H. B. M. Jonkers. History, principles and application of the SPRINT method. The Journal of Systems and Software, 41(3):199– 219, June 1, 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/41/3/6020.pdf.

Friedman:1992:HAR

[FK92] Batya Friedman and Peter H. Kahn, Jr. Human agency and responsible computing: Implications for computer system design. The Journal of Systems and Software, 17(1):7–14, January 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fletcher:1995:RCR

[Fle95] Peter Fletcher. Readers’ corner:

The role of experiments in computer science. The Journal of Systems and Software, 30(1–2): 161–163, July/August 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frakes:1991:EET

[FLN91] W. B. Frakes, D. J. Lubinsky, and D. N. Neal. Experimental evaluation of a test coverage analyzer for C and C++. The Journal of Systems and Software, 16 (2):135–139, October 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fugini:1987:CMA

[FM87] Mariagrazia Fugini and Giancarlo Martella. Conceptual modeling of authorization in database systems. The Journal of Systems and Software, 7(1): 3–13, March 1987. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fenton:1990:DSB

[FM90a] Norman Fenton and Austin Melton. Deriving structurally based software measures. The Journal of Systems and Software, 12(3):177–187, July 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Finance:1990:FSI

[FM90b] Jean Pierre Finance and Ali Mili. A foreword to the Special Issue of the Journal of Systems and Software on the Role of Languages in Programming (the Tenth Tunisian-French Seminar of Computer Science). The Journal of Systems and Software, 12(2):83–85, May 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fenelon:1993:ITS

[FM93] Peter Fenelon and John A. Mc-Dermid. An integrated tool set for software safety analysis. The Journal of Systems and Software, 21(3):279–290, June 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fdida:1986:QSR

[FMP86] S. Fdida, D. Mailles, and G. Pujolle. Queueing systems with resource sharing. The Journal of Systems and Software, 6(1–2): 23–29, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Foulk:1985:APN

[FN85] Patrick W. Foulk and Salwa M. Nassar. Analysis of parallelism in nested do loops. The Journal of Systems and Software, 5 (1):73–80, February 1985. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Florin:1986:OPU

[FN86] G. Florin and S. Natkin. Oneplace unbounded stochastic Petri nets: Ergodic criteria and steady-state solutions. The Journal of Systems and Software, 6 (1–2):103–115, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fenton:1999:SMS

[FN99] Norman E. Fenton and Martin Neil. Software metrics: successes, failures and new directions. The Journal of Systems and Software, 47(2–3):149–157, July 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/47/2-3/
6150.pdf.

Fruth:1996:PBO

[FPW96] Lisa Spicknall Fruth, James M.

Purtilo, and Elizabeth L. White. A pattern-based object-linking mechanism for component-based software development environments. The Journal of Systems and Software, 32(3):227– 235, March 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fraisse:1986:EPA

[Fra86] P. Fraisse. Efficient parallel algorithm for the transitive closure of a graph. The Journal of Systems and Software, 6(1–2):165– 168, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frankel:1990:HKB

[Fra90] Ric Frankel. A hybrid knowledgebased approach to structured selection. The Journal of Systems and Software, 13(1):13–23, September 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Finney:1998:MCS

[FRF98] Kate Finney, Keith Rennolls, and Alex Fedorec. Measuring the comprehensibility of Z specifications. The Journal of Systems and Software, 42 (1):3–15, July 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/42/1/6023.pdf.

Fritzson:1983:SDT

[Fri83] Peter Fritzson. Symbolic debugging through incremental compilation in an integrated environment. The Journal of Systems and Software, 3(4):285–294, December 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frieder:1990:FTH

[Fri90] Ophir Frieder. Fault tolerance on a hypercube: a database application. The Journal of Systems and Software, 13(1):25–38, September 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Farr:1988:TSM

[FS88] William H. Farr and Oliver D.

Smith. Tool for statistical modeling and estimation of reliability functions for software: Smerfs. The Journal of Systems and Software, 8(1):47–55, January 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frieder:1991:DUC

[FS91] Ophir Frieder and Mark E. Segal. On dynamically updating a computer program: From concept to prototype. The Journal of Systems and Software, 14(2): 111–128, February 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Felix-Simpson:1987:IAM

[FSA87] Susan C. Felix-Simpson and H. Pat Artis. Implementing an analytic model for Customer Information Control Systems (CICS). The Journal of Systems and Software, 7(1):51–60, March 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fiutem:1999:PAP

[FTAM99] R. Fiutem, P. Tonella, G. Antoniol, and E. Merlo. Pointsto analysis for program understanding. The Journal of Systems and Software, 44(3):213– 227, January 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.

                                        com/cas/tree/store/jss/sub/
                                        1999/44/3/6086.pdf.

Fuggetta:1999:RMS

[Fug99] Alfonso Fuggetta. Rethinking the modes of software engineering research. The Journal of Systems and Software, 47(2–3): 133–138, July 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1999/47/2-3/6148.pdf.

Furbach:1993:FSM

[Fur93] Ulrich Furbach. Formal specification methods for reactive systems. The Journal of Systems and Software, 21(2):129– 139, May 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ford:1990:ICB

[FW90] Ray Ford and Mary Pfreundschuh Wagner. Incremental concurrent builds for modular systems ^{,}. The Journal of Systems and Software, 13(3):157– 176, November 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Finnie:1997:CSE

[FWD97] G. R. Finnie, G. E. Wittig, and J.-M. Desharnais. A comparison of software effort estimation techniques: Using function points with neural networks, case-based reasoning and regression models. The Journal of Systems and Software, 39(3):

281–289, December 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frankl:1997:AUV

[FWH97] Phyllis G. Frankl, Stewart N.

Weiss, and Cang Hu. All-uses vs mutation testing: an experimental comparison of effectiveness. The Journal of Systems and Software, 38(3):235– 253, September 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Finnie:1993:PSD

[FWP93] Gavin R. Finnie, Gerhard E.

Wittig, and Doncho I. Petkov. Prioritizing software development productivity factors using the analytic hierarchy process. The Journal of Systems and Software, 22(2):129–139, August 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Foreman:1993:SEC

[FZ93] Lynn M. Foreman and Stuart H.

Zweben. A study of the effectiveness of control and data flow testing strategies. The Journal of Systems and Software, 21(3): 215–228, June 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fang:1995:MTP

[FZHS95] Ming Fang, Hussein S. M. Zedan, and Chris J. Ho-Stuart. A model for timed-probabilistic behaviors. The Journal of Systems and Software, 28(3):239– 251, March 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Griswold:1995:MDT

[GA95] William G. Griswold and Darren C. Atkinson. Managing design trade-offs for a program understanding and transformation tool. The Journal of Systems and Software, 30(1–2):99–116, July/ August 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gupta:1992:CPA

[GAK92] Meenakshi Gupta, Manju Agarwal, and Ashok Kumar. A comparative performance analysis of a multidisk system under headof-the-line and alternating priority disciplines. The Journal of Systems and Software, 19(1):59– 68, September 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gantenbein:1991:DBS

[Gan91] Rex E. Gantenbein. Dynamic binding in strongly typed programming languages. The Journal of Systems and Software, 14 (1):31–38, January 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gaspoz:1996:MDD

[Gas96] Jean-Paul Gaspoz. Methodology for the development of distributed telecommunications services. The Journal of Systems and Software, 33(3):253– 271, June 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Guan:1992:MPS

[GAW92] Sheng-Uei Guan and Hussein Abdel-Wahab. Multiuser processes and shared capability lists for shared workspace collaboration. The Journal of Systems and Software, 19(1):3–12, September 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Guan:1991:JOO

[GAWC91] Sheng-Uei Guan, Hussein Abdel-Wahab, and Peter Calingaert. Jointly-owned objects for collaboration: Operating-system support and protection model. The Journal of Systems and Software, 16(2):85–95, October 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gowen:1994:ATV

[GC94] Lon D. Gowen and James S.

Collofello. Assessing traditional verification’s effectiveness on safety-critical software systems. The Journal of Systems and Software, 26(2):103– 115, August 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Grasso:1986:PAC

[GDF86] P. A. Grasso, T. S. Dillon, and K. E. Forward. Performance analysis of common bus multimicroprocessor systems. The Journal of Systems and Software, 6 (1–2):71–79, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gentleman:1983:HAH

[GH83] W. Morven Gentleman and Henry Hoeksma. Hardware assisted high-level debugging. The Journal of Systems and Software, 3(4):309–314, December 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gaviotis:1991:CSE

[GHC91] I. Gaviotis, A. Hatzimanikatis, and D. Christodoulakis. Communicating in software engineering environments: The GRASPIN paradigm. The Journal of Systems and Software, 15(1):83–89, April 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Girardi:1995:UER

[GI95] M. R. Girardi and B. Ibrahim. Using English to retrieve software. The Journal of Systems and Software, 30(3):249– 270, September 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Giese:1979:PCC

[Gie79] Clarence Giese. Partitioning considerations for complex computer based weapon systems. The Journal of Systems and Software, 1(1):3–18, ???? 1979. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gilb:1988:PND

[Gil88] Tom Gilb. The pre-natal death of the CIS project: a software disaster story. The Journal of Systems and Software, 8(3):161– 163, June 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gantenbein:1988:DID

[GJ88] Rex E. Gantenbein and Douglas W. Jones. The design and implementation of a dynamic binding feature for a high-level language. The Journal of Systems and Software, 8(4):259–273, September 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Guidec:1996:OOF

[GJP96] F. Guidec, J.-M. Jézéquel, and J.-L. Pacherie. An objectoriented framework for supercomputing. The Journal of Systems and Software, 33(3):239– 251, June 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Geenens:1991:ISC

[GK91a] Ronald Geenens and Mustafa Kamal. An industry study of computer practices. The Journal of Systems and Software, 16 (1):75–80, September 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gerlach:1991:FDH

[GK91b] James H. Gerlach and Feng-Yang Kuo. Formal development of hybrid user-computer interfaces with advanced forms of user assistance. The Journal of Systems and Software, 16(3): 169–183, November 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gerogiannis:1998:CSC

[GKP98] Vasilis C. Gerogiannis, Achilles D. Kameas, and Panayotis E. Pintelas. Comparative study and categorization of high-level Petri nets. The Journal of Systems and Software, 43(2):133– 160, November 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/2/6056.pdf.

Glass:1986:ECD

[Gla86] Robert L. Glass. Editor’s corner: “Dangerous and misleading”: A look at software research via the Parnas papers. The Journal of Systems and Software, 6(3):217– 218, August 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1988:ECM

[Gla88a] Robert L. Glass. Editor’s corner: a mythology of technology transfer. The Journal of Systems and Software, 8(4):257, September 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1988:ECR

[Gla88b] Robert L. Glass. Editor’s corner:

Robert L. Glass: Software technology transfer: a multiflawed process (the road to productivity is full of potholes). The Journal of Systems and Software, 8(5): 339–341, December 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1988:ECS

[Gla88c] Robert L. Glass. Editor’s corner:

Software productivity improvement: Who’s doing what? The Journal of Systems and Software, 8(3):159–160, June 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1989:EC

[Gla89a] Robert L. Glass. Editor’s corner.

The Journal of Systems and Software, 9(1):1–2, January 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1989:ECHa

[Gla89b] Robert L. Glass. Editor’s corner: How about next year? A look at a study of technology maturation. The Journal of Systems and Software, 9(3):167–168, March 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1989:ECHb

[Gla89c] Robert L. Glass. Editor’s corner:

How can computer science truly become a science, and software engineering truly become engineering? The Journal of Systems and Software, 10(2):75–76, September 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1989:ECM

[Gla89d] Robert L. Glass. Editor’s corner: My trivial/brilliant concept called “problem solving”. The Journal of Systems and Software, 9(2):77–78, February 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1989:ECSc

[Gla89e] Robert L. Glass. Editor’s corner: Should a journal publish a virus? The Journal of Systems and Software, 10(4):229, November 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1989:ECSb

[Gla89f] Robert L. Glass. Editor’s corner: Software metrics: of lightning rods and built-up tension. The Journal of Systems and Software, 10(3):157–158, October 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1989:ECSa

[Gla89g] Robert L. Glass. Editor’s corner: Some thoughts on software errors. The Journal of Systems and Software, 9(4):243–244, May 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1989:ECL

[Gla89h] Robert L. Glass. Editor’s corner: The legend of the bad software project. The Journal of Systems and Software, 10(1):1, July 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1989:TRB

[Gla89i] Robert L. Glass. The temporal relationship between theory and practice. The Journal of Systems and Software, 10(1):65–67, July 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1990:ECL

[Gla90a] Robert L. Glass. Editor’s corner: At last — some measured results of the “structured revolution”. The Journal of Systems and Software, 13(1):1, September 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1990:ECA

[Gla90b] Robert L. Glass. Editor’s corner: Automatic programming — a cocktail party myth? The Journal of Systems and Software, 13(2):85–86, October 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1990:ECMb

[Gla90c] Robert L. Glass. Editor’s corner:

Mikhail Gorbachev and software productivity (!?). The Journal of Systems and Software, 13(3):

151–152, November 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1990:ECSa

[Gla90d] Robert L. Glass. Editor’s corner: Software design: Is there madness in a method? The Journal of Systems and Software, 11(1): 1–2, January 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1990:ECSb

[Gla90e] Robert L. Glass. Editor’s corner: Software maintenance is a solution — not a problem. The Journal of Systems and Software, 11 (2):77–78, February 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1990:ECSc

[Gla90f] Robert L. Glass. Editor’s corner: Some pros and cons on (of all things!) references. The Journal of Systems and Software, 11(3): 157, March 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1990:ECMa

[Gla90g] Robert L. Glass. Editor’s corner: The mystery of the little Chinese boy and his computer. The Journal of Systems and Software, 12 (1):1–2, April 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1990:ECT

[Gla90h] Robert L. Glass. Editor’s corner: Theory versus practice — revisited. The Journal of Systems and Software, 12(2):81–82, May 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1991:ECCa

[Gla91a] Robert L. Glass. Editor’s corner: 4GLs and CASE: What’s the payoff? The Journal of Systems and Software, 14(3):131– 132, March 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1991:ECR

[Gla91b] Robert L. Glass. Editor’s corner: a report from the best and brightest. The Journal of Systems and Software, 14(1):1– 2, January 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1991:ECF

[Gla91c] Robert L. Glass. Editor’s corner:

Formal methods vs. heuristics: Clarifying a controversy. The Journal of Systems and Software, 15(2):103–105, May 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1991:ECQ

[Gla91d] Robert L. Glass. Editor’s corner:

Quality software: What is management’s role? The Journal of Systems and Software, 16(2):83– 84, October 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1991:ECSa

[Gla91e] Robert L. Glass. Editor’s corner: Software maintenance: It’s time to educate the educators. The Journal of Systems and Software, 14(2):65–66, February 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1991:ECSd

[Gla91f] Robert L. Glass. Editor’s corner: Software productivity: What is management’s role? The Journal of Systems and Software, 16(3): 167–168, November 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1991:ECSc

[Gla91g] Robert L. Glass. Editor’s corner: Structured research? (A partly tongue-in-cheek look). The Journal of Systems and Software, 16 (1):1–2, September 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1991:ECCb

[Gla91h] Robert L. Glass. Editor’s corner: The computing communication chasm: Some new insight into why it is there, and what can be done about it. The Journal of Systems and Software, 15 (1):1–2, April 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1991:ECSb

[Gla91i] Robert L. Glass. Editor’s corner: The (solved, unsolved) problem of literature searches. The Journal of Systems and Software, 15 (3):203–204, July 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1992:CAT

[Gla92a] Robert L. Glass. A comparative analysis of the topic areas of computer science, software engineering, and information systems. The Journal of Systems and Software, 19(3): 277–289, November 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1992:ECF

[Gla92b] Robert L. Glass. Editor’s corner:

Finally — someone who understands where software quality fits in. The Journal of Systems and Software, 19(3):205–206, November 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1992:ECBb

[Gla92c] Robert L. Glass. Editor’s corner: On becoming obsolete. The Journal of Systems and Software, 18(1):3–4, April 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1992:ECBa

[Gla92d] Robert L. Glass. Editor’s corner: The best and worst of software in the 1980s. The Journal of Systems and Software, 17 (2):109–110, February 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1992:ECIa

[Gla92e] Robert L. Glass. Editor’s corner: The importance of software quality in the 1990s. The Journal of Systems and Software, 17 (3):203–204, March 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1992:ECIb

[Gla92f] Robert L. Glass. Editor’s corner: The importance of the application domain cluster. The Journal of Systems and Software, 18(3): 205–206, July 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1992:ECL

[Gla92g] Robert L. Glass. Editor’s corner: The link between software quality and software maintenance. The Journal of Systems and Software, 19(1):1–2, September 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1992:ECW

[Gla92h] Robert L. Glass. Editor’s corner: We have lost our way. The Journal of Systems and Software, 18(2):111–112, May 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1993:ECC

[Gla93a] Robert L. Glass. Editor’s corner: Can English majors write maintenance documentation? The Journal of Systems and Software, 21(1):1–2, April 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1993:ECD

[Gla93b] Robert L. Glass. Editor’s corner: in defense of adhocracy. The Journal of Systems and Software, 22(3):149–150, September 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1993:ECO

[Gla93c] Robert L. Glass. Editor’s corner: Object-orientation: Theory and practice. The Journal of Systems and Software, 20(1):1– 2, January 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1993:ECU

[Gla93d] Robert L. Glass. Editor’s corner: Of understanding, acceptance, … and formal methods. The Journal of Systems and Software, 21(2):115–116, May 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1993:ECS

[Gla93e] Robert L. Glass. Editor’s corner: Software estimation is not a rational process. The Journal of Systems and Software, 23(1): 1–2, October 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1993:ECG

[Gla93f] Robert L. Glass. Editor’s corner: The gold we’ve failed to mine. The Journal of Systems and Software, 22(2):79–80, August 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1993:ECM

[Gla93g] Robert L. Glass. Editor’s corner: The many flavors of testing. The Journal of Systems and Software, 20(2):105–106, February 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1993:ECW

[Gla93h] Robert L. Glass. Editor’s corner: Which do you think? Modern methods will lead to less software maintenance, or more? The Journal of Systems and Software, 23(3):209–210, December 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1993:EDW

[Gla93i] Robert L. Glass. Error detection: Which is better, reviews or testing? The Journal of Systems and Software, 22(1):1–2, July 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1994:ASS

[Gla94a] Robert L. Glass. Assessment of systems and software engineering scholars and institutions. The Journal of Systems and Software, 27(1):63–67, October 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1994:ECTb

[Gla94b] Robert L. Glass. Editor’s corner: a tabulation of topics where software practice leads software theory. The Journal of Systems and Software, 25(3):219– 222, June 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1994:ECDa

[Gla94c] Robert L. Glass. Editor’s corner: Discipline and creativity: An odd couple. The Journal of Systems and Software, 26(3):195, September 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1994:ECDb

[Gla94d] Robert L. Glass. Editor’s corner:

Discipline is a dirty word: a story about the software life cycle. The Journal of Systems and Software, 27(1):1–2, October 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1994:ECM

[Gla94e] Robert L. Glass. Editor’s corner: My readers write: Some thoughts on formal methods. The Journal of Systems and Software, 27(2):71–72, November 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1994:ECO

[Gla94f] Robert L. Glass. Editor’s corner: Object-orientation: Theory and Practice (II). The Journal of Systems and Software, 24(2):91– 93, February 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1994:ECF

[Gla94g] Robert L. Glass. Editor’s corner: The faking of software design. The Journal of Systems and Software, 26(2):101– 102, August 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1994:ECTa

[Gla94h] Robert L. Glass. Editor’s corner: Theory + practice: a disturbing example. The Journal of Systems and Software, 25(2): 125–126, May 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1994:ECU

[Gla94i] Robert L. Glass. Editor’s corner: User satisfaction. The Journal of Systems and Software, 24(1): 1–2, January 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:ECC

[Gla95a] Robert L. Glass. Editor’s corner: a close escape from greatness. The Journal of Systems and Software, 28(2):93–95, February 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:ECT

[Gla95b] Robert L. Glass. Editor’s corner: a theory about software’s practice. The Journal of Systems and Software, 28(3):187– 188, March 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:ECA

[Gla95c] Robert L. Glass. Editor’s corner:

An assessment of systems and software engineering scholars and institutions, 1993 and 1994. The Journal of Systems and Software, 31(1):3–6, October 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:ECB

[Gla95d] Robert L. Glass. Editor’s corner:

Beyond formal methods. The Journal of Systems and Software, 29(2):95–96, May 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:ECS

[Gla95e] Robert L. Glass. Editor’s corner: In search of an obvious yet radical idea. The Journal of Systems and Software, 31(1):1– 2, October 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:ECP

[Gla95f] Robert L. Glass. Editor’s corner: Plenty of pitfalls: There are numbers and then there are numbers. The Journal of Systems and Software, 31(2):95–96, November 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:RCE

[Gla95g] Robert L. Glass. Readers’ corner: an exchange with an anonymous reader, and a call for help from either researchers or librarians. The Journal of Systems and Software, 30(1–2):160, July/ August 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:SAF

[Gla95h] Robert L. Glass. Software automation: Fact or fraud? The Journal of Systems and Software, 31(3):183–184, December 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:SBC

[Gla95i] Robert L. Glass. Structure-based critique of contemporary computing research. The Journal of Systems and Software, 28(1):3– 7, January 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1995:TCS

[Gla95j] Robert L. Glass. Turning the corner… to the same old street: a fundamental nonchange in JSS policy. The Journal of Systems and Software, 28(1):1–2, January 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ASS

[Gla96a] Robert L. Glass. An assessment of systems and software engineering scholars and institutions (1993–1995). The Journal of Systems and Software, 35 (1):85–89, October 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECA

[Gla96b] Robert L. Glass. Editor’s corner: About software engineering hype. The Journal of Systems and Software, 34(1):1, July 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECC

[Gla96c] Robert L. Glass. Editor’s corner: Computer support for creativity. The Journal of Systems and Software, 34(2):85–87, August 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECF

[Gla96d] Robert L. Glass. Editor’s corner: Formal methods: a dramatic (success, failure) story. The Journal of Systems and Software, 32(2):99–100, February 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECMa

[Gla96e] Robert L. Glass. Editor’s corner: Mathematics and the computer scientist. The Journal of Systems and Software, 34(3): 173–176, September 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECMb

[Gla96f] Robert L. Glass. Editor’s corner: Methodologies: Bend to fit? The Journal of Systems and Software, 35(2):93–94, November 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECO

[Gla96g] Robert L. Glass. Editor’s corner: Object-orientation claims: Naturalness, seamlessness seem doubtful. The Journal of Systems and Software, 33(1):1–2, April 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECSa

[Gla96h] Robert L. Glass. Editor’s corner: Study supports existence of software crisis; management issues appear to be prime cause. The Journal of Systems and Software, 32(3):183–184, March 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECN

[Gla96i] Robert L. Glass. Editor’s corner: The non-goal-oriented committee meeting. The Journal of Systems and Software, 35(3): 167–168, December 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECSb

[Gla96j] Robert L. Glass. Editor’s corner: The strange case of the proofreader’s pencil. The Journal of Systems and Software, 35(1):1– 2, October 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1996:ECW

[Gla96k] Robert L. Glass. Editor’s corner: What has happened to the 1995 “Nonchange” in JSS. The Journal of Systems and Software, 32(1):1–2, January 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ASS

[Gla97a] Robert L. Glass. Assessment of systems and software engineering scholars and institutions (1993– 1996). The Journal of Systems and Software, 39(1):83–88, October 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:CID

[Gla97b] Robert L. Glass. COBOL: Is it dying or thriving? The Journal of Systems and Software, 39(3): 197–199, December 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ECP

[Gla97c] Robert L. Glass. Editor’s corner:

Are programmers really “out of control”? The Journal of Systems and Software, 36(2):101– 102, February 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ECCa

[Gla97d] Robert L. Glass. Editor’s corner: Client/server: Boom or bust? The Journal of Systems and Software, 37(2):99–100, May 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ECCb

[Gla97e] Robert L. Glass. Editor’s corner: Computer science Ph.D. production — more institutions, fewer degrees, more students (!). The Journal of Systems and Software, 39(2):91, November 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ECG

[Gla97f] Robert L. Glass. Editor’s corner: “good enough” software — a unique American advantage. The Journal of Systems and Software, 38(2):105– 106, August 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ECTb

[Gla97g] Robert L. Glass. Editor’s corner: Is there anything “timehonored” in the field of software? The Journal of Systems and Software, 38(3):195– 196, September 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ECSb

[Gla97h] Robert L. Glass. Editor’s corner: Some brainstorming thoughts about the date crisis problem. The Journal of Systems and Software, 37(3):177–179, June 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ECSa

[Gla97i] Robert L. Glass. Editor’s corner: Some thoughts on automatic code generation. The Journal of Systems and Software, 37(1):1–3, April 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ECTc

[Gla97j] Robert L. Glass. Editor’s corner: Top scholars and institutions study reaching maturity: a lesson re-learned. The Journal of Systems and Software, 39(1): 1–2, October 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:ECTa

[Gla97k] Robert L. Glass. Editor’s corner:

Two years later: What has happened to the 1995 “nonchange” in JSS policy? The Journal of Systems and Software, 36(1): 1, January 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:PSW

[Gla97l] Robert L. Glass. Pilot studies:

What, why and how. The Journal of Systems and Software, 36 (1):85–97, January 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1997:SRP

[Gla97m] Robert L. Glass. Software reflections — a pioneer’s view of the history of the field. The Journal of Systems and Software, 36 (3):203–209, March 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1998:AC

[Gla98a] Robert L. Glass. Anecdote corner. The Journal of Systems and Software, 43(1):65–176, October 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1998:ASS

[Gla98b] Robert L. Glass. An assessment of systems and software engineering scholars and institutions (1993–1997). The Journal of Systems and Software, 43(1): 59–64, October 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/43/1/6065.pdf. See corri-

gendum [Gla99b].

Glass:1998:EECa

[Gla98c] Robert L. Glass. Editorial — Editor’s corner: a success/failure story. The Journal of Systems and Software, 43(2):83–84, November 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1998:EECb

[Gla98d] Robert L. Glass. Editorial — Editor’s corner. Webwork: Is it systems development? how can we learn about doing it better? The Journal of Systems and Software, 44(1):1–2, December 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1998:EWW

[Gla98e] Robert L. Glass. Editorial: a word of warning about reuse. The Journal of Systems and Software, 41(3):145–146, June 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1998:ECC

[Gla98f] Robert L. Glass. Editor’s corner: Cleaning up the trash after Y2K gets “solved”. The Journal of Systems and Software, 43(3): 161–163, November 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1998:ECE

[Gla98g] Robert L. Glass. Editor’s corner.

expert systems: Failure or success? The Journal of Systems and Software, 43(1):1–2, October 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1998:ECH

[Gla98h] Robert L. Glass. Editor’s corner: How long do information systems last? The Journal of Systems and Software, 40(2):91, February 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1998:ECS

[Gla98i] Robert L. Glass. Editor’s Corner: Software runaways — Some surprising findings. The Journal of Systems and Software, 41 (2):75–77, May 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/2/6008.pdf.

Glass:1998:ECW

[Gla98j] Robert L. Glass. Editor’s corner: “who cares?” technologies in practice. The Journal of Systems and Software, 41(1): 1–2, April 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1998:ECY

[Gla98k] Robert L. Glass. Editor’s corner: yet another date crisis — Y2k is only the beginning. The Journal of Systems and Software, 40(1): 1–2, January 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1999:ASS

[Gla99a] Robert L. Glass. An assessment of systems and software engineering scholars and institutions (1994–1998). The Journal of Systems and Software, 49 (1):81–86, December 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/25/32/article.pdf;

http://www.elsevier.nl/gej-
ng/10/29/11/46/25/abstract.

html. See corrigendum [Gla00].

Glass:1999:CAS

[Gla99b] Robert L. Glass. Corrigendum to “An assessment of systems and software engineering scholars and institutions (1993–1997)” [Journal of Systems and Software 43 (1998) 59– 64]. The Journal of Systems and Software, 48(1):73, August 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/
29/11/37/17/23/;   http://
www.elsevier.nl/gej-ng/10/
29/11/37/17/23/article.pdf.

See [Gla98b].

Glass:1999:EEA

[Gla99c]  Robert L. Glass.  Examin-

ing the effects of the “Application Revolution”. The Journal of Systems and Software, 46 (1):1–5, April 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/1/6110.pdf.

Glass:1999:CSE

[Gla99d] Robert L. Glass. Of counting schemes and end-of-century partying. The Journal of Systems and Software, 47(1):1– 2, May 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.

                                        com/cas/tree/store/jss/sub/
                                        1999/47/1/6132.pdf.

Glass:2000:CAS

[Gla00] Robert L. Glass. Corrigendum to: an assessment of systems and software engineering scholars and institutions (1994–1998) [The Journal of Systems and Software 49 (1) (1999) 81–86]. The Journal of Systems and Software, 51(3):275, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic).  URL http://www.
                                        elsevier.nl/gej-ng/10/29/
                                        11/50/28/33/article.pdf;
                                        http://www.elsevier.nl/gej-
                                        ng/10/29/11/50/28/abstract.

html. See [Gla99a].

Ghezzi:1990:TLL

[GMM90] Carlo Ghezzi, Dino Mandrioli, and Angelo Morzenti. TRIO: a logic language for executable specifications of real-time systems. The Journal of Systems and Software, 12(2):107– 123, May 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ghezzi:1994:VTR

[GMP94] Carlo Ghezzi, Sandro Morasca, and Mauro Pezzè. Validating timing requirements for time basic net specifications. The Journal of Systems and Software, 27 (2):97–117, November 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Goel:1980:SED

[Goe80] Amrit L. Goel. Software error detection model with applications. The Journal of Systems and Software, 1(3):243–249, ???? 1980. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Goel:1984:ISI

[Goe84] Amrit L. Goel. Introduction to the special issue on the Fifth Minnowbrook Workshop on Software Performance Evaluation. The Journal of Systems and Software, 4(4):267, November 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gomaa:1989:SDM

[Gom89] H. Gomaa. A software design method for distributed real-time applications. The Journal of Systems and Software, 9(2):81– 94, February 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gomaa:1994:SDM

[Gom94] Hassan Gomaa. Software design methods for the design of largescale real-time systems. The Journal of Systems and Software, 25(2):127–146, May 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gomaa:1995:RSR

[Gom95] Hassan Gomaa. Reusable software requirements and architectures for families of systems. The Journal of Systems and Software, 28(3):189–202, March 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gonzalez:1995:UMS

[Gon95] Renato R. Gonzalez. A unified metric of software: Complexity: Measuring productivity, quality, and value. The Journal of Systems and Software, 29(1): 17–37, April 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gorla:1991:PHD

[Gor91] Narasimhaiah Gorla. Performance of a hybrid DBMS using a case study. The Journal of Systems and Software, 14(3):199– 214, March 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gotterbarn:1990:WRS

[Got90] Donald Gotterbarn. A workshop report: Software Engineering Ethics. The Journal of Systems and Software, 11(3):215– 216, March 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gotterbarn:1992:ECP

[Got92a] Donald Gotterbarn. Editor’s corner: Professional computer ethics. The Journal of Systems and Software, 17(1):5–6, January 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gotterbarn:1992:UAC

[Got92b] Donald Gotterbarn. The use and abuse of computer ethics. The Journal of Systems and Software, 17(1):75–80, January 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gotterbarn:1993:GEC

[Got93] Donald Gotterbarn. Guest Editor’s corner: Object-oriented software development: a full life cycle. The Journal of Systems and Software, 23(2):83–84, November 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Golshani:1998:UIM

[GP98] Forouzan Golshani and Antonio Pizzarello. The use of immense memories for database systems: a pragmatic approach. The Journal of Systems and Software, 41 (1):41–50, April 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/1/6005.pdf.

Gorla:1997:ESS

[GR97] Narasimhaiah Gorla and Ravi Ramakrishnan. Effect of software structure attributes on software development productivity. The Journal of Systems and Software, 36(2):191–199, February 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gopinath:1992:DBD

[GRS92] Prabha Gopinath, Rajiv Ramnath, and Karsten Schwan. Data base design for real-time adaptations. The Journal of Systems and Software, 17(2):155– 167, February 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Goutas:1991:GDB

[GSC91] S. Goutas, P. Soupos, and D. Christodoulakis. The GRASPIN data base. The Journal of Systems and Software, 15(1):75–82, April 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gulezian:1991:RCC

[Gul91] Ronald Gulezian. Reformulating and calibrating COCOMO. The Journal of Systems and Software, 16(3):235–242, November 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gulezian:1992:RC

[Gul92] Ronald Gulezian. Readers’ corner. The Journal of Systems and Software, 18(1):2, April 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gulezian:1996:HRS

[Gul96] Ronald Gulezian. Handling regression subsets in software modeling. The Journal of Systems and Software, 33(1):81–86, April 1996. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic). See response [BT97].

Glass:1992:TTS

[GV92] Robert L. Glass and Iris Vessey.

Toward a taxonomy of software application domains: History. The Journal of Systems and Software, 17(2):189–199, February 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glykas:1999:FMO

[GV99] Michalis Glykas and George Valiris. Formal methods in object oriented business modelling. The Journal of Systems and Software, 48(1):27– 41, August 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/29/11/37/17/
20/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/
29/11/37/17/20/article.pdf.

Griss:1995:HDS

[GW95] Martin L. Griss and Kevin D.

Wentzel. Hybrid domain-specific kits. The Journal of Systems and Software, 30(3):213– 230, September 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Habermann:1985:ADO

[Hab85] A. Nico Habermann. Automatic deletion of obsolete information. The Journal of Systems and Software, 5(2):145–154, May 1985.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1986:MPA

[Hać86a] Anna Hać. Modelling parallel access to shared resources in a distributed file system using queueing networks. The Journal of Systems and Software, 6(1–2): 61–69, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1986:PID

[Hać86b] Anna Hać. Performancereliability issues in distributed file systems. The Journal of Systems and Software, 6(3):219–224, August 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1988:MMC

[Hać88] Anna Hać. A multiprocessor model with classes of processes and priorities. The Journal of Systems and Software, 8(3):247– 251, June 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1989:BPE

[Haç89a] A. Haç. A benchmark for performance evaluation of a distributed file system. The Journal of Systems and Software, 9(2):273–285, February 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1989:VPM

[Hać89b] Anna Hać. A validated performance model for distributed file systems. The Journal of Systems and Software, 10(3):169– 185, October 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1991:DAA

[Hać91] Anna Hać. Design algorithms for asynchronous write operations in disk-buffer-cache memory. The Journal of Systems and Software, 16(3):243–253, November 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1992:PAP

[Hać92] Anna Hać. Performance analysis of the priority queues for the load-building algorithms. The Journal of Systems and Software, 19(1):69–74, September 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1993:DAM

[Hać93] Anna Hać. A distributed algorithm for managing resources in a hierarchical network. The Journal of Systems and Software, 22(3):191–206, September 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1994:DMA

[Hać94] Anna Hać. Distributed multicasting algorithm with congestion control and message routing. The Journal of Systems and Software, 24(1):49–65, January 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hager:1991:SCR

[Hag91] J. A. Hager. Software cost reduction methods in practice: a postmortem analysis. The Journal of Systems and Software, 14(2):67– 77, February 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Halang:1992:RTS

[Hal92] Wolfgang A. Halang. Real-time systems: Another perspective. The Journal of Systems and Software, 18(1):101–108, April 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hamlet:1981:HEC

[Ham81] Richard Hamlet. Hard-to-use evaluation criteria for software engineering. The Journal of Systems and Software, 2(2):89–96, June 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1981:ETI

[Har81] P. G. Harrison. Efficient tabledriven implementation of the finite state machine. The Journal of Systems and Software, 2(3): 201–211, September 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1988:ISI

[Har88a] Warren Harrison. Introduction to the special issue on software engineering. The Journal of Systems and Software, 8(1):1– 2, January 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1988:MSM

[Har88b] Warren Harrison. Mae: a syntactic metric analysis environment. The Journal of Systems and Software, 8(1):57–62, January 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1990:FSI

[Har90a] Warren Harrison. A foreword to the special issue on using software metrics. The Journal of Systems and Software, 13(2):87– 88, October 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1990:GEI

[Har90b] Warren Harrison. Guest Editor’s introduction: The Oregon Workshop on Software Metrics. The Journal of Systems and Software, 12(3):175, July 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1993:GEI

[Har93] Warren Harrison. Guest Editor’s introduction. The Journal of Systems and Software, 20 (3):205–206, March 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1994:GEI

[Har94] Warren Harrison. Guest Editor’s introduction. The Journal of Systems and Software, 24(3): 209, March 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harris:1995:WRO

[Har95a] Daniel P. Harris. Controversy corner: Who really owns “your” software programs? The Journal of Systems and Software, 29(2): 191–193, May 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1995:GEC

[Har95b] Warren Harrison. Guest Editor’s corner. The Journal of Systems and Software, 29(1):1, April 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hardgrave:1997:AOO

[Har97] Bill C. Hardgrave. Adopting object-oriented technology: Evolution or revolution? The Journal of Systems and Software, 37(1):19–25, April 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hartson:1998:HCI

[Har98]  H. Rex Hartson.  Human–

computer interaction: Interdisciplinary roots and trends. The Journal of Systems and Software, 43(2):103–118, November 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/2/6054.pdf.

Harrold:1999:TES

[Har99] Mary Jean Harrold. Testing evolving software. The Journal of Systems and Software, 47(2– 3):173–181, July 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/47/2-3/6152.pdf.

Hasselbring:1998:PLA

[Has98]  Wilhelm Hasselbring.  The

ProSet–Linda approach to prototyping parallel systems. The Journal of Systems and Software, 43(3):187–196, November 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/3/6060.pdf.

Hatton:1999:RFF

[Hat99] Les Hatton. Repetitive failure, feedback and the lost art of diagnosis. The Journal of Systems and Software, 47(2–3):183–188, July 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/47/2-3/
6153.pdf.

Hayes:1986:PPE

[Hay86]  Ann H. Hayes.  Parallel-

processing experiences on the Denelcor HEP computer. The Journal of Systems and Software, 6(1–2):7–10, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hartson:1983:MER

[HB83] H. Rex Hartson and Earl J. Balliet. Model of enforcement relationships among database access control dependencies. The Journal of Systems and Software, 3(3):201–217, September 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Howatt:1989:RDA

[HB89] J. W. Howatt and A. L. Baker. Rigorous definition and analysis of program complexity meaures — an example using nesting. The Journal of Systems and Software, 10(2):139–150, September 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Halliday:1994:ETS

[HBCC94] Michael Halliday, Inderpal Bhandari, Jarir Chaar, and Ram Chillarege. Experiences in transferring a software process improvement methodology to production laboratories. The Journal of Systems and Software, 26(1):61–68, July 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Howard:1999:EMI

[HBJ^{+}99] Geoffry S. Howard, Thomas Bodnovich, Thomas Janicki, Jens Liegle, Steven Klein, Paul Albert, and David Cannon. The efficacy of matching information systems development methodologies with application characteristics — an empirical study. The Journal of Systems and Software, 45(3):177– 195, March 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/3/6104.pdf.

Harrison:1986:DNC

[HC86] Warren Harrison and Curtis Cook. Are deeply nested conditionals less readable? The Journal of Systems and Software, 6(4):335–341, November 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1987:MMM

[HC87] Warren Harrison and Curtis Cook. Micro/macro measure of software complexity. The Journal of Systems and Software, 7(3):213–219, September 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ho:1991:RMP

[HCC91] Tsu-Feng Ho, Wah-Chun Chan, and Chyan-Goei Chung. Reliability modeling of parallel programs with modular structure using stochastic Petri nets. The Journal of Systems and Software, 16(2):121–128, October 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hirschhorn:1984:PBS

[HD84] S. Hirschhorn and A. M. Davis. Parallels between software and VLSI engineering. The Journal of Systems and Software, 4(1): 27–37, April 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Heeman:1990:IPE

[Hee90] Frans C. Heeman. Incremental parsing of expressions. The Journal of Systems and Software, 13 (1):55–69, September 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Heim:1995:CRM

[Hei95] Michael Heim. Controversy revisited: The Mythseeker Project. The Journal of Systems and Software, 28(1):89, January 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Henry:1988:THP

[Hen88] Sallie Henry. Technique for hiding proprietary details while providing sufficient information for researchers; or, do you recognize this well-known algorithm? The Journal of Systems and Software, 8(1):3–11, January 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Henninger:1995:IAT

[Hen95] Scott Henninger. Information access tools for software reuse. The Journal of Systems and Software, 30(3):231–247, September 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hetzel:1995:SSS

[Het95] William C. Hetzel. The sorry state of software practice measurement and evaluation. The Journal of Systems and Software, 31(2):171–179, November 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huynh:1992:WMF

[HFK92] Khoa D. Huynh, Eduardo B. Fernandez, and Taghi M. Khoshgoftaar. A workload model for frame-based real-time applications on distributed systems. The Journal of Systems and Software, 18(3):255–271, July 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Henry:1991:CGT

[HG91] Sallie Henry and Roger Goff.

Comparison of a graphical and a textual design language using software quality metrics. The Journal of Systems and Software, 14(3):133–146, March 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Horspool:1987:ADD

[HH87] R. Nigel Horspool and Ronald M.

Huberman. Analysis and development of demand prepaging policies. The Journal of Systems and Software, 7(3):183–194, September 1987. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Heiat:1997:MEE

[HH97] Abbas Heiat and Nafisseh Heiat. A model for estimating efforts required for developing small-scale business applications. The Journal of Systems and Software, 39 (1):7–14, October 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hilburn:1999:GDS

[HHB^{+}99] Thomas B. Hilburn, Greg Hislop, Donald J. Bagert, Michael Lutz, Susan Mengel, and Michael Mc-Cracken. Guidance for the development of software engineering education programs. The Journal of Systems and Software, 49(2–3):163–169, December 30, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

elsevier.nl/gej-ng/10/29/
11/46/26/32/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

Huang:1997:EBI

[HHL^{+}97] Chung-Ming Huang, Jenq-Muh Hsu, Huei-Yang Lai, Duen-Tay Huang, and Jao-Chiang Pong. An Estelle-based incremental protocol design system. The Journal of Systems and Software, 36(2):115–135, February 1997.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hix:1994:CRE

[HHSR94] Deborah Hix, H. Rex Hartson, Antonio C. Siochi, and David Ruppert. Customer responsibility for ensuring usability: Requirements on the user interface development process. The Journal of Systems and Software, 25 (3):241–255, June 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hashim:1992:PKB

[HHZ92] F. Hashim, S. E.-D. Habib, and M. Zaki. A PC knowledge-based environment for VLSI layout verification. The Journal of Systems and Software, 18(1):19–31, April 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hislop:1998:AES

[His98] Gregory W. Hislop. Analyzing existing software for software reuse. The Journal of Systems and Software, 41(1): 33–40, April 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/1/6004.pdf.

Hac:1990:DLB

[HJ90a] Anna Hać and Xiaowei Jin. Dynamic load balancing in a distributed system using a senderinitiated algorithm. The Journal of Systems and Software, 11 (2):79–94, February 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1990:PCC

[HJ90b] Anna Hać and Theodore J. Johnson. A performance comparison of a closely-coupled and a loosely-coupled architecture. The Journal of Systems and Software, 13(3):201–208, November 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1991:DAD

[HJ91] Anna Hać and Xiaowei Jin. A decentralized algorithm for dynamic load balancing with file transfer. The Journal of Systems and Software, 16(1):37–52, September 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1991:AFR

[HJS91] Anna Hać, Xiaowei Jin, and Jo-Han Soo. Algorithms for file replication in a distributed system. The Journal of Systems and Software, 14(3):173– 181, March 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hasselbring:1998:COP

[HK98] Wilhelm Hasselbring and Andreas Kröber. Combining OMT with a prototyping approach. The Journal of Systems and Software, 43(3):177– 185, November 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/3/6059.pdf.

Horspool:1983:IBS

[HL83] R. Nigel Horspool and Jeffrey M.

Laks. Improved block sequencing method for program restructuring. The Journal of Systems and Software, 3(3):245–250, September 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Henry:1990:IML

[HL90] Sallie Henry and John Lewis. Integrating metrics into a largescale software development environment. The Journal of Systems and Software, 13(2):89–95, October 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Horspool:1993:TBM

[HL93] R. Nigel Horspool and Michael R.

Levy. Translator-based multiparadigm programming. The Journal of Systems and Software, 23(1):39–49, October 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hac:1994:DLB

[HL94a] Anna Hać and Raghavendra Rao Loka. A distributed loadbuilding algorithm for parallel compilation of files in a software application. The Journal of Systems and Software, 26(2):167– 177, August 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Horng:1994:SAO

[HL94b] Jorng-Tzong Horng and Baw-Jhiune Liu. Some aspects of operations in an object-oriented data base based on graphs. The Journal of Systems and Software, 24(2):155–179, February 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:1998:MCE

[HL98] S. J. Huang and R. Lai. On measuring the complexity of an Estelle specification. The Journal of Systems and Software, 40 (2):165–181, February 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:1999:EII

[HLAB99] John Harrison, Wie Ming Lim, Anna Andrusiewicz, and Anthony Berglas. Evaluation of the ITOC information system design recovery tool: a case study. The Journal of Systems and Software, 44(3):229– 240, January 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1999/44/3/6087.pdf.

Hwang:1999:CDC

[HLC99] Ren-Junn Hwang, Wei-Bin Lee, and Chin-Chen Chang. A concept of designing cheater identification methods for secret sharing. The Journal of Systems and Software, 46(1): 7–11, April 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/1/6111.pdf.

Huang:1998:SMP

[HMC98] Yin-Fu Huang, Chien-Chie Mao, and Jyh-Her Chen. A simulation model for the primary copy strategy in distributed INGRES. The Journal of Systems and Software, 41(1):51–61, April 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/1/6006.pdf.

Hawryszkiewycz:1996:CAS

[HMG96] I. T. Hawryszkiewycz, L. A. Maciaszek, and J. R. Getta. Coordination and artifact semantics in asynchronous distributed cooperation. The Journal of Systems and Software, 33(2):179– 188, May 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hadjiefthymiades:1999:SRD

[HMP99] Stathes Hadjiefthymiades, Drakoulis Martakos, and Costas Petrou. Stateful relational database gateways for the World Wide Web. The Journal of Systems and Software, 48(3):177– 187, November 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic).  URL http://www.
                                        elsevier.nl/gej-ng/10/29/
                                        11/37/24/19/article.pdf;
                                        http://www.elsevier.nl/gej-
                                        ng/10/29/11/37/24/abstract.
                                        html.

Ho:1996:LMB

[HO96] W. Wilson Ho and Ronald A. Olsson. A layered model for building debugging and monitoring tools. The Journal of Systems and Software, 34(3):211– 222, September 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hakuta:1997:SSP

[HO97] Mitsuari Hakuta and Masato Ohminami. A study of software portability evaluation. The Journal of Systems and Software, 38 (2):145–154, August 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hoang:1994:GEC

[Hoa94] Pham Hoang. Guest Editor’s corner: Fault tolerance in realtime systems. The Journal of Systems and Software, 25(1):1–2, April 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hon:1990:ASQ

[Hon90] Samuel E. Hon III. Assuring software quality through measurements: a buyer’s perspective. The Journal of Systems and Software, 13(2):117–130, October 1990. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrold:1997:AFM

[HOT97] Mary Jean Harrold, A. Jefferson Offutt, and Kanupriya Tewary. An approach to fault modeling and fault seeding using the program dependence graph. The Journal of Systems and Software, 36(3):273–295, March 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Howden:1980:FTD

[How80] William E. Howden. Functional testing and design abstractions. The Journal of Systems and Software, 1(4):307–313, ???? 1980. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Humenik:1990:PPE

[HP90] Keith Humenik and Roger S.

Pinkham. Production probability estimators for context-free grammars. The Journal of Systems and Software, 12(1):43–53, April 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Humenik:1992:TEC

[HP92] K. Humenik and R. S. Pinkham.

Transposition errors in contextfree languages. The Journal of Systems and Software, 18(3): 273–280, July 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hayne:1995:GDB

[HR95] Stephen Hayne and Sudha Ram. Group data base design: Addressing the view modeling problem. The Journal of Systems and Software, 28(2):97–116, February 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huh:1996:CMF

[HR96] Soon-Young Huh and David A. Rosenberg. A change management framework: Dependency maintenance and change notification. The Journal of Systems and Software, 34(3):231– 246, September 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Henry:1995:QES

[HRS95] Joel Henry, Allan Rossman, and John Snyder. Quantitative evaluation of software process improvement. The Journal of Systems and Software, 28(2):169– 177, February 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hops:1995:DAC

[HS95] Jonathan M. Hops and Joseph S. Sherif. Development and application of composite complexity models and a relative complexity metric in a software maintenance environment. The Journal of Systems and Software, 31(2): 157–169, November 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Henry:1999:UBL

[HS99] Sallie M. Henry and K. Todd Stevens. Using Belbin’s leadership role to improve team effectiveness: an empirical investigation. The Journal of Systems and Software, 44(3):241– 250, January 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/44/3/6088.pdf.

Hsieh:1991:DEC

[Hsi91a] C. Samuel Hsieh. Distributed evaluation of commutative and associative functions. The Journal of Systems and Software, 16 (1):53–58, September 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hsieh:1991:SCD

[Hsi91b] C. Samuel Hsieh. Slice, chunk and dataflow anomaly as DAT-ALOG rules. The Journal of Systems and Software, 16(3): 197–203, November 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hakuta:1997:SSE

[HTO97] Mitsuari Hakuta, Fujimitsu Tone, and Masato Ohminami. A software size estimation model and its evaluation. The Journal of Systems and Software, 37(3): 253–263, June 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hsieh:1996:CSP

[HU96] C. Samuel Hsieh and Elizabeth A. Unger. On the control structure of a program slice. The Journal of Systems and Software, 34(2):123– 126, August 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hsieh:1992:UPD

[HUMT92] C. Samuel Hsieh, Elizabeth A. Unger, and Ramon A. Mata-Toledo. Using program dependence graphs for information flow control. The Journal of Systems and Software, 17(3):227– 232, March 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hurley:1993:MPI

[Hur93] W. David Hurley. A method for predicting the impact of user interface design on software design. The Journal of Systems and Software, 22(3):179–190, September 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hague:1994:DRT

[HW94] Waqar Hague and Johnny Wong. Distributed real-time nested transactions. The Journal of Systems and Software, 27(2):85–95, November 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Humenik:1994:GEC

[HY94] Keith Humenik and Yelena Yesha. Guest Editors’ corner.

The Journal of Systems and Software, 27(3):157–158, December 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hwang:1995:TPE

[HY95] Min-Shiang Hwang and Wei-Pang Yang. A two-phase encryption scheme for enhancing database security. The Journal of Systems and Software, 31(3): 257–265, December 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hamilton:1979:RBD

[HZ79] M. Hamilton and S. Zeldin. The relationship between design and verification. The Journal of Systems and Software, 1(1):29–56, ???? 1979. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hamilton:1983:FLC

[HZ83] M. Hamilton and S. Zeldin.

Functional life cycle model and its automation: Use. it. The Journal of Systems and Software, 3(1):25–62, March 1983. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Haley:1984:DAW

[HZ84] Allen Haley and Stuart Zweben.

Development and application of a white box approach to integration testing. The Journal of Systems and Software, 4(4):309–315, November 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Itzfeldt:1991:PGC

[IKCN91] Wolf D. Itzfeldt, Bernd Krämer, and Marie-Luise Christ-Neumann. The project GRASPIN and its CASE environment prototypes: An overview. The Journal of Systems and Software, 15(1):5– 16, April 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Iannello:1995:PAD

[IMM95] Giulio Iannello, Antonino Mazzeo, and Nicola Mazzocca. Performance analysis of distributed memory computers with parallel node architecture. The Journal of Systems and Software, 29(2): 107–120, May 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Isoda:1995:ESR

[Iso95] Sadahiro Isoda. Experiences of a software reuse project. The Journal of Systems and Software, 30(3):171–186, September 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Isoda:1998:CCR

[Iso98] Sadahiro Isoda. A criticism on the capture-and-recapture method for software reliability assurance. The Journal of Systems and Software, 43(1):3– 102, October 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/1/6045.pdf.

Itzkovitz:1998:TMA

[ISS98] Ayal Itzkovitz, Assaf Schuster, and Lea Shalev. Thread migration and its applications in distributed shared memory systems. The Journal of Systems and Software, 42(1):71–87, July 1, 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.cs.technion.ac.il/Labs/
Millipede/papers/jss.pdf;
 http://www.elsevier.com/
cas/tree/store/jss/sub/1998/
42/1/6028.pdf.

Isoda:1995:EIS

[IYKO95] Sadahiro Isoda, Shuichiro Yamamoto, Hiroaki Kuroki, and Atsuko Oka. Evaluation and introduction of the structured methodology and a CASE tool. The Journal of Systems and Software, 28(1):49–58, January 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jackson:1998:FMT

[Jac98] Michael Jackson. Formal methods and traditional engineering. The Journal of Systems and Software, 40(3):191– 194, March 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jarzabek:1993:DMD

[Jar93] Stan Jarzabek. Domain modeldriven software reengineering and maintenance. The Journal of Systems and Software, 20(1):37– 51, January 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Javey:1988:LCS

[Jav88] S. Javey. A language construct for the specification of discontinuities. The Journal of Systems and Software, 8(5):409–417, December 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jenson:1991:PEP

[JB91] Richard L. Jenson and Jon W. Bartley. Parametric estimation of programming effort: An object-oriented model. The Journal of Systems and Software, 15(2):107–114, May 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jan:1998:PWP

[JC98] Jinn-Ke Jan and Yu-Yii Chen. “paramita wisdom” password authentication scheme without verification tables. The Journal of Systems and Software, 42 (1):45–57, July 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/42/1/6026.pdf.

Jia:1999:CMM

[JCJ99] X. Jia, J. Cao, and W. Jia. A classification of multicast mechanisms: implementations and applications. The Journal of Systems and Software, 45(2):

99–112, March 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/2/6098.pdf.

Jeffery:1987:SDP

[Jef87] D. R. Jeffery. Software development productivity model for MIS environments. The Journal of Systems and Software, 7(2):115– 125, June 1987. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeffrey:1991:HSA

[Jef91] H. Joel Jeffrey. Human systems analysis in the software engineering curriculum. The Journal of Systems and Software, 14 (3):147–153, March 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeffrey:1992:PDM

[Jef92] H. Joel Jeffrey. Pragmatic design of meetings and presentations. The Journal of Systems and Software, 19(1):85–96, September 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeffrey:1996:AED

[Jef96] H. Joel Jeffrey. Addressing the essential difficulties of software engineering. The Journal of Systems and Software, 32(2):157– 179, February 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeng:1999:TID

[Jen99] Bingchiang Jeng. Toward an integration of data flow and domain testing. The Journal of Systems and Software, 45(1): 19–30, February 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/1/6092.pdf.

Jeng:1999:AAD

[JF99] Bingchiang Jeng and István Forgács. An automatic approach of domain test data generation. The Journal of Systems and Software, 49(1):97– 112, December 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/25/35/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/25/abstract.
html.

Johnson:1999:OOM

[JH99] Richard A. Johnson and Bill C. Hardgrave. Object-oriented methods: current practices and attitudes. The Journal of Systems and Software, 48(1):5– 12, August 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/29/11/37/17/
18/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/
29/11/37/17/18/article.pdf.

Jia:1999:COG

[Jia99] Weijia Jia. Communicating object group and protocols for distributed systems. The Journal of Systems and Software, 45(2): 113–126, March 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/2/6099.pdf.

Jeffery:1985:MPP

[JL85] D. Ross Jeffery and Michael J.

Lawrence. Managing programming productivity. The Journal of Systems and Software, 5 (1):49–58, February 1985. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jirachiefpattana:1997:ESE

[JL97] A. Jirachiefpattana and R. Lai.

EVEN: a software environment for Estelle specification verification. The Journal of Systems and Software, 39(2):119–143, November 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jaoua:1990:UEA

[JM90] Ali Jaoua and Ali Mili. The use of executable assertions for error detection and damage assessment. The Journal of Systems and Software, 12(1):15–37, April 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jayaprakash:1996:CAF

[JM96] S. Jayaprakash and C. R. Muthukrishnan. Critical analysis of “A fault-tolerant algorithm for mutual exclusion in a distributed system”. The Journal of Systems and Software, 35(3): 217–221, December 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [WTS95].

Jmaiel:1996:SCP

[Jma96] Mohamed Jmaiel. Specification of communication protocols using temporal logic. The Journal of Systems and Software, 33(3): 299–312, June 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Juristo:1999:FAG

[JMM99] Natalia Juristo, José L. Morant, and Ana M. Moreno. A formal approach for generating OO specifications from natural language. The Journal of Systems and Software, 48(2):139– 153, October 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/37/23/22/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/23/abstract.
html.

Jajodia:1984:TER

[JN84] Sushil Jajodia and Peter A. Ng. Translation of entity-relationship diagrams into relational structures. The Journal of Systems and Software, 4(2–3):123– 133, July 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jajodia:1984:ISI

[JNY84] Sushil Jajodia, Peter Ann-Beng Ng, and Raymond T. Yeh. Introduction to the special issue on the use of entity-relationship concepts in databases and related software. The Journal of Systems and Software, 4(2–3): 95–98, July 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jennings:1983:APE

[JO83] Steven F. Jennings and Arthur E.

Oldehoeft. Analysis of program execution on a recursive streamoriented data flow architecture. The Journal of Systems and Software, 3(2):147–154, June 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Joshi:1983:SDR

[Jos83] Ramchandra D. Joshi. Software development for reliable software systems. The Journal of Systems and Software, 3(2):107–121, June 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Joyce:1987:IIS

[Joy87] Daniel Joyce. Identification and investigation of software design guidelines for using encapsulation units. The Journal of Systems and Software, 7(4):287–295, December 1987. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Joyce:1994:EFG

[Joy94] Daniel T. Joyce. Experience with a fourth-generation language system for software development: An educator’s perspective. The Journal of Systems and Software, 25(2):193–200, May 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeffrey:1994:RDM

[JP94] H. Joel Jeffrey and Anthony O. Putman. Relationship definition and management: Tools for requirements analysis. The Journal of Systems and Software, 24 (3):277–294, March 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jones:1990:IDE

[JS90] A. Russell Jones and Esther R. Steinberg. Interface design: an educational perspective. The Journal of Systems and Software, 11(3):199–204, March 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jaragh:1999:SCP

[JS99] Mansour Jaragh and Kassem Saleh. Synthesis of communications protocol converters using the timed Petri net model. The Journal of Systems and Software, 47(1):53–69, May 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/47/1/6139.pdf.

Jan:1997:SEV

[JT97] Jinn-Ke Jan and Chih-Chang Tai. A secure electronic voting protocol with IC cards. The Journal of Systems and Software, 39(2):93–101, November 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jones:1998:FMR

[JTW98] Sara Jones, David Till, and Ann M. Wrightson. Formal methods and requirements engineering: Challenges and synergies. The Journal of Systems and Software, 40(3):263– 273, March 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jard:1983:ATS

[JvB83] Claude Jard and Gregor v. Bochmann.

Approach to testing specifications. The Journal of Systems and Software, 3(4):315–323, December 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jagadeesan:1998:SBT

[JVP^{+}98] Lalita Jategaonkar Jagadeesan, Lawrence G. Votta, Adam Porter, Carlos Puchol, and J. Christopher Ramming. Specificationbased testing of reactive software: a case study in technology transfer. The Journal of Systems and Software, 40(3):249– 262, March 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khoshnevisan:1996:SEM

[KA96] Hessam Khoshnevisan and Mohamad Afshar. Space-efficient memo-functions. The Journal of Systems and Software, 35(1):29– 41, October 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kallman:1992:DCE

[Kal92] Ernest A. Kallman. Developing a code for ethical computer use. The Journal of Systems and Software, 17(1):69–74, January 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khoshgoftaar:1997:ITB

[KAL97] Taghi M. Khoshgoftaar, Edward B. Allen, and David L. Lanning. An information theorybased approach to quantifying the contribution of a software metric. The Journal of Systems and Software, 36(2):103– 113, February 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kampfner:1989:SAD

[Kam89] Roberto R. Kampfner. A synthetic approach to the design of information-systems software. The Journal of Systems and Software, 10(1):3–14, July 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kamkar:1995:OCC

[Kam95] Mariam Kamkar. An overview and comparative classification of program slicing techniques. The Journal of Systems and Software, 31(3):197–214, December 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Karatza:1994:SSS

[Kar94] Helen D. Karatza. Simulation study of a system with two processors linked in tandem. The Journal of Systems and Software, 26(3):285–292, September 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Karatza:1998:TRR

[Kar98] Helen D. Karatza. Task routing and resequencing in a multiprocessor system. The Journal of Systems and Software, 41 (3):189–197, June 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/3/6019.pdf.

Karimi:1996:SMS

[KB96] Jahangir Karimi and Peter L.

Briggs. Software maintenance support for knowledge-based systems. The Journal of Systems and Software, 34(3):191– 210, September 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kang:1998:UDA

[KB98] Byung-Kyoo Kang and James M. Bieman. Using design abstractions to visualize, quantify, and restructure software. The Journal of Systems and Software, 42(2):175–187, August 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/2/6036.pdf.

Karimi:1996:PTC

[KC96] Jahangir Karimi and Craig Lee Carpenter. A partitioning technique for concurrent software design. The Journal of Systems and Software, 33(1):15–35, April 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Krovi:1998:UCR

[KC98] Ravindra Krovi and Akhilesh Chandra. User cognitive representations: The case for an object oriented model. The Journal of Systems and Software, 43(3): 165–176, November 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/3/6058.pdf.

Kusumoto:1998:PAT

[KCK^{+}98] Shinji Kusumoto, Atsushi Chimura, Tohru Kikuno, Ken ichi Matsumoto, and Yukio Mohri. A promising approach to twoperson software review in educational environment. The Journal of Systems and Software, 40 (2):115–123, February 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kavi:1991:SCP

[KD91] Krishna M. Kavi and Akshay K.

Deshpande. Specification of concurrent processes using a dataflow model of computation and partially ordered events. The Journal of Systems and Software, 16(2):107–120, October 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kendall:1980:DIC

[Ken80] Kenneth E. Kendall. A decentralized information and control system for blood management. The Journal of Systems and Software, 1(4):299–306, ???? 1980. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kent:1984:FBD

[Ken84] W. Kent. Fact-based data analysis and design. The Journal of Systems and Software, 4(2–3): 99–121, July 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kerr:1992:ESP

[Ker92] Roger M. Kerr. Expert systems in production scheduling: Lessons from a failed implementation. The Journal of Systems and Software, 19(2):123–130, October 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kung:1996:RTO

[KGH^{+}96] David C. Kung, Jerry Gao, Pei Hsia, Yasufumi Toyoshima, and Cris Chen. On regression testing of object-oriented programs. The Journal of Systems and Software, 32(1):21–40, January 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kafura:1981:SQM

[KH81] Dennis Kafura and Sallie Henry. Software quality metrics based on interconnectivity. The Journal of Systems and Software, 2 (2):121–131, June 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kan:1996:MCA

[KH96] Chieh-Ying Kan and Xudong He. A method for constructing algebraic Petri nets. The Journal of Systems and Software, 35(1):15– 27, October 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kee:1997:ECA

[KH97] Kok Kin Kee and Salim Hariri. Efficient communication algorithms for ring multicomputers. The Journal of Systems and Software, 37(3):215–225, June 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Katchabaw:1999:MDA

[KHL^{+}99] Michael J. Katchabaw, Stephen L.

Howard, Hanan L. Lutfiyya, Andrew D. Marshall, and Michael A. Bauer. Making distributed applications manageable through instrumentation. The Journal of Systems and Software, 45(2):81– 97, March 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/2/6097.pdf.

Klein:1999:UPE

[KJ99] Gary Klein and James J. Jiang.

User perception of expert system advice. The Journal of Systems and Software, 48(2):155– 161, October 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/37/23/23/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/23/abstract.
html.

Klein:1997:ISE

[KJB97] Gary Klein, James J. Jiang, and Joseph Balloun. Information system evaluation by system typology. The Journal of Systems and Software, 37(3):181– 186, June 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koch:1981:QSP

[KK81] Harvey S. Koch and Peter Kubat. Quick and simple procedures to assess software reliability and facilitate project management. The Journal of Systems and Software, 2(3):271–276, September 1981. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kaiser:1985:IPP

[KK85] Gail E. Kaiser and Elaine Kant. Incremental parsing without a parser. The Journal of Systems and Software, 5(2):121–144, May 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kusumoto:1996:EET

[KKiMT96] Shinji Kusumoto, Tohru Kikuno, Ken ichi Matsumoto, and Koji Torii. Experimental evaluation of time allocation procedure for technical reviews. The Journal of Systems and Software, 35(2): 119–126, November 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Korel:1990:DSC

[KL90] Bogdan Korel and Janusz Laski. Dynamic slicing of computer programs. The Journal of Systems and Software, 13(3):187– 195, November 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kramer:1991:TFM

[KL91] Bernd Krämer and Luqi. Toward formal models of software engineering processes. The Journal of Systems and Software, 15(1):

63–74, April 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khoshgoftaar:1995:NNA

[KL95] Taghi M. Khoshgoftaar and David L. Lanning. A neural network approach for early detection of program modules having high risk in the maintenance phase. The Journal of Systems and Software, 29(1):85–91, April 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kinnunen:1996:MTM

[KL96] Kimmo Kinnunen and Mauri Leppänen. O/A matrix and a technique for methodology engineering. The Journal of Systems and Software, 33(2):141– 152, May 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koutny:1989:SER

[KM89] M. Koutny and L. V. Mancini.

Synchronizing events in replicated systems. The Journal of Systems and Software, 9(3):183– 190, March 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khanna:1992:SVA

[KM92] Rashi Khanna and Bruce McMillin.

SMILI: Visualization of asynchronous massively parallel programs. The Journal of Systems and Software, 19(3):261–275, November 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khoshgoftaar:1994:AAU

[KML94] Taghi M. Khoshgoftaar, John C. Munson, and David L. Lanning. Alternative approaches for the use of metrics to order programs by complexity. The Journal of Systems and Software, 24 (3):211–221, March 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Keller-McNulty:1989:RRS

[KMM89] Sallie Keller-McNulty and Mark S. McNulty. Response to: Resolving the software science anomaly. The Journal of Systems and Software, 10(1):69–71, July 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Keller-McNulty:1991:SMS

[KMMG91] Sallie Keller-McNulty, Mark S. McNulty, and David A. Gustafson. Stochastic models for software science. The Journal of Systems and Software, 16(1):59–68, September 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kemayel:1991:CFP

[KMO91] L. Kemayel, A. Mili, and I. Ouederni. Controllable factors for programmer productivity: a statistical study. The Journal of Systems and Software, 16 (2):151–163, October 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kellner:1999:SPS

[KMR99] Marc I. Kellner, Raymond J.

Madachy, and David M. Raffo. Software process simulation modeling: Why? What? How? The Journal of Systems and Software, 46(2–3):91–105, April 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/46/2-3/
6118.pdf.

Kim:1997:SFT

[KN97] Hyun C. Kim and V. S. S. Nair.

Software fault tolerance for distributed object based computing. The Journal of Systems and Software, 39(2):103–117, November 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kulkarni:1986:MPR

[KNT86] V. G. Kulkarni, V. F. Nicola, and K. S. Trivedi. On modelling the performance and reliability of multimode computer systems. The Journal of Systems and Software, 6(1–2):175– 182, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kandelin:1995:VOO

[KO95] Nils A. Kandelin and Daniel E.

O’Leary. Verification of objectoriented systems: Domaindependent and domain-independent approaches. The Journal of Systems and Software, 29(3):261– 269, June 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Komorowski:1988:DLP

[Kom88] Henryk Jan Komorowski. A declarative logic programming environment. The Journal of Systems and Software, 8(2):77– 89, March 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kornman:1983:PMP

[Kor83] Brent D. Kornman. Pattern matching and pattern-directed invocation in systems programming languages. The Journal of Systems and Software, 3(1):95– 102, March 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koriem:1999:FSD

[Kor99a] Samir M. Koriem. Fast and simple decomposition techniques for the reliability analysis of interconnection networks. The Journal of Systems and Software, 45 (2):155–171, March 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1999/45/2/6102.pdf.

Koriem:1999:NPE

[Kor99b] Samir M. Koriem. R-nets for the performance evaluation of hard real-time systems. The Journal of Systems and Software, 46 (1):41–58, April 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/1/6114.pdf.

Koskimies:1991:HLT

[KP91] Kai Koskimies and Jukka Paakki. High-level tools for language implementation. The Journal of Systems and Software, 15(2):115–131, May 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koriem:1993:FTA

[KP93] Samir M. Koriem and L. M. Patnaik. Fault-tolerance analysis of hypercube systems using Petri net theory. The Journal of Systems and Software, 21(1):71–88, April 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kameas:1997:FAI

[KP97a] A. Kameas and P. Pintelas. The functional architecture and interaction model of a GENerator of intelligent TutORing applications. The Journal of Systems and Software, 36(3):233– 245, March 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koriem:1997:GSH

[KP97b] Samir M. Koriem and L. M. Patnaik. A generalized stochastic high-level Petri net model for performance analysis. The Journal of Systems and Software, 36 (3):247–265, March 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Konana:1998:TMM

[KR98] Prabhudev Konana and Sudha Ram. Transaction management mechanisms for active and realtime databases: a comprehensive protocol and a performance study. The Journal of Systems and Software, 42(3):205– 225, September 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/42/3/6039.pdf.

Kramer:1991:IGS

[Krä91a]  Bernd Krämer.  Introducing

the GRASPIN specification language SEGRAS. The Journal of Systems and Software, 15(1): 17–31, April 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kramer:1991:SPP

[Krä91b] Bernd Krämer. A sort of parametric polymorphism for algebraic specifications. The Journal of Systems and Software, 15(1): 33–42, April 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Krishnan:1993:SSI

[Kri93] Padmanabhan Krishnan. Specification of systems with interrupts. The Journal of Systems and Software, 21(3):291– 304, June 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ketabchi:1996:AOT

[KS96] Mohammad A. Ketabchi and Kamyar J. Sadeghi. Applying object technology to software analysis and maintenance system development. The Journal of Systems and Software, 32(1):41– 56, January 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kudo:1989:QDP

[KSFT89] Hideo Kudo, Yuji Sugiyama, Mamoru Fujii, and Koji Torii. Quantifying a design process based on experiments. The Journal of Systems and Software, 9 (2):129–136, February 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kelly:1992:ADD

[KSH92] John C. Kelly, Joseph S. Sherif, and Jonathan Hops. An analysis of defect densities found during software inspections. The Journal of Systems and Software, 17 (2):111–117, February 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Katz:1984:EVS

[KSS84] Randy Katz, Walt Scacchi, and P. Subrahmanyam. Environments for VLSI and software engineering. The Journal of Systems and Software, 4(1):13–26, April 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:1993:IOO

[KSW93] Won Kim, Yongmoo Suh, and Andrew B. Whinston. An IBIS and object-oriented approach to scientific research data management. The Journal of Systems and Software, 23(2):183– 197, November 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kitchenham:1985:SPD

[KT85] Barbara A. Kitchenham and N. R. Taylor. Software project development cost estimation. The Journal of Systems and Software, 5(4):267–278, November 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kung:1991:PIM

[Kun91a] Chenho Kung. Process interface modeling and consistency checking. The Journal of Systems and Software, 15(2):185– 191, May 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kung:1991:RDK

[Kun91b] Chenho Kung. Response to “Deliberations on Kung’s process interface modeling”. The Journal of Systems and Software, 15(2):199, May 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [CB91].

Kung:1995:EVF

[Kun95] David C. Kung. An executable visual formalism for objectoriented conceptual modeling. The Journal of Systems and Software, 31(1):33–43, October 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kuo:1994:MDE

[Kuo94] Feng-Yang Kuo. A methodology for deriving an entityrelationship model based on a data flow diagram. The Journal of Systems and Software, 24 (2):139–154, February 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kusalik:1990:SSC

[Kus90] Anthony J. Kusalik. Secondary storage in a concurrent logic programming environment. The Journal of Systems and Software, 11(1):31–44, January 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kumar:1991:TMD

[KW91] Muruganandan Kumar and Johnny Wong. Transaction models for design environments. The Journal of Systems and Software, 16(3):219–228, November 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kumar:1993:TMD

[KW93] Muruganandan Kumar and Johnny Wong. Transaction management in design databases. The Journal of Systems and Software, 22(1):3–15, July 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kalla:1999:ANR

[KWME99] Malleswar Kalla, Johnny S. K. Wong, Armin R. Mikler, and Stephen Elbert. Achieving nonrepudiation of Web based transactions. The Journal of Systems and Software, 48(3):165– 175, November 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/37/24/18/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/24/abstract.
html.

Kavi:1992:RTS

[KY92] Krishna M. Kavi and Seung-Min Yang. Real-time systems design methodologies: an introduction and a survey. The Journal of Systems and Software, 18(1):85– 99, April 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kamel:1991:MIH

[KZ91] Magdi N. Kamel and Moshe Zviran. A methodology for integrating heterogeneous data bases in a hospital environment. The Journal of Systems and Software, 15(3):251–260, July 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:1997:SEP

[LAHS97] Chi Y. Lin, Tarek Abdel-Hamid, and Joseph S. Sherif. Software-Engineering Process Simulation model (SEPS). The Journal of Systems and Software, 38(3): 263–277, September 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:1995:UPA

[Lai95] R. Lai. On using PROMPT for the automatic implementation of the ISO ACSE protocol. The Journal of Systems and Software, 28(2):143–155, February 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:1997:EUE

[Lai97a] R. Lai. An experience in using EDT to process the ISO transaction processing Estelle specification. The Journal of Systems and Software, 36(2):137– 145, February 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:1997:EE

[Lai97b] R. Lai. From Estelle to Erlang.

The Journal of Systems and Software, 37(3):235–251, June 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:1997:PRC

[Lai97c] R. Lai. Performance results for the CSMA/CD protocol using GreatSPN. The Journal of Systems and Software, 37(1):75–90, April 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:1997:SSV

[Lai97d] R. Lai. A success story of verifying a real complex protocol. The Journal of Systems and Software, 36(1):69–83, January 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:1999:TMI

[Lai99] R. Lai. Towards more industrially relevant academic researches into testing of communicating systems. The Journal of Systems and Software, 47(2–3):89–96, July 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

                                        /www.elsevier.com/cas/tree/
                                        store/jss/sub/1999/47/2-3/
                                        6143.pdf.

Lakhotia:1993:USE

[Lak93] Arun Lakhotia. Understanding someone else’s code: Analysis of experiences. The Journal of Systems and Software, 23(3): 269–275, December 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lakhotia:1997:UFE

[Lak97] Arun Lakhotia. A unified framework for expressing software subsystem classification techniques. The Journal of Systems and Software, 36(3):211– 231, March 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lam:1997:ARR

[Lam97] W. Lam. Achieving requirements reuse: a domain-specific approach from avionics. The Journal of Systems and Software, 38(3):197–209, September 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lanphar:1990:QPM

[Lan90] Robert Lanphar. Quantitative process management in software engineering, a reconciliation between process and product views. The Journal of Systems and Software, 12(3):243–248, July 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Land:1998:CBA

[Lan98a] Frank Land. A contingency based approach to requirements elicitation and systems development. The Journal of Systems and Software, 40(1):3–6, January 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Land:1998:IAO

[Lan98b] Frank Land. From IT application to organizational transformation. The Journal of Systems and Software, 40(2):183– 184, February 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Laski:1990:DFT

[Las90] Janusz Laski. Data flow testing in STAD. The Journal of Systems and Software, 12(1):3– 14, April 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lawrence:1981:PMO

[Law81] M. J. Lawrence. Programming methodology, organizational environment, and programming productivity. The Journal of Systems and Software, 2(3):257– 269, September 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lou:1998:AEM

[LC98] Der-Chyuan Lou and Chin-Chen Chang. An adaptive exponentiation method. The Journal of Systems and Software, 42 (1):59–69, July 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/1/6027.pdf.

Luqi:1998:SSP

[LCZ98] Luqi, Carl K. Chang, and Hong Zhu. Specifications in software prototyping. The Journal of Systems and Software, 42(2): 125–140, August 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/2/6032.pdf.

Lamb:1987:DPM

[LDN87] David Alex Lamb, Ivor Durham, and Joseph M. Newcomer. Drat: a program for maintaining listings. The Journal of Systems and Software, 7(2):163–170, June 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

LEcuyer:1987:FFR

[L’E87] Pierre L’Ecuyer. Formal formatting rules for Pascal programs. The Journal of Systems and Software, 7(4):311–322, December 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leary:1995:ABE

[Lea95] John R. Leary. An architectural basis for evolving software systems. The Journal of Systems and Software, 30(1–2):27– 43, July/August 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:1993:SMS

[Lee93] Heeseok Lee. A structured methodology for software development effort prediction using the analytic hierarchy process. The Journal of Systems and Software, 21(2):179–186, May 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lehman:1980:ULE

[Leh80] M. M. Lehman. On understanding laws, evolution, and conservation in the large-program life cycle. The Journal of Systems and Software, 1(3):213–221, ???? 1980. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lennartsson:1995:RC

[Len95] Bengt Lennartsson. Readers’ corner. The Journal of Systems and Software, 29(1):92, April 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Letovsky:1987:CPP

[Let87] Stanley Letovsky. Cognitive processes in program comprehension. The Journal of Systems and Software, 7(4):325–339, December 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leung:1992:OSR

[Leu92] Yiu-Wing Leung. Optimum software release time with a given cost budget. The Journal of Systems and Software, 17(3):233– 242, March 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leung:1997:DRA

[Leu97]  Yiu-Wing Leung.  Dynamic

resource-allocation for softwaremodule testing. The Journal of Systems and Software, 37(2): 129–139, May 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:1991:RTS

[LF91] Ching-Cheng Lee and H. A. Fatmi. Run-time support For parallel functional programming on shared-memory multiprocessors. The Journal of Systems and Software, 16(1):69–74, September 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lu:1996:VHS

[LF96] Jian Lu and Zongming Fei. Verification of HOS software specification by a hierarchical software understanding approach. The Journal of Systems and Software, 32(2):135–142, February 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

LeCharlier:1998:SNI

[LF98] Baudouin Le Charlier and Pierre Flener. Specifications are necessarily informal or: Some more myths of formal methods. The Journal of Systems and Software, 40(3):275–296, March 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liang:1999:FTO

[LFY^{+}99] D. Liang, C.-L. Fang, S.-M.

Yuan, C. Chen, and G. E. Jan. A fault–tolerant object service on CORBA. The Journal of Systems and Software, 48(3):197– 211, November 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/37/24/21/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/24/abstract.
html.

Lin:1997:GUI

[LG97] Yi-Bing Lin and Joe Geigel. A graphical user interface design for network simulation. The Journal of Systems and Software, 36(2):181–190, February 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leveson:1983:SFT

[LH83] Nancy G. Leveson and Peter R. Harvey. Software fault tree analysis. The Journal of Systems and Software, 3(2):173–181, June 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:1990:MTS

[LH90] J. A. N. Lee and Xudong He. A methodology for test selection. The Journal of Systems and Software, 13(3):177–185, November 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:1993:OOM

[LH93] Wei Li and Sallie Henry. Objectoriented metrics that predict maintainability. The Journal of Systems and Software, 23(2): 111–122, November 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:1995:MVR

[LH95] Jane W. S. Liu and Rhan Ha. Methods for validating real-time constraints. The Journal of Systems and Software, 30(1–2):85– 98, July/August 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lattanzi:1998:SRU

[LH98] Mark Lattanzi and Sallie Henry.

Software reuse using C++ classes: The question of inheritance. The Journal of Systems and Software, 41(2):127– 132, May 1, 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/2/6013.pdf.

Lam:1995:MDA

[LHC95] Kwok-Yan Lam, Lucas Hui, and Siu-Leung Chung. Multivariate data analysis software for enhancing system security. The Journal of Systems and Software, 31(3):267–275, December 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lam:1996:DRM

[LHC96] Kwok-Yan Lam, Lucas Hui, and Siu-Leung Chung. A data reduction method for intrusion detection. The Journal of Systems and Software, 33(1):101– 108, April 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:1998:AMS

[Li98] Wei Li. Another metric suite for object-oriented programming. The Journal of Systems and Software, 44(2):155– 162, December 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/2/6080.pdf. See corri-

gendum [Li99].

Li:1999:CAM

[Li99] Wei Li. Corrigendum to “Another metric suite for objectoriented programming” [Journal of Systems and Software 44 (1998) 155–162]. The Journal of Systems and Software, 48(1):75, August 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.elsevier.nl/gej-ng/10/
29/11/37/17/24/;   http://
www.elsevier.nl/gej-ng/10/
29/11/37/17/24/article.pdf.

See [Li98].

Leventhal:1992:AVC

[LIC92] Laura M. Leventhal, Keith E. Instone, and David W. Chilson. Another view of computer science ethics: Patterns of responses among computer scientists. The Journal of Systems and Software, 17(1):49–60, January 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Linberg:1999:SDP

[Lin99] Kurt R. Linberg. Software developer perceptions about software project failure: a case study. The Journal of Systems and Software, 49(2–3):177–192, December 30, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

elsevier.nl/gej-ng/10/29/
11/46/26/34/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

Lipow:1979:PSF

[Lip79] Myron Lipow. Prediction of software failures. The Journal of Systems and Software, 1(1):71– 75, ???? 1979. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Littlewood:1980:LVM

[Lit80] B. Littlewood. The Littlewood– Verrall model for software reliability compared with some rivals. The Journal of Systems and Software, 1(3):251–258, ???? 1980. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Litecky:1990:SMP

[Lit90]  Chuck Litecky.  Spreadsheet

macro programming: a critique with emphasis on Lotus 1-2-3. The Journal of Systems and Software, 13(3):197–200, November 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:1993:FRS

[Liu93] Shaoying Liu. A formal requirements specification method based on data flow analysis. The Journal of Systems and Software, 21(2):141–149, May 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:1995:ICF

[Liu95] Shaoying Liu. Internal consistency of FRSM specifications. The Journal of Systems and Software, 29(2):167–175, May 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:1998:QAA

[Liu98] Xiaoqing Frank Liu. A quantitative approach for assessing the priorities of software quality requirements. The Journal of Systems and Software, 42(2): 105–113, August 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/2/6030.pdf.

Lloyd:1999:CIP

[LJ99] Karl Brett Lloyd and David John Jankowski. A cognitive information processing and information theory approach to diagram clarity: a synthesis and experimental investigation. The Journal of Systems and Software, 45(3): 203–214, March 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/3/6106.pdf.

Lopez:1996:GFD

[LJM96] Antonio M. Lopez, Jr., Bobby Junkin, and Lloyd McGregor. Gross feature detection in remotely sensed CAMS imagery using a multi-paradigm software approach. The Journal of Systems and Software, 34(2):151– 159, August 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lloyd:1993:TED

[LK93] William S. Lloyd and Phil Kearns. Tracing the execution of distributed programs. The Journal of Systems and Software, 21(3):201–214, June 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Laughery:1985:HFS

[LL85] K. Ronald Laughery, Jr. and Kenneth R. Laughery, Sr. Human factors in software engineering: a review of the literature. The Journal of Systems and Software, 5(1):3–14, February 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:1997:VIA

[LL97a] R. Lai and X. Li. Verification of the ISO ACSE protocol using deductive machine. The Journal of Systems and Software, 39(3): 223–247, December 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:1997:PNB

[LL97b] Gang-Soo Lee and Jin-Seok Lee. Petri net based models for specification and analysis of cryptographic protocols. The Journal of Systems and Software, 37(2): 141–159, May 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:1998:AND

[LL98] R. Lai and X. Li. Applying a new decomposition method to verify communication protocols. The Journal of Systems and Software, 40(1):29–50, January 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lo:1999:AII

[LL99] A. Lo and R. Lai. An automatic implementation of the ISO FTAM protocol based on an integrated specification of Estelle and ASN.1. The Journal of Systems and Software, 48(2):119– 137, October 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/37/23/21/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/23/abstract.
html.

Lee:1994:DSP

[LM94] Gang-Soo Lee and Tadao Murata. A β-distributed stochastic Petri net model for software project time/cost management. The Journal of Systems and Software, 26(2):149– 165, August 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:1996:MOA

[LM96] Shaoying Liu and John A. Mc-Dermid. A model-oriented approach to safety analysis using fault trees and a support system. The Journal of Systems and Software, 35(2):151–164, November 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Licker:1992:DES

[LO92] Paul Licker and Bruce Olsen. Dimensions of expert system management. The Journal of Systems and Software, 19(2):171– 183, October 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lohse:1984:EES

[Loh84] John B. Lohse. Experimental evaluation of software design principles: an investigation into the effect of module coupling on system modifiability. The Journal of Systems and Software, 4(4):301–308, November 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lokan:1996:ESP

[Lok96] Christopher J. Lokan. Early size prediction for C and Pascal programs. The Journal of Systems and Software, 32(1):65–72, January 1996. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:1993:OOO

[LP93] Yang Kyu Lee and Sung Joo Park. OPNets: an objectoriented high-level Petri net model for real-time system modeling. The Journal of Systems and Software, 20(1):69–86, January 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lederer:1995:CIS

[LP95] Albert L. Lederer and Jayesh Prasad. Causes of inaccurate software development cost estimates. The Journal of Systems and Software, 31(2):125– 134, November 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Littman:1987:MMS

[LPLS87] David C. Littman, Jeannine Pinto, Stanley Letovsky, and Elliot Soloway. Mental models and software maintenance. The Journal of Systems and Software, 7(4):341–355, December 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lehman:1999:IFG

[LR99] M. M. Lehman and J. F. Ramil. The impact of feedback in the global software process. The Journal of Systems and Software, 46(2–3):123–134, April 15, 1999. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/46/2-3/
6121.pdf.

Lanovaz:1992:OOI

[LS92] Daniel Lanovaz and Duane Szafron. An object-oriented inference engine for PROLOG. The Journal of Systems and Software, 19(1):13–25, September 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lam:1997:IHA

[LS97] Kwok-Yan Lam and Tom Salkield.

Implementing a highly available network directory service. The Journal of Systems and Software, 37(1):41–47, April 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lindvall:1998:HWD

[LS98] Mikael Lindvall and Kristian Sandahl. How well do experienced software developers predict software change? The Journal of Systems and Software, 43 (1):19–27, October 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/1/6047.pdf.

Lidtke:1999:ISC

[LS99] Doris K. Lidtke and Gorden Stokes. An information systems-centric curriculum, ISCC ’99. The Journal of Systems and Software, 49(2–3):171– 175, December 30, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/26/33/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

Liu:1995:PFM

[LSD95] Shaoying Liu, Victoria Stavridou, and Bruno Dutertre. The practice of formal methods in safety-critical systems. The Journal of Systems and Software, 28(1):77–87, January 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Laitinen:1997:EMS

[LTHR97] Kari Laitinen, Jorma Taramaa, Markku Heikkila, and Neil C. Rowe. Enhancing maintainability of source programs through disabbreviation. The Journal of Systems and Software, 37(2): 117–128, May 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:1992:IES

[LTT92] Jim-Min Lin, Shang Rong Tsai, and Li-Ming Tseng. Integrating existing software packages using the virtual machine technique. The Journal of Systems and Software, 18(3):207–218, July 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Luegenbiehl:1992:CPM

[Lue92] Heinz C. Luegenbiehl. Computer professionals: Moral autonomy and a code of ethics. The Journal of Systems and Software, 17 (1):61–68, January 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lutz:1996:TSR

[Lut96] Robyn R. Lutz. Targeting safetyrelated errors during software requirements analysis. The Journal of Systems and Software, 34(3): 223–230, September 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lanubile:1997:EPQ

[LV97] Filippo Lanubile and Giuseppe Visaggio. Evaluating predictive quality models derived from software measures: Lessons learned. The Journal of Systems and Software, 38(3):225–234, September 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Luckham:1993:POE

[LVB^{+}93] David C. Luckham, James Vera, Doug Bryan, Larry Augustin, and Frank Belz. Partial orderings of event sets and their application to prototyping concurrent, timed systems. The Journal of Systems and Software, 21(3): 253–265, June 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lassez:1981:CES

[LvSL81] J. L. Lassez, D. van der Knijff, J. Shepherd, and C. Lassez. A critical examination of software science. The Journal of Systems and Software, 2(2):105– 112, June 1981. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lei:1997:EDP

[LZL97] Su Te Lei, Kang Zhang, and Kei-Chun Li. Experience with the design of a performance tuning tool for parallel programs. The Journal of Systems and Software, 39(1):27–37, October 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McGarry:1989:MAS

[MA89] Frank E. McGarry and William W. Agresti. Measuring Ada for software development in the Software Engineering Laboratory. The Journal of Systems and Software, 9(2):149–159, February 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Medeiros:1994:IIC

[MA94] Claudia Bauzer Medeiros and Márcia Jacobina Andrade. Implementing integrity control in active data bases. The Journal of Systems and Software, 27(3): 171–181, December 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

MacDonell:1991:RSC

[Mac91] S. G. MacDonell. Rigor in software complexity measurement experimentation. The Journal of Systems and Software, 16(2): 141–149, October 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Maier:1996:IMU

[Mai96] Mark W. Maier. Integrated modeling: a unified approach to system engineering. The Journal of Systems and Software, 32(2): 101–119, February 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Martin:1981:ICP

[Mar81] Edith W. Martin. Introduction to collected papers on computing standards for embedded computer systems. The Journal of Systems and Software, 2(4): 317–318, December 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Marti:1984:IDP

[Mar84] Robert W. Marti. Integrating database and program descriptions using an er-data dictionary. The Journal of Systems and Software, 4(2–3):185–195, July 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Matley:1986:MPC

[Mat86] Ben G. Matley. Modest proposal: Computer education can be raised to professional status. The Journal of Systems and Software, 6(4):379–389, November 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mathews:1996:OFO

[Mat96] John Mathews. Organizational foundations of object-oriented programming. The Journal of Systems and Software, 34(3): 247–253, September 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mazlack:1981:NLS

[Maz81] Lawrence J. Mazlack. Natural language symbol string storage. The Journal of Systems and Software, 2(1):71–79, February 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mostow:1984:ATS

[MB84] Jack Mostow and Bob Balzer. Application of a transformational software development methodology to VLSI design. The Journal of Systems and Software, 4(1):51–61, April 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mashiko:1997:UGP

[MB97] Yasuhiro Mashiko and Victor R. Basili. Using the GQM paradigm to investigate influential factors for software process improvement. The Journal of Systems and Software, 36(1):17–32, January 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Marsan:1986:PFS

[MBCD86] Marco Ajmone Marsan, Gianfranco Balbo, Giovanni Chiola, and Susanna Donatelli. On the product-form solution of a class of multiple-bus multiprocessor system models. The Journal of Systems and Software, 6 (1–2):117–124, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mead:1999:IUC

[MBL^{+}99] Nancy Mead, Kathy Beckman, Jimmy Lawrence, George O’Mary, Cynthia Parish, Perla Unpingco, and Hope Walker. Industry/university collaborations: different perspectives heighten mutual opportunities. The Journal of Systems and Software, 49(2–3):155–162, December 30, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

elsevier.nl/gej-ng/10/29/
11/46/26/31/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

Mathaisel:1991:CCS

[MC91] Dennis F. X. Mathaisel and Clare L. Comm. Course and classroom scheduling: An interactive computer graphics approach. The Journal of Systems and Software, 15(2):149– 157, May 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Matocha:1998:TDT

[MC98] Jeff Matocha and Tracy Camp. A taxonomy of distributed termination detection algorithms. The Journal of Systems and Software, 43(3):207–221, November 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/43/3/6062.pdf.

McFarland:1992:BBE

[McF92] Michael C. McFarland. Broadening the basis for ethical analysis. The Journal of Systems and Software, 17(1):39–40, January 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Millen:1981:EAH

[MD81] Jonathan K. Millen and David L. Drake. Experiment with Affirm and Hdm. The Journal of Systems and Software, 2(2):159–175, June 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Moreau:1989:OOG

[MD89] Dennis R. Moreau and Wayne D. Dominick. Object-oriented graphical information systems: Research plan and evaluation metrics. The Journal of Systems and Software, 10(1):23–28, July 1989. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Mincer-Daszkiewicz:1991:PBM

[MD91] Janina Mincer-Daszkiewicz. Program I/O behavior: Models and their applications. The Journal of Systems and Software, 14 (1):51–62, January 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Merriman:1987:AIS

[Mer87] Michael F. Merriman. Automated interactive simulation modeling system: Aisim. The Journal of Systems and Software, 7(1):61–72, March 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Meyer:1988:CTC

[Mey88a] Bertrand Meyer. Cépage: Toward computer-aided design of software. The Journal of Systems and Software, 8(5):419–429, December 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Meyer:1988:ELE

[Mey88b] Bertrand Meyer. Eiffel: a language and environment for software engineering. The Journal of Systems and Software, 8(3): 199–246, June 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Marlin:1990:CCT

[MF90] Chris Marlin and Dennis Freidel. Comparing communication in two languages employing buffered message-passing. The Journal of Systems and Software, 12(2):87–105, May 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McMullin:1981:EDA

[MG81] Paul R. McMullin and John D. Gannon. Evaluating a data abstraction testing system based on formal specifications. The Journal of Systems and Software, 2 (2):177–186, June 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Matzen:1997:FLM

[MGH97] R. W. Matzen, K. M. George, and G. E. Hedrick. A formal language model for parsing SGML. The Journal of Systems and Software, 36(2):147– 166, February 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mili:1987:UEA

[MGJT87] Ali Mili, Sihem Guemara, Ali Jaoua, and Paul Torres. On the use of executable assertions in structured programs. The Journal of Systems and Software, 7 (1):15–28, March 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Moser:1999:CEB

[MHSM99] Simon Moser, Brian Henderson-Sellers, and Vojislav B. Mišić. Cost estimation based on business models. The Journal of Systems and Software, 49(1):33– 42, December 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/25/28/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/25/abstract.
html.

Miller:1992:PMP

[MIH92] L. L. Miller, S. R. Inglett, and A. R. Hurson. PASS — a multiuser parallel file system based on microcomputers. The Journal of Systems and Software, 19 (1):75–83, September 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mills:1989:MSE

[Mil89] Everald E. Mills. The Master of Software Engineering (MSE) program at Seattle University. The Journal of Systems and Software, 10(4):245–252, November 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mili:1996:BDO

[Mil96a] Hafedh Mili. On behavioral descriptions in object-oriented modeling. The Journal of Systems and Software, 34(2):105– 121, August 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mills:1996:EES

[Mil96b] Kevin L. Mills. An experimental evaluation of specification techniques for improving functional testing. The Journal of Systems and Software, 32(1):83–95, January 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Millet:1998:NF

[Mil98] Ido Millet. Normalization on the fly. The Journal of Systems and Software, 40(1):85–88, January 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McCrosky:1989:DPA

[MJ89] Carl McCrosky and M. A. Jenkins. Design of a processor for array-theoretic computation. The Journal of Systems and Software, 9(1):67–74, January 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Munson:1990:ARC

[MK90] John C. Munson and Taghi M. Khoshgoftaar. Applications of a relative complexity metric for software project management. The Journal of Systems and Software, 12(3):283–291, July 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Munson:1993:MDS

[MK93] John C. Munson and Taghi M. Kohshgoftaar. Measurement of data structure complexity. The Journal of Systems and Software, 20(3):217–225, March 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mutka:1995:TAP

[ML95] Matt W. Mutka and Jong-Pyng Li. A tool for allocating periodic real-time tasks to a set of processors. The Journal of Systems and Software, 29(2):135– 148, May 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

MacEwen:1981:AHT

[MM81] Glenn H. MacEwen and T. Patrick Martin. Abstraction hierarchies in top-down design. The Journal of Systems and Software, 2(3): 213–224, September 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McColl:1992:EEN

[MM92] R. Beth McColl and James C.

McKim Jr. Evaluating and extending NPath as a software complexity measure. The Journal of Systems and Software, 17 (3):275–279, March 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Masiero:1993:DIG

[MM93a] P. C. Masiero and C. A. A.

Meira. Development and instantiation of a generic application generator. The Journal of Systems and Software, 23(1):27– 37, October 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McKim:1993:CID

[MM93b] James C. McKim, Jr. and David A. Mondou. Class interface design: Designing for correctness. The Journal of Systems and Software, 23(2):85–94, November 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Morell:1993:SMT

[MM93c] Larry Morell and Branson Murrill. Semantic metrics through error flow analysis. The Journal of Systems and Software, 20 (3):253–265, March 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mackey:1995:SMR

[MM95] Stephen R. Mackey and Lynn M. Meredith. Software migration and reengineering: a pilot project in reengineering. The Journal of Systems and Software, 30(1–2):137–150, July/ August 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Morrison:1992:EST

[MMSH92] Mike Morrison, Joline Morrison, Olivia R. Liu Sheng, and Kunihiko Higa. Environment selection, training, and implementation in high-level expert systems environments: Experiences and guidelines. The Journal of Systems and Software, 19(2):147– 152, October 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Musa:1984:CTD

[MO84] John D. Musa and Kazuhira Okumoto. A comparison of time domains for software reliability models. The Journal of Systems and Software, 4(4):277–287, November 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mizuno:1990:ACL

[MO90] Masaaki Mizuno and Arthur E.

Oldehoeft. An access control language for object-oriented programming systems. The Journal of Systems and Software, 13 (1):3–12, September 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mohanty:1981:EMS

[Moh81] Siba N. Mohanty. Entropy metrics for software design evaluation. The Journal of Systems and Software, 2(1):39–46, February 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Moores:1998:ACM

[Moo98]  Trevor T. Moores.  Apply-

ing complexity measures to rule-based Prolog programs. The Journal of Systems and Software, 44(1):45–52, December 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/1/6070.pdf.

Morganti:1986:CDF

[Mor86] M. Morganti. Communications in distributed fault-tolerant computing systems. The Journal of Systems and Software, 6(1–2): 213–216, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Morisio:1999:MPS

[Mor99] M. Morisio. Measurement processes are software, too. The Journal of Systems and Software, 49(1):17–31, December 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic).  URL http://www.
                                        elsevier.nl/gej-ng/10/29/
                                        11/46/25/27/article.pdf;
                                        http://www.elsevier.nl/gej-
                                        ng/10/29/11/46/25/abstract.
                                        html.

Mostow:1984:DBF

[Mos84a] Jack Mostow. A decision-based framework for comparing hardware compilers. The Journal of Systems and Software, 4(1):39– 50, April 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mostow:1984:DFC

[Mos84b] Jack Mostow. Decision-based framework for comparing hardware compilers. The Journal of Systems and Software, 4(1):153– 164, April 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Motschnig-Pitrik:1996:ANA

[Mot96] Renate Motschnig-Pitrik. Analyzing the notions of attribute, aggregate, part and member in data/knowledge modeling. The Journal of Systems and Software, 33(2):113–122, May 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Moynihan:1996:ECO

[Moy96] Tony Moynihan. An experimental comparison of objectorientation and functional-decomposition as paradigms for communicating system functionality to users. The Journal of Systems and Software, 33(2):163–169, May 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Moore:1989:TPS

[MP89] Freeman L. Moore and Phillip R.

Purvis. Training practicing software engineers at Texas Instruments. The Journal of Systems and Software, 10(4):253– 260, November 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Motschnig-Pitrik:1990:FSC

[MP90] R. Motschnig-Pitrik. A framework for the support of a common structural level for software- , data base-, and knowledgebased systems. The Journal of Systems and Software, 12(2): 125–137, May 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Makki:1994:NSO

[MP94] Kia Makki and Niki Pissinou.

A new storage organization for temporal data bases. The Journal of Systems and Software, 27(3):159–169, December 1994.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Murrell:1995:FSR

[MP95] S. Murrell and R. Plant. Formal semantics for rule-based systems. The Journal of Systems and Software, 29(3):251–259, June 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Miellou:1986:IMP

[MPS86] J.-C. Miellou, G.-R. Perrin, and P. Spiteri. Inexpensive method of performance evaluation for subdomain decomposition parallel algorithms of three-dimensional elliptical problems. The Journal of Systems and Software, 6(1–2): 169–173, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McHenry:1980:STI

[MR80] Robert C. McHenry and Jerry A. Rand. Software technology and integration engineering. The Journal of Systems and Software, 1(3):187–193, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mili:1983:RMI

[MR83] Ali Mili and Donna Reese. Representation and manipulation of information systems: a simplified approach. The Journal of Systems and Software, 3(1):85– 93, March 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Markowitz:1984:ERA

[MR84] Victor M. Markowitz and Yoav Raz. Entity-relationship algebra and its semantic description capabilities. The Journal of Systems and Software, 4(2–3):147– 162, July 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Marie:1986:AMM

[MR86] Raymond Marie and Gerardo Rubino. Approximation for a multiclass. /M/1/Fifo queue imbedded in a closed queueing network. The Journal of Systems and Software, 6(1–2):31–39, May 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Misic:1999:ASA

[MR99] Mark M. Misic and Nancy L.

Russo. An assessment of systems analysis and design courses. The Journal of Systems and Software, 45(3):197– 202, March 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/3/6105.pdf.

Mili:1994:PSC

[MRW^{+}94] Hafedh Mili, Roy Radai, Wang Weigang, Karl Strickland, Cornelia Boldyreff, Lene Olsen, Jan Witt, Jurgen Heger, Wolfgang Scherr, and Peter Elzer. Practitioner and SoftClass: a comparative study of two software reuse research projects. The Journal of Systems and Software, 25(2): 147–170, May 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Manley:1979:EIa

[MS79a] John H. Manley and Alan B. Salisbury. Editors’ introduction. The Journal of Systems and Software, 1(1):1, ???? 1979. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Manley:1979:EIb

[MS79b] John H. Manley and Alan B. Salisbury. Editors’ introduction. The Journal of Systems and Software, 1(2):105, ???? 1979. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Majoros:1981:SPT

[MS81] Maria Majoros and Harry M. Sneed. Softest program test system. The Journal of Systems and Software, 2(4):289–296, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mishra:1990:FTM

[MS90] Shivakant Mishra and Pradip K. Srimani. Fault-tolerant mutual exclusion algorithms. The Journal of Systems and Software, 11 (2):111–129, February 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mehta:1997:MTA

[MS97] Dinesh P. Mehta and Sartaj Sahni. Models, techniques, and algorithms for finding, selecting, and displaying patterns in strings and other discrete objects. The Journal of Systems and Software, 39(3):201–221, December 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Morrey:1998:TSC

[MSHB98] Ian Morrey, Jawed Siddiqi, Richard Hibberd, and Graham Buckberry. A toolset to support the construction and animation of formal specifications. The Journal of Systems and Software, 41(3):147– 160, June 1, 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/3/6016.pdf.

Mili:1990:OOM

[MSI90] Hafedh Mili, John Sibert, and Yoav Intrator. An objectoriented model based on relations. The Journal of Systems and Software, 12(2):139– 155, May 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Misic:1998:EEC

[MT98] Vojislav B. Mišić and Dejan N.

Tešić. Estimation of effort and complexity: an object-oriented case study. The Journal of Systems and Software, 41(2): 133–143, May 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/41/2/6014.pdf.

Mata-Toledo:1992:FAS

[MTG92] Ramon A. Mata-Toledo and David A. Gustafson. A factor analysis of software complexity measures. The Journal of Systems and Software, 17(3):267– 273, March 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Miyazaki:1994:RRD

[MTON94] Y. Miyazaki, M. Terakado, K. Ozaki, and H. Nozaki. Robust regression for developing software estimation models. The Journal of Systems and Software, 27(1):3–16, October 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mata-Toledo:1997:VRS

[MTW97] Ramón A. Mata-Toledo and Matthew A. Willis. Visualization of random sequences using the chaos game algorithm. The Journal of Systems and Software, 39(1):3–6, October 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mueller:1986:DAS

[Mue86] Bruno Mueller. Decomposition approach for the stationary analysis of fault tolerant queueing systems. The Journal of Systems and Software, 6(1–2):199– 204, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Murphy:1999:TSP

[Mur99] Michael G. Murphy. Teaching software project management: a response–interaction approach. The Journal of Systems and Software, 49(2–3):145– 148, December 30, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/26/29/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

Musa:1980:SRM

[Mus80] John D. Musa. Software reliability measurement. The Journal of Systems and Software, 1(3):223– 241, ???? 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Morell:1993:FDS

[MV93] Larry J. Morell and Jeffrey M.

Voas. A framework for defining semantic metrics. The Journal of Systems and Software, 20 (3):245–251, March 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mostert:1995:TIC

[MvS95] D. N. J. Mostert and S. H. von Solms. A technique to include computer security, safety, and resilience requirements as part of the requirements specification. The Journal of Systems and Software, 31(1):45–53, October 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mehrotra:1995:AKB

[MW95] Mala Mehrotra and Chris Wild. Analyzing knowledge-based systems with multiviewpoint clustering analysis. The Journal of Systems and Software, 29(3): 235–249, June 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mikler:1997:QVF

[MWH97] Armin R. Mikler, Johnny S. K. Wong, and Vasant Honavar. Quo vadis — a framework for intelligent routing in large communication networks. The Journal of Systems and Software, 37(1): 61–73, April 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mikler:1998:OOA

[MWH98] Armin R. Mikler, Johnny S. K. Wong, and Vasant Honavar. An object oriented approach to simulating large communication networks. The Journal of Systems and Software, 40(2):151– 164, February 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Myrvold:1990:DAS

[Myr90] Alan Myrvold. Data analysis for software metrics. The Journal of Systems and Software, 12(3): 271–275, July 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Navabi:1992:HLL

[Nav92] Zainalabedin Navabi. A highlevel language for design and modeling of hardware. The Journal of Systems and Software, 18(1):5–18, April 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nusenoff:1993:GST

[NB93] Ronald E. Nusenoff and Dennis C. Bunde. A guidebook and a spreadsheet tool for a corporate metrics program. The Journal of Systems and Software, 23(3): 245–255, December 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Norcio:1988:DCS

[NC88] A. F. Norcio and L. J. Chmura.

Designing complex software. The Journal of Systems and Software, 8(3):165–184, June 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nesi:1996:MFO

[NC96] P. Nesi and M. Campanai. Metric framework for object-oriented real-time systems specification languages. The Journal of Systems and Software, 34(1):43–65, July 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Naumann:1980:DIR

[NDM80] J. David Naumann, Gordon B.

Davis, and James D. McKeen. Determining information requirements: a contingency method for selection of a requirements assurance strategy. The Journal of Systems and Software, 1(4):273–281, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nechvatal:1996:PKB

[Nec96] James Nechvatal. A public-key-based key escrow system. The Journal of Systems and Software, 35(1):73–83, October 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Neilsen:1997:PNK

[Nei97] Mitchell L. Neilsen. Properties of nondominated K-coteries. The Journal of Systems and Software, 37(1):91–96, April 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nelson:1981:FPA

[Nel81] Eldred Nelson. Functional programming analysis. The Journal of Systems and Software, 2(3): 225–235, September 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Narayanaswamy:1991:FFC

[NG91] K. Narayanaswamy and Neil M. Goldman. A flexible framework for cooperative distributed software development. The Journal of Systems and Software, 16 (2):97–105, October 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ng:1993:VOO

[Ng93] K. W. Ng. A visual objectoriented concurrent LISP multiparadigm environment. The Journal of Systems and Software, 22(2):91–106, August 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ng:1999:RBV

[Ng99] Joseph Kee-Yin Ng. A reserved bandwidth video smoothing algorithm for MPEG transmission. The Journal of Systems and Software, 48(3):233– 245, November 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/37/24/23/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/24/abstract.
html.

Ni:1997:ETT

[Ni97] David Chi-Liang Ni. Enumeration and traceability tools for UNIX^{T M} and WINDOWS^{T M} environments. The Journal of Systems and Software, 39(1):15– 25, October 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nitsche:1996:VBA

[Nit96] Ulrich Nitsche. Verification and behavior abstraction towards a tractable verification technique for large distributed systems. The Journal of Systems and Software, 33(3):273–285, June 1996.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nitsche:1998:AFV

[Nit98] Ulrich Nitsche. Application of formal verification and behaviour abstraction to the service interaction problem in intelligent networks. The Journal of Systems and Software, 40(3):227– 248, March 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Naslund:1999:UIC

[NL99] Torbjörn Näslund and Jonas Löwgren. Usability inspection in contract-based systems development — a contextual assessment. The Journal of Systems and Software, 45(3):233– 240, March 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1999/45/3/6109.pdf.

Neilsen:1993:QBA

[NM93] Mitchell L. Neilsen and Masaaki Mizuno. A quorum-based algorithm for parameterized semaphore operations. The Journal of Systems and Software, 21 (1):103–112, April 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Notkin:1985:ABG

[Not85a] David Notkin. Annotated bibliography of Gandalf literature.

The Journal of Systems and Software, 5(2):173–176, May 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Notkin:1985:GP

[Not85b] David Notkin. Gandalf Project.

The Journal of Systems and Software, 5(2):91–105, May 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nesi:1998:EEP

[NQ98] P. Nesi and T. Querci. Effort estimation and prediction of object-oriented systems. The Journal of Systems and Software, 42(1):89–102, July 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/1/6021.pdf.

Nyari:1983:SPA

[NS83] Erika Nyari and Harry Sneed.

Sofspec: a pragmatic approach to automated specification verification. The Journal of Systems and Software, 3(3):193–200, September 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Narayanaswamy:1987:DFS

[NS87] K. Narayanaswamy and Walt Scacchi. Database foundation to support software system evolution. The Journal of Systems and Software, 7(1):37–49, March 1987. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Naik:1992:VPC

[NS92] Kshirasagar Naik and Behcet Sarikaya. Verification of protocol conformance test cases using reachability analysis. The Journal of Systems and Software, 19 (1):41–57, September 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nakata:1984:IED

[NY84] Shuji Nakata and Go Yamazaki. Ismos: an experimental database-oriented tool generator. The Journal of Systems and Software, 4(2–3):219–238, July 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Osterweil:1979:ALC

[OBS79] Leon J. Osterweil, John R. Brown, and Leon G. Stucki. AS-SET: a life cycle verification and visibility system. The Journal of Systems and Software, 1(1): 77–86, ???? 1979. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oman:1990:DCT

[OC90] Paul W. Oman and Curtis R. Cook. Design and code traceability using a PDL metrics tool. The Journal of Systems and Software, 12(3):189–198, July 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oman:1991:PST

[OC91] Paul W. Oman and Curtis R.

Cook. A programming style taxonomy. The Journal of Systems and Software, 15(3):287– 301, July 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oman:1989:EPE

[OCCN89] Paul W. Oman, Curtis, R. Cook, and Murthi Nanja. Effects of programming experience in debugging semantic errors. The Journal of Systems and Software, 9(3):197–207, March 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Okumoto:1980:ORT

[OG80] Kazu Okumoto and Amrit L.

Goel. Optimum release time for software systems based on reliability and cost criteria. The Journal of Systems and Software, 1 (4):315–318, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oman:1994:CTP

[OH94] Paul Oman and Jack Hagemeister. Construction and testing of polynomials predicting software maintainability. The Journal of Systems and Software, 24 (3):251–266, March 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Odani:1990:HBS

[OHBR90] Masayasu Odani, Sun Young Hwang, Tom Blank, and Tom Rokicki. The Hermod behavioral synthesis system. The Journal of Systems and Software, 13(1):39– 54, September 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Offutt:1993:SMS

[OHK93] A. Jefferson Offutt, Mary Jean Harrold, and Priyadarshan Kolte. A software metric system for module coupling. The Journal of Systems and Software, 20 (3):295–308, March 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Orailoglu:1994:SFT

[OK94] Alex Orailoğlu and Ramesh Karri. Synthesis of faulttolerant and real-time microarchitectures. The Journal of Systems and Software, 25(1):73–84, April 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Obara:1997:MAT

[OKOM97] E. Obara, T. Kawasaki, Y. Ookawa, and N. Maeda. Metrics and analyses in the test phase of largescale software. The Journal of Systems and Software, 38(1):37– 46, July 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Offutt:1999:GTD

[OL99] A. Jefferson Offutt and Shaoying Liu. Generating test data from SOFL specifications. The Journal of Systems and Software, 49(1):49–62, December 15, 1999.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/25/30/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/25/abstract.
html.

ONeill:1983:IEP

[O’N83] Don O’Neill. Integration engineering perspective. The Journal of Systems and Software, 3(1): 77–83, March 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oravec:1992:GEI

[OP92] Jo Ann Oravec and Robert Plant. Guest Editors’ introduction. The Journal of Systems and Software, 19(2):111, October 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Onorato:1987:PND

[OS87] Lisa A. Onorato and Roger W.

Schvaneveldt. Programmernonprogrammer differences in specifying procedures to people and computers. The Journal of Systems and Software, 7(4): 357–369, December 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

OHara-Schettino:1998:DNM

[OSG98] Elizabeth O’Hara-Schettino and Hassan Gomaa. Dynamic navigation in multiple view software specifications and designs. The Journal of Systems and Software, 41(2):93–103, May 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/2/6010.pdf.

Ostroff:1992:FMS

[Ost92] Jonathan S. Ostroff. Formal methods for the specification and design of real-time safety critical systems. The Journal of Systems and Software, 18(1):33–60, April 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.cs.yorku.
ca/People/Jonathan.Ostroff/
survey/combined-paper/combined-
paper.html.

Oravec:1992:IWC

[OT92] Jo Ann Oravec and Larry Travis. If we could do it over, we’d … learning from lessthan-successful expert system projects. The Journal of Systems and Software, 19(2):113– 122, October 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ostrand:1984:CCS

[OW84] Thomas J. Ostrand and Elaine J. Weyuker. Collecting and categorizing software error data in an industrial environment. The Journal of Systems and Software, 4(4):289–300, November 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Owolabi:1996:DOE

[Owo96] O. Owolabi. Dictionary organizations for efficient similarity retrieval. The Journal of Systems and Software, 34(2):127– 132, August 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ozkul:1997:EAL

[Ozk97] Tarik Ozkul. Estimating acquisition loop time for time-critical real-time data acquisition applications for the purpose of selecting appropriate hardware and software. The Journal of Systems and Software, 37(3):227– 233, June 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Polat:1999:MAT

[PA99] Faruk Polat and Reda Alhajj.

A multi-agent tuple-space based problem solving framework. The Journal of Systems and Software, 47(1):11–17, May 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/47/1/6134.pdf.

Panzl:1981:MES

[Pan81] David J. Panzl. Method for evaluating software development techniques. The Journal of Systems and Software, 2(2):133–137, June 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Parkinson:1986:PAP

[Par86] D. Parkinson. Performance analysis in a 4096 processor environment. The Journal of Systems and Software, 6(1–2):11–15, May 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Parnas:1998:FMT

[Par98]  David Lorge Parnas.  “for-

mal methods” technology transfer will fail. The Journal of Systems and Software, 40(3):195– 198, March 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pasquini:1996:EVD

[Pas96] A. Pasquini. Empirical validation of a domain test selection criterion. The Journal of Systems and Software, 35(2): 107–117, November 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Paul:1992:RC

[Pau92] Raymond A. Paul. Readers’ corner. The Journal of Systems and Software, 17(1):1–3, January 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Parrish:1993:AFG

[PBC93] Allen S. Parrish, Richard B. Borie, and David W. Cordes. Automated flow graph-based testing of object-oriented software modules. The Journal of Systems and Software, 23(2):95–109, November 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Poo:1998:CSM

[PC98a] Danny C. C. Poo and Mui Ken Chung. CASE and software maintenance practices in Singapore. The Journal of Systems and Software, 44(2):97– 105, December 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/2/6075.pdf.

Poo:1998:SEP

[PC98b] Danny C. C. Poo and Mui Ken Chung. Software engineering practices in Singapore. The Journal of Systems and Software, 41(1):3–15, April 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/1/6002.pdf.

Poggi:1998:UPD

[PD98] Agostino Poggi and Giulio Destri. Using PVM to develop a distributed object-oriented language for heterogeneous processing. The Journal of Systems and Software, 40(2):139–150, February 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Perkusich:1994:EFT

[PdC94] Angelo Perkusich, J. C. A. de Figueiredo, and S. K. Chang.

Embedding fault-tolerant properties in the design of complex software systems. The Journal of Systems and Software, 25(1): 23–37, April 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Perkusich:1997:GNP

[PdF97] Angelo Perkusich and Jorge C. A. de Figueiredo. G-Nets: a Petri net based approach for logical and timing analysis of complex software systems. The Journal of Systems and Software, 39 (1):39–59, October 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Prieto-Diaz:1986:MIL

[PDN86] Rubén Prieto-Díaz and James M. Neighbors. Module interconnection languages. The Journal of Systems and Software, 6(4): 307–334, November 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pfleeger:1995:MMG

[Pfl95] Shari Lawrence Pfleeger. Maturity, models and goals: How to build a metrics plan. The Journal of Systems and Software, 31(2):143–155, November 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pfleeger:1999:UIT

[Pfl99] S. L. Pfleeger. Understanding and improving technology transfer in software engineering. The Journal of Systems and Software, 47(2–3):111–124, July 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/47/2-3/
6146.pdf.

Pombortsis:1986:AMM

[PH86] A. Pombortsis and C. Halatsis.

Analysis of a modified model for synchronous multiprocessor systems. The Journal of Systems and Software, 6(4):361–365, November 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pombortsis:1993:PCS

[PH93] A. Pombortsis and C. Halatsis.

Performance of circuit-switched interconnection networks under nonuniform traffic patterns. The Journal of Systems and Software, 20(2):189–201, February 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pham:1994:ODV

[Pha94] Hoang Pham. On the optimal design of N-version software system subject to constraints. The Journal of Systems and Software, 27(1):55–61, October 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Phister:1981:MSD

[Phi81] Montgomery Phister, Jr. Model of the software development process. The Journal of Systems and Software, 2(3):237–255, September 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Philip:1998:SDG

[Phi98] George C. Philip. Software design guidelines for event-driven programming. The Journal of Systems and Software, 41(2): 79–91, May 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/2/6009.pdf.

Pintelas:1989:OSS

[PK89] Panayotis E. Pintelas and Vasilios Kallistros. An overview of some software design languages. The Journal of Systems and Software, 10(2):125–138, September 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Post:1998:CEC

[PKK98] Gerald Post, Albert Kagan, and Robert T. Keim. A comparative evaluation of CASE tools. The Journal of Systems and Software, 44(2):87– 96, December 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/2/6074.pdf.

Powell:1983:DMD

[PL83] Michael L. Powell and Mark A. Linton. Database model of debugging. The Journal of Systems and Software, 3(4):295–300, December 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Poo:1992:ESM

[PL92] Danny C. C. Poo and Paul J.

Layzell. An evolutionary structural model for software maintenance. The Journal of Systems and Software, 18(2):113– 123, May 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Padgham:1994:UIM

[PL94] Lin Padgham and Jonas Löwgren.

A user interface management approach for object-oriented data base applications. The Journal of Systems and Software, 27(3): 183–193, December 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Poo:1996:TIO

[PL96] Danny C. C. Poo and Shwu-Yi Lee. TarTAn: Interweaving objects with rules in information systems development. The Journal of Systems and Software, 33(1):3–14, April 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pfahl:1999:ISD

[PL99] Dietmar Pfahl and Karl Lebsanft. Integration of system dynamics modelling with descriptive process modelling and goal-oriented measurement. The Journal of Systems and Software, 46(2–3):135–150, April 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/46/2-3/
6122.pdf.

Plant:1992:ESD

[Pla92]  R. T. Plant.  Expert sys-

tem development and testing: a knowledge engineer’s perspective. The Journal of Systems and Software, 19(2):141–146, October 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Plant:1995:GEC

[Pla95] Robert Plant. Guest Editor’s corner: special issue on software quality in knowledge-based systems. The Journal of Systems and Software, 29(3):197– 198, June 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Papazoglou:1990:OOA

[PM90a] M. P. Papazoglou and L. Marinos. An object-oriented approach to distributed data management. The Journal of Systems and Software, 11(2):95–109, February 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pfleeger:1990:SMP

[PM90b] Shari Lawrence Pfleeger and Clement McGowan. Software metrics in the process maturity framework. The Journal of Systems and Software, 12(3):255– 261, July 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pissinou:1994:CAT

[PM94] Niki Pissinou and Kia Makki.

A coherent architecture for a temporal object data base management system. The Journal of Systems and Software, 27(3): 195–205, December 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

P:1999:OTA

[PM99] Ajith Tom P. and C. Siva Ram Murthy. Optimal task allocation in distributed systems by graph matching and state space search. The Journal of Systems and Software, 46(1): 59–75, April 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/1/6115.pdf.

Powell:1999:SLC

[PMB99] Antony Powell, Keith Mander, and Duncan Brown. Strategies for lifecycle concurrency and iteration — a system dynamics approach. The Journal of Systems and Software, 46(2–3): 151–161, April 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/2-3/6123.pdf.

Park:1996:FPS

[PNK96] Dong-Won Park, Swaminathan Natarajan, and Arkady Kanevsky. Fixed-priority scheduling of realtime systems using utilization bounds. The Journal of Systems and Software, 33(1):57–63, April 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Poo:1993:IES

[Poo93] Danny C. C. Poo. Implementing an evolutionary structural software model. The Journal of Systems and Software, 22(2):81– 90, August 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Popescu:1992:RMK

[Pop92] Ilié Popescu. A relational model for knowledge representation in expert systems. The Journal of Systems and Software, 18(2): 147–155, May 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Porter:1993:UMD

[Por93]  Adam A. Porter.  Using

measurement-driven modeling to provide empirical feedback to software developers. The Journal of Systems and Software, 20 (3):237–243, March 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Poulin:1995:PSR

[Pou95] Jeffrey S. Poulin. Populating software repositories: Incentives and domain-specific software. The Journal of Systems and Software, 30(3):187– 199, September 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Powell:1986:HAD

[Pow86] D. R. Powell. A hierarchical approach to distributed computing system dependability evaluation. The Journal of Systems and Software, 6(1–2):183– 198, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Prasad:1994:RSA

[PP94] Aarthi Prasad and E. K. Park.

Reuse system: an artificial intelligence—based approach. The Journal of Systems and Software, 27(3):207–221, December 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pressman:1990:CCD

[Pre90] Israel Pressman. Controversy corner: The decision makers. The Journal of Systems and Software, 11(1):71–74, January 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Preece:1995:TQA

[Pre95] Alun D. Preece. Toward a quality assessment framework for knowledge-based systems. The Journal of Systems and Software, 29(3):219–234, June 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Porter:1990:ETG

[PS90] Adam A. Porter and Richard W. Selby. Evaluating techniques for generating metric-based classification trees. The Journal of Systems and Software, 12(3):209– 218, July 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pintelas:1991:CSF

[PT91] P. Pintelas and S. Tragoudas. A comparative study of five language independent programming environments. The Journal of Systems and Software, 14(1):3– 15, January 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Probert:1984:HTE

[PU84a] Robert L. Probert and Hasan Ural. High-level testing and example-directed development of software specifications. The Journal of Systems and Software, 4(4):2–3, November 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Probert:1984:HLT

[PU84b] Robert L. Probert and Hasan Ural. High-level testing and example-directed development of software specifications. The Journal of Systems and Software, 4(4):317–325, November 1984. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pulk:1990:CCI

[Pul90] Brian E. Pulk. Controversy corner: Improving software project management. The Journal of Systems and Software, 13(3): 231–235, November 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pombortsis:1994:CPA

[PV94] A. Pombortsis and I. Vlahavas.

A contribution to the problem of avoiding congestion in multistage networks in the presence of unbalanced traffic. The Journal of Systems and Software, 26(3): 273–284, September 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Parnas:1987:ADR

[PW87] D. L. Parnas and D. M. Weiss.

Active design reviews: Principles and practices. The Journal of Systems and Software, 7(4): 259–265, December 1987. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Purtilo:1992:FPA

[PW92] James M. Purtilo and Elizabeth L. White. A flexible program adaptation system: Case studies in Ada. The Journal of Systems and Software, 17(2): 129–143, February 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Portman:1994:DIR

[PZ94] M. Portman and H. S. M. Zedan.

The development of imprecise real-time systems. The Journal of Systems and Software, 27 (2):73–83, November 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Quintas:1994:CCS

[Qui94] Paul Quintas. Controversy Corner: Software engineering policy and practice: Lessons from the Alvey Program. The Journal of Systems and Software, 24(1):67– 88, January 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Reid:1991:CCC

[RA91] Richard A. Reid and Peter L. Asprey. Controversy corner: Computer performance management: Conceptual considerations from an operations research perspective. The Journal of Systems and Software, 14(3):215– 219, March 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rumerstorfer:1996:BFS

[RA96] Hubert Rumerstorfer and Josef Altmann. BOOM — a first step to an object-oriented fourth generation system. The Journal of Systems and Software, 33(2): 171–177, May 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Redmond:1990:SMU

[RAC90] James A. Redmond and Reynold Ah-Chuen. Software metrics — a user’s perspective. The Journal of Systems and Software, 13 (2):97–110, October 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rader:1984:VSE

[Rad84] Jock A. Rader. VLSI and software engineering: Guest Editor’s introduction. The Journal of Systems and Software, 4(1): 3–6, April 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rahm:1992:FWA

[Rah92] Erhard Rahm. A framework for workload allocation in distributed transaction processing systems. The Journal of Systems and Software, 18(2):171– 190, May 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rajlich:1985:SRR

[Raj85] Vaclav Rajlich. Stepwise refinement revisited. The Journal of Systems and Software, 5(1):81– 88, February 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rajlich:1994:DGM

[Raj94] Vaclav Rajlich. Decomposition/ generalization methodology for object-oriented programming. The Journal of Systems and Software, 24(2):181–186, February 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ramarao:1990:EFT

[Ram90] K. V. S. Ramarao. Efficient fault-tolerant broadcasts. The Journal of Systems and Software, 11(2):131–141, February 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Raveling:1981:SOD

[Rav81] Jerry Raveling. Status and outlook for DoD configuration management requirements. The Journal of Systems and Software, 2(4):363–370, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Robillard:1989:IMN

[RB89] Pierre N. Robillard and Germinal Boloix. The interconnectivity metrics: a new metric showing how a program is organized. The Journal of Systems and Software, 10(1):29–39, July 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rijsenbrij:1993:PDP

[RB93a] D. B. B. Rijsenbrij and A. H. Bauer. Project diagnosis: a proper start is half the battle. The Journal of Systems and Software, 21(1):17–25, April 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rijsenbrij:1993:QSS

[RB93b] D. B. B. Rijsenbrij and A. H. Bauer. A quality system for a software house. The Journal of Systems and Software, 23(3): 211–224, December 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ramesh:1999:ECR

[RB99] V. Ramesh and Glenn J.

Browne. Expressing casual relationships in conceptual database schemas. The Journal of Systems and Software, 45(3):225– 232, March 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/3/6108.pdf.

Robson:1991:APC

[RBCM91] D. J. Robson, K. H. Bennett, B. J. Cornelius, and M. Munro. Approaches to program comprehension. The Journal of Systems and Software, 14(2):79–84, February 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rader:1995:OUC

[RBM95] Jock A. Rader, Alan W. Brown, and Ed J. Morris. Operational use of CASE integration: An investigation of the state of the practice. The Journal of Systems and Software, 28(1):59–68, January 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ram:1989:ADD

[RC89] Sudha Ram and Clark L. Chastain. Architecture of distributed data base systems. The Journal of Systems and Software, 10 (2):77–95, September 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rising:1994:IHM

[RC94] Linda S. Rising and Frank W. Calliss. An information-hiding metric. The Journal of Systems and Software, 26(3):211– 220, September 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rus:1999:SPS

[RCL99] Ioana Rus, James Collofello, and Peter Lakey. Software process simulation for reliability management. The Journal of Systems and Software, 46(2–3): 173–182, April 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1999/46/2-3/6125.pdf.

Rong-Chau:1993:PMA

[RCSD93] Liu Rong-Chau and Wang Sheng-De. Performance modeling and analysis of load balancing policies with priority queueing. The Journal of Systems and Software, 20(2):169–187, February 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Reck:1993:FSA

[Rec93] Martin Reck. Formally specifying an automated trade execution system. The Journal of Systems and Software, 21(3):245– 252, June 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Reed:1985:CST

[Ree85] Alan Reed. Correct: a simulated text editor. The Journal of Systems and Software, 5(3):179– 184, August 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Reifer:1987:SRU

[Rei87] Donald J. Reifer. Softcost-R: User experiences and lessons learned at the age of one. The Journal of Systems and Software, 7(4):279–286, December 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Reifer:1990:ARF

[Rei90a] Donald J. Reifer. Asset-R:

a function point sizing tool for scientific and real-time systems. The Journal of Systems and Software, 11(3):159– 171, March 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Reifer:1990:CCD

[Rei90b]  Donald J. Reifer.  Contro-

versy corner: a debate, Zucconi vs. Reifer: Is Asia really a threat? The Journal of Systems and Software, 12(1):74– 77, April 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See also [Zuc90b, Zuc90a].

Reynolds:1980:ECS

[Rey80] John H. Reynolds. Evaluation of contemporary software engineering techniques for a large Fortran simulation. The Journal of Systems and Software, 1(2):131– 140, February 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Reynolds:1984:MMC

[Rey84] Robert G. Reynolds. Metrics to measure the complexity of partial programs. The Journal of Systems and Software, 4(1):75– 91, April 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Reynolds:1989:PMS

[Rey89] Robert G. Reynolds. The partial metrics system: a tool to support the metrics-driven design of pseudocode programs. The Journal of Systems and Software, 9 (4):287–295, May 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ruspini:1984:III

[RF84] Enrique H. Ruspini and Robert Fraley. Id: an intelligent information dictionary system. The Journal of Systems and Software, 4(2–3):197–205, July 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rodriguez:1979:DFB

[RG79] Jorge E. Rodriguez and Sol J. Greenspan. Directed flowgraphs: The basis of a specification and construction methodology for real-time systems. The Journal of Systems and Software, 1 (1):19–27, ???? 1979. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Riddle:1981:GEI

[Rid81] William E. Riddle. Guest Editor’s introduction: Tool and methodology evaluation. The Journal of Systems and Software, 2(2):85–88, June 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ryu:1993:CIA

[RMC93] K. Ryu, S. R. Maeng, and J. W. Cho. Concurrency and inheritance in actor-based objectoriented languages. The Journal of Systems and Software, 20 (1):53–67, January 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Robertson:1998:ESL

[Rob98] David Robertson. An empirical study of the LSS specification toolkit in use. The Journal of Systems and Software, 42(2): 115–123, August 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/2/6031.pdf.

Rodrigue:1986:SPD

[Rod86] Garry Rodrigue. Solving partial differential equations on multiprocessors. The Journal of Systems and Software, 6(1–2):147– 154, May 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rogers:1989:CAM

[Rog89] McCagie B. Rogers. Considerations for applying multiple-level protocols to satellite communication networks. The Journal of Systems and Software, 10(3): 201–212, October 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rogers:1994:MPC

[Rog94] McCagie Brooks Rogers. The MYTHSEEKER project: Concepts and realization. The Journal of Systems and Software, 25 (3):271–298, June 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Romanovsky:1998:SAA

[Rom98] A. Romanovsky. A study of atomic action schemes intended for standard Ada. The Journal of Systems and Software, 43(1): 29–44, October 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/43/1/6048.pdf.

Romanovsky:1999:CDS

[Rom99] Alexander Romanovsky. Class diversity support in objectoriented languages. The Journal of Systems and Software, 48 (1):43–57, August 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/29/11/37/17/
21/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/
29/11/37/17/21/article.pdf.

Rosselet:1987:LDB

[Ros87] A. Rosselet. Language definitionbased compiler development. The Journal of Systems and Software, 7(2):145–161, June 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rotenstreich:1989:RDP

[Rot89] Shmuel Rotenstreich. Reliability of distributed programs under overloads. The Journal of Systems and Software, 9(1):29– 40, January 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Roweth:1986:DPA

[Row86] D. Roweth. Design and performance analysis of Transputer arrays. The Journal of Systems and Software, 6(1–2):21–22, May 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rehder:1997:SCS

[RPL97] Bob Rehder, Nancy Pennington, and Adrienne Y. Lee. Scoring the completeness of software designs. The Journal of Systems and Software, 36(1):33–68, January 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rajlich:1998:CSE

[RR98] Vaclav Rajlich and Shivkumar Ragunathan. A case study of evolution in object oriented and heterogeneous architectures. The Journal of Systems and Software, 43(2):85– 91, November 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/2/6052.pdf.

Rine:1998:IRS

[RS98] David C. Rine and Robert M. Sonnemann. Investments in reusable software. A study of software reuse investment success factors. The Journal of Systems and Software, 41(1): 17–32, April 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/1/6003.pdf.

Rai:1998:SQA

[RST98] Arun Rai, Haidong Song, and Marvin Troutt. Software quality assurance: an analytical survey and research prioritization. The Journal of Systems and Software, 40(1):67–83, January 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Robert:1986:PSB

[RT86] Y. Robert and M. Tchuente. Parallel solution of band triangular linear systems on VLSI arrays with limited fan-out. The Journal of Systems and Software, 6 (1–2):155–163, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rubinovitz:1993:DIQ

[RT93] Harvey Rubinovitz and Bhavani Thuraisingham. Design and implementation of a query processor for a trusted distributed data base management system. The Journal of Systems and Software, 21(1):49–69, April 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Russell:1990:ISS

[Rus90] Meg Russell. International survey of software measurement education and training. The Journal of Systems and Software, 12 (3):233–241, July 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rombach:1992:TFL

[RUV92] H. Dieter Rombach, Bradford T.

Ulery, and Jon D. Valett. Toward full life cycle control: Adding maintenance measurement to the SEL. The Journal of Systems and Software, 18(2):125– 138, May 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Roeseler:1991:MQC

[Rv91] Armin Roeseler and Anneliese von Mayrhauser. Managing the quality of computing services: a user-oriented approach using utility theory and service-level indices. The Journal of Systems and Software, 16(3):185– 196, November 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Roeseler:1992:USL

[Rv92] Armin Roeseler and Anneliese von Mayrhauser. Using servicelevel indices to manage the quality of computing services: a case study. The Journal of Systems and Software, 18(2):139– 145, May 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Roeseler:1993:PBA

[Rv93] Armin Roeseler and Anneliese von Mayrhauser. A productionbased approach to performance evaluation of computing technology. The Journal of Systems and Software, 20(2):153–162, February 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Raffo:1999:SPS

[RVM99] David M. Raffo, Joseph V. Vandeville, and Robert H. Martin. Software process simulation to achieve higher CMM levels. The Journal of Systems and Software, 46(2–3):163–172, April 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/46/2-3/
6124.pdf.

Ravoor:1997:MTP

[RW97] Suresh B. Ravoor and Johnny S. K. Wong. Multithreaded transaction processing in distributed systems. The Journal of Systems and Software, 38 (2):107–117, August 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Raz:1993:PCA

[RY93] Tzvi Raz and Alan Yaung. Process clustering with an algorithm based on a coupling metric. The Journal of Systems and Software, 22(3):217–223, September 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Raffa:1994:RPS

[RZ94] M. Raffa and G. Zollo. The role of professionals in small Italian software firms. The Journal of Systems and Software, 26(1): 19–30, July 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sarwar:1993:FLP

[SAA93] S. Mansoor Sarwar and Marwan H. Abu-Amara. Functional languages: a performance study. The Journal of Systems and Software, 20(2):163– 167, February 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saleh:1994:ERP

[SAASA94] Kassem Saleh, Imtiaz Ahmad, Khaled Al-Saqabi, and Anjali Agarwal. An efficient recovery procedure for fault tolerance in distributed systems. The Journal of Systems and Software, 25(1): 39–50, April 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sage:1995:SES

[Sag95] Andrew P. Sage. Systems engineering and systems management for reengineering. The Journal of Systems and Software, 30(1–2):3–25, July/August 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sahraoui:1994:STA

[Sah94] A. E. K. Sahraoui. Some timing aspects of software development for reactive systems: The two-axis paradigm. The Journal of Systems and Software, 25(1): 51–57, April 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saiedian:1998:GEC

[Sai98] Hossein Saiedian. Guest Editor’s corner: research directions in formal methods technology transfer. The Journal of Systems and Software, 40(3):187– 189, March 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saiedian:1999:SEE

[Sai99] Hossein Saiedian. Software engineering education and training for the next millennium. The Journal of Systems and Software, 49(2–3):113–115, December 30, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

elsevier.nl/gej-ng/10/29/
11/46/26/25/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

Sakai:1984:ERB

[Sak84] Hirotaka Sakai. Entity-relationship behavior modeling in conceptual schema design. The Journal of Systems and Software, 4(2–3): 135–145, July 1984. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Salisbury:1980:TCH

[Sal80] Alan B. Salisbury. TACFIRE:

a case history of a weapon system software development. The Journal of Systems and Software, 1(2):155–175, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Samson:1993:KBT

[Sam93] Dolly Samson. Knowledge-based test planning: Framework for a knowledge-based system to prepare a system test plan from system requirements. The Journal of Systems and Software, 20(2): 115–124, February 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sanden:1995:DCS

[San95] Bo I. Sanden. Designing control systems with entity-life modeling. The Journal of Systems and Software, 28(3):225– 237, March 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Swanson:1988:UCS

[SB88] E. Burton Swanson and Cynthia M. Beath. Use of case study data in software management research. The Journal of Systems and Software, 8(1):63– 71, January 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Subramanian:1993:DRS

[SB93] Girish H. Subramanian and Steven Breslawski. Dimensionality reduction in software development effort estimation. The Journal of Systems and Software, 21(2):187–196, May 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Subramanian:1995:EAS

[SB95] Girish H. Subramanian and Steven Breslawski. An empirical analysis of software effort estimate alterations. The Journal of Systems and Software, 31(2): 135–141, November 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Staalhane:1997:SCQ

[SBA97] T. Stålhane, P. C. Borgersen, and K. Arnesen. In search of the customer’s quality view. The Journal of Systems and Software, 38(1):85–93, July 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Son:1998:DTD

[SBB98] Sang H. Son, Robert C.

Beckinger, and David A. Baker. DRDB: Towards distributed real-time database services for time-critical active applications. The Journal of Systems and Software, 42(3):193– 204, September 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/42/3/6038.pdf.

Stansifer:1994:MCP

[SBM94] Ryan Stansifer, Mike Beaven, and Dan C. Marinescu. Modeling concurrent programs with colored Petri nets. The Journal of Systems and Software, 26 (2):129–148, August 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shatz:1988:PNF

[SC88] S. M. Shatz and W. K. Cheng.

A Petri net framework for automated static analysis of Ada tasking behavior. The Journal of Systems and Software, 8(5): 343–359, December 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saiedian:1999:TEF

[SC99] Hossein Saiedian and Kalyani Chennupati. Towards an evaluative framework for software process improvement models. The Journal of Systems and Software, 47(2–3):139–148, July 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/47/2-3/
6149.pdf.

Scanlan:1988:LPU

[Sca88] David A. Scanlan. Learner preference for using structured flowcharts vs. pseudocode when comprehending short, relatively complex algorithms: a summary analysis. The Journal of Systems and Software, 8(2):145–155, March 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Scacchi:1999:ESP

[Sca99] Walt Scacchi. Experience with software process simulation and modeling. The Journal of Systems and Software, 46(2–3):183– 192, April 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/2-3/6126.pdf.

Scholtz:1993:OOP

[SCG^{+}93] Jean Scholtz, Shyam Chidamber, Robert Glass, Al Goerner, Mary Beth Rosson, Mike Stark, and Iris Vessey. Object-oriented programming: The promise and the reality. The Journal of Systems and Software, 23(2): 199–204, November 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schuman:1981:NSD

[Sch81] Antonia D. Schuman. New software documentation standard: Report on the Joint Logistic Commanders Joint Policy Coordinating Group on Computer Resource Management, Computer Software Management Subgroup, Software Workshop. The Journal of Systems and Software, 2(4):329–336, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schmidt:1991:PAN

[Sch91] Heinz W. Schmidt. Prototyping and analysis of non-sequential systems using predicate-event nets. The Journal of Systems and Software, 15(1):43–62, April 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schneberger:1997:DCE

[Sch97]  Scott L. Schneberger.  Dis-

tributed computing environments: Effects on software maintenance difficulty. The Journal of Systems and Software, 37(2): 101–116, May 1997. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sum:1986:AOS

[SCK86] R. N. Sum, Jr., R. H. Campbell, and W. J. Kubitz. An approach to operating system testing. The Journal of Systems and Software, 6(3):273–284, August 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sindre:1995:RAS

[SCK95] Guttorm Sindre, Reidar Conradi, and Even-Andre Karlsson. The REBOOT approach to software reuse. The Journal of Systems and Software, 30(3):201– 212, September 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sarwar:1994:NFC

[SD94] S. Mansoor Sarwar and James A. Davis. New families of combinators for efficient list manipulation. The Journal of Systems and Software, 27(2):137–146, November 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Seddio:1993:ITM

[Sed93] Carl Seddio. Integrating test metrics within a software engineering measurement program at Eastman Kodak Company: a follow-up case study. The Journal of Systems and Software, 20 (3):227–235, March 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Seidewitz:1989:GOO

[Sei89] E. Seidewitz. General objectoriented software development: Background and experience. The Journal of Systems and Software, 9(2):95–108, February 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Selby:1993:IAT

[Sel93] Richard W. Selby. Interconnectivity analysis techniques for error localization in large systems. The Journal of Systems and Software, 20(3):267– 272, March 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stytz:1992:DAS

[SF92] M. R. Stytz and O. Frieder. Dynamic adaptive surface rendering within a distributed memory multiprocessor. The Journal of Systems and Software, 19(3): 247–259, November 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Storey:1999:CDE

[SFM99] M.-A. D. Storey, F. D. Fracchia, and H. A. Müller. Cognitive design elements to support the construction of a mental model during software exploration. The Journal of Systems and Software, 44(3):171– 185, January 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/44/3/6083.pdf.

Schwan:1989:ARA

[SG89] K. Schwan and C. Gaimon. Automating resource allocation for multiprocessors. The Journal of Systems and Software, 9(2):51– 66, February 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Scheff:1991:UDB

[SG91] Benson H. Scheff and Thomas Georgon. Using documentation blueprints to produce mandated DoD data items. The Journal of Systems and Software, 14(2): 103–110, February 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saacks-Giguette:1993:FBD

[SGL93] Marguerite Saacks-Giguette and Antonio M. Lopez, Jr. A framebased design for the TIMS and CAMS metadata for a stennis information management system. The Journal of Systems and Software, 20(1):87–92, January 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shock:1998:CSS

[SH98] Robert C. Shock and Thomas C. Hartrum. A classification scheme for software modules. The Journal of Systems and Software, 42 (1):29–44, July 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/42/1/6025.pdf.

Sheu:1989:DSD

[She89] Phillip C. Sheu. Describing semantic data bases with logic. The Journal of Systems and Software, 9(1):19–27, January 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sheu:1990:KBA

[She90] Phillip C.-Y. Sheu. A knowledgebased approach for high-level programming of concurrent systems. The Journal of Systems and Software, 11(1):21–29, January 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sherer:1994:MSF

[She94] Susan A. Sherer. Measuring software failure risk: Methodology and an example. The Journal of Systems and Software, 25(3): 257–269, June 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sherer:1995:SFP

[She95] Susan A. Sherer. Software fault prediction. The Journal of Systems and Software, 29(2):97–105, May 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shoja:1991:DFL

[Sho91] G. C. Shoja. A distributed facility for load sharing and parallel processing among workstations. The Journal of Systems and Software, 14(3):163– 172, March 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shum:1999:EFT

[Shu99] Kam Hong Shum. Effective fault tolerance for agent-based cluster computing. The Journal of Systems and Software, 48(3):189– 196, November 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/37/24/20/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/24/abstract.
html.

Shepperd:1994:CTM

[SI94] M. Shepperd and D. C. Ince. A critique of three metrics. The Journal of Systems and Software, 26(3):197–210, September 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shahmehri:1995:UCA

[SKF95] Nahid Shahmehri, Mariam Kamkar, and Peter Fritzson. Usability criteria for automated debugging systems. The Journal of Systems and Software, 31(1):55– 70, October 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sobol:1996:PCR

[SKS96] Marion G. Sobol, Albert Kagan, and Hirohisa Shimura. Performance criteria for relational databases in different normal forms. The Journal of Systems and Software, 34(1):31–42, July 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Skuce:1991:LSM

[Sku91] Douglas Skuce. A language and system for making definitions of technical concepts. The Journal of Systems and Software, 14 (1):39–50, January 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stark:1994:SMS

[SKV94] George E. Stark, Louise C. Kern, and C. W. Vowell. A software metric set for program maintenance management. The Journal of Systems and Software, 24 (3):239–249, March 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schick:1980:USP

[SL80] G. J. Schick and Chi Yuan Lin.

Use of a subjective prior distribution for the reliability of computer software. The Journal of Systems and Software, 1(3):259– 266, ???? 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shin:1996:PMA

[SL96] Hyunsik Shin and Jinjoo Lee. A process model of application software package acquisition and implementation. The Journal of Systems and Software, 32(1):57– 64, January 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Salisbury:1980:EI

[SM80] Alan B. Salisbury and John H.

Manley. Editors’ introduction. The Journal of Systems and Software, 1(4):271, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Salisbury:1981:EIa

[SM81a] Alan B. Salisbury and John H. Manley. Editors’ introduction. The Journal of Systems and Software, 2(1):1, February 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Salisbury:1981:EIb

[SM81b] Alan B. Salisbury and John H. Manley. Editors’ introduction. The Journal of Systems and Software, 2(2):83, June 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Salisbury:1981:EIc

[SM81c] Alan B. Salisbury and John H. Manley. Editors’ introduction. The Journal of Systems and Software, 2(3):191, September 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Salisbury:1981:EId

[SM81d] Alan B. Salisbury and John H. Manley. Editors’ introduction. The Journal of Systems and Software, 2(4):279, December 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Salisbury:1983:EI

[SM83] Alan B. Salisbury and John H. Manley. Editors’ introduction. The Journal of Systems and Software, 3(1):1, March 1983. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Spangler:1992:SFC

[SM92a] William E. Spangler and Jerrold H. May. Success and failure in cooperative expert systems development: a tale of two projects. The Journal of Systems and Software, 19(2):131– 140, October 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Summers:1992:CCC

[SM92b] Craig Summers and Eric Markusen.

Controversy corner: Computers, ethics, and collective violence. The Journal of Systems and Software, 17(1):91–103, January 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shieh:1996:OOA

[SMCL96] Ce-Kuen Shieh, Su-Cheong Mac, Tzu-Chiang Chang, and Chung-Ming Lai. An object-oriented approach to develop software faulttolerant mechanisms for parallel programming systems. The Journal of Systems and Software, 32(3):215–225, March 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Siegel:1994:CIC

[SMS94] Michael Siegel, Stuart Madnick, and Edward Sciore. Context interchange in a client-server architecture. The Journal of Systems and Software, 27(3):223–232, December 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shieh:1998:IPD

[SMU98] Ce-Kuen Shieh, Su-Cheong Mac, and Jyh-Chang Ueng. Improving the performance of distributed shared memory systems via parallel file input/output. The Journal of Systems and Software, 44 (1):3–15, December 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/44/1/6067.pdf.

Samadzadeh:1991:SSM

[SN91] M. H. Samadzadeh and K. Nandakumar. A study of software metrics. The Journal of Systems and Software, 16(3): 229–234, November 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sneed:1983:SSE

[Sne83] Harry M. Sneed. Softening software engineering system. The Journal of Systems and Software, 3(1):63–76, March 1983. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Snyder:1979:IUC

[Sny79] Samuel S. Snyder. Influence of U.S. Cryptologic Organizations on the digital computer industry. The Journal of Systems and Software, 1(1):87–102, ???? 1979. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Snyder:1991:STG

[Sny91] Robin M. Snyder. Specifying textual to graphical conversion. The Journal of Systems and Software, 16(1):17–28, September 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Soloway:1987:SSE

[Sol87] Elliot Soloway. Special section on empirical studies of programming. The Journal of Systems and Software, 7(4):323, December 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Song:1993:LTG

[Son93] J. Song. A layout tool for Glotos.

The Journal of Systems and Software, 21(2):151–162, May 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stotts:1994:PFA

[SP94] P. David Stotts and William Pugh. Parallel finite automata for modeling concurrent software systems. The Journal of Systems and Software, 27(1):27–43, October 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Spafford:1992:CHB

[Spa92] Eugene H. Spafford. Are computer hacker break-ins ethical? The Journal of Systems and Software, 17(1):41–47, January 1992.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saleh:1999:DOC

[SPK99] Kassem Saleh, Robert Probert, and Hassib Khanafer. The distributed object computing paradigm: concepts and applications. The Journal of Systems and Software, 47(2–3):125–131, July 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/47/2-3/
6147.pdf.

Sherif:1998:MOO

[SS98] Joseph S. Sherif and Peter Sanderson. Metrics for objectoriented software projects. The Journal of Systems and Software, 44(2):147–154, December 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/2/6079.pdf.

Schuba:1998:PEC

[SSK98] Christoph L. Schuba, Eugene H. Spafford, and Berry Kercheval. Prototyping experiences with classical IP and ARP over signaled ATM connections. The Journal of Systems and Software, 44(1):31– 43, December 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.

com/cas/tree/store/jss/sub/
1998/44/1/6069.pdf.

Sikkel:1999:RWC

[SSvdW99] Klaas Sikkel, Ton A. M. Spil, and Rob L. W. van de Weg. A real-world case study in information technology for undergraduate students. The Journal of Systems and Software, 49(2– 3):117–123, December 30, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/26/26/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

Shieh:1989:ESA

[ST89] Ce Kuen Shieh and Li Ming Tseng. Extending a stand-alone personal computer to integrate multiple operating systems concurrently. The Journal of Systems and Software, 9(1):41–49, January 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stavely:1983:MPS

[Sta83] Allan M. Stavely. Modeling and projection in software development. The Journal of Systems and Software, 3(2):137–146, June 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stavely:1985:IMS

[Sta85] Allan M. Stavely. Inference from models of software systems. The Journal of Systems and Software, 5(3):185–191, August 1985. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stavely:1990:AAC

[Sta90] Allan M. Stavely. Algorithms for analyzing concurrent software systems using derivatives. The Journal of Systems and Software, 11(1):3–20, January 1990. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stark:1993:IOO

[Sta93a] Mike Stark. Impacts of objectoriented technologies: Seven years of software engineering. The Journal of Systems and Software, 23(2):163–169, November 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stavely:1993:ESI

[Sta93b] Allan M. Stavely. An empirical study of iteration in applications software. The Journal of Systems and Software, 22(3): 167–177, September 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stavridou:1999:ISI

[Sta99] Victoria Stavridou. Integration in software intensive systems. The Journal of Systems and Software, 48(2):91–104, October 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/29/
                                        11/37/23/19/article.pdf;

http://www.elsevier.nl/gej-
ng/10/29/11/37/23/abstract.
html.

Sedlmeyer:1983:KBF

[STJ83] Robert L. Sedlmeyer, William B.

Thompson, and Paul E. Johnson. Knowledge-based fault localization in debugging. The Journal of Systems and Software, 3(4): 301–307, December 1983. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stoyenko:1992:ESA

[Sto92] Alexander D. Stoyenko. The evolution and state-of-the-art of real-time languages. The Journal of Systems and Software, 18(1): 61–83, April 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stuebing:1983:IWS

[Stu83] H. G. Stuebing. Industrialization of weapon system software. The Journal of Systems and Software, 3(2):163–171, June 1983. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Subramanian:1993:EES

[Sub93] Girish H. Subramanian. An empirical examination of software development modes. The Journal of Systems and Software, 23 (1):3–7, October 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Swigger:1988:DPP

[SW88] Kathleen M. Swigger and Layne F. Wallace. A discussion of past programming errors and their effect on learning assembly language. The Journal of Systems and Software, 8(5):395–399, December 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saiedian:1993:COO

[SW93] Hossein Saiedian and Stan Wileman. A concurrent objectoriented framework for simulation of network protocols. The Journal of Systems and Software, 23(2):139–150, November 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shah:1994:TMO

[SW94a] Pinaki Shah and Johnny Wong. Transaction management in an object-oriented database system. The Journal of Systems and Software, 24(2):115–124, February 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Staalhane:1994:QRC

[SW94b] Tor Stålhane and Kari J. Wedde. The quest for reliability: a case study. The Journal of Systems and Software, 26(1):69–76, July 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Semmel:1995:GEC

[SW95a] R. D. Semmel and M. Wilson. Guest Editors’ corner: Reengineering complex systems. The Journal of Systems and Software, 30(1–2):1–2, July/August 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Semmel:1995:IRD

[SW95b] R. D. Semmel and R. P. Winkler. Integrating reengineered databases to support data fusion. The Journal of Systems and Software, 30(1–2):127–135, July/August 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shah:1996:CCO

[SW96] Pinaki Shah and Johnny Wong.

Concurrency control in an object-oriented data base system. The Journal of Systems and Software, 35(3):169–183, December 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Smith:1999:PMI

[SW99] C. U. Smith and L. G. Williams.

A performance model interchange format. The Journal of Systems and Software, 49(1):63– 80, December 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/25/31/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/25/abstract.
html.

Subramanian:1997:EEF

[SYB97] Girish H. Subramanian, Gayle J.

Yaverbaum, and Stephen J.

Brandt. An empirical evaluation of factors influencing expert systems effectiveness. The Journal of Systems and Software, 38(3): 255–261, September 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shi:1998:EMC

[SZ98] Q. Shi and N. Zhang. An effective model for composition of secure systems. The Journal of Systems and Software, 43(3): 233–244, November 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/43/3/6064.pdf.

Taivalsaari:1993:NO

[Tai93] Antero Taivalsaari. On the notion of object. The Journal of Systems and Software, 21(1):3– 16, April 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Takahashi:1997:SQC

[Tak97] Ryouei Takahashi. Software quality classification model based on McCabe’s complexity measure. The Journal of Systems and Software, 38(1):61–69, July 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tang:1996:NDO

[Tan96] Shung-Ming Tang. A new data organizing algorithm for parallel searching. The Journal of Systems and Software, 32(2):121– 133, February 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tardy:1992:SSA

[Tar92] Jean E. Tardy. Strategies for software acquisition. The Journal of Systems and Software, 18 (3):281–285, July 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tausworthe:1980:WBS

[Tau80] Robert C. Tausworthe. Work breakdown structure in software project management. The Journal of Systems and Software, 1 (3):181–186, ???? 1980. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tausworthe:1992:CCI

[Tau92] Robert C. Tausworthe. Controversy corner: Information models of software productivity: Limits on productivity growth. The Journal of Systems and Software, 19(2):185–201, October 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Turley:1995:CEN

[TB95] Richard T. Turley and James M.

Bieman. Competencies of exceptional and nonexceptional software engineers. The Journal of Systems and Software, 28(1):19– 38, January 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Terwilliger:1989:EEI

[TC89a] Robert B. Terwilliger and Roy H. Campbell. ENCOMPASS: An environment for the incremental development of software. The Journal of Systems and Software, 10(1):41–53, July 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Terwilliger:1989:PES

[TC89b] Robert B. Terwilliger and Roy H. Campbell. PLEASE: Executable specifications for incremental software development. The Journal of Systems and Software, 10(2):97–112, September 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tang:1993:URH

[TC93] S. M. Tang and L. R. Chow. Using rank hashing function to develop a data allocating scheme for a two-disk system. The Journal of Systems and Software, 22 (2):107–115, August 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Toyn:1998:PLT

[TCMJ98] I. Toyn, D. M. Cattrall, J. A. McDermid, and J. L. Jacob. A practical language and toolkit for high-integrity tools. The Journal of Systems and Software, 41 (3):161–173, June 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/41/3/6017.pdf.

Thibodeau:1980:LCP

[TD80] Robert Thibodeau and E. N.

Dodson. Life cycle phase interrelationships. The Journal of Systems and Software, 1(3):203– 211, ???? 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thomas:1997:AER

[TDB97] William M. Thomas, Alex Delis, and Victor R. Basili. An analysts of errors in a reuse-oriented development environment. The Journal of Systems and Software, 38(3):211–224, September 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thompson:1999:PNG

[TE99] J. Barrie Thompson and Helen M. Edwards. Providing new graduate opportunities: experiences with a UK Master’s level computing conversion course. The Journal of Systems and Software, 49(2–3):135– 143, December 30, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/26/28/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

Turner:1999:CBF

[TFLW99] C. Reid Turner, Alfonso Fuggetta, Luigi Lavazza, and Alexander L. Wolf. A conceptual basis for feature engineering. The Journal of Systems and Software, 49(1): 3–15, December 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic).  URL http://www.
                                        elsevier.nl/gej-ng/10/29/
                                        11/46/25/26/article.pdf;
                                        http://www.elsevier.nl/gej-
                                        ng/10/29/11/46/25/abstract.
                                        html.

Thayer:1980:OSU

[Tha80] Richard H. Thayer. Organizational structures used in software development by the U.S. aerospace industry. The Journal of Systems and Software, 1(4): 283–297, ???? 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thimbleby:1994:CCO

[Thi94] Harold Thimbleby. Controversy corner: an organizational solution to piracy and viruses. The Journal of Systems and Software, 25(2):207–215, May 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tian:1996:IAT

[Tia96] Jeff Tian. An integrated approach to test tracking and analysis. The Journal of Systems and Software, 35(2):127–140, November 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tian:1999:MCI

[Tia99] Jeff Tian. Measurement and continuous improvement of software reliability throughout software life-cycle. The Journal of Systems and Software, 47(2–3): 189–195, July 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/47/2-3/6154.pdf.

Thomsen:1987:TPL

[TK87] Kristine Stougaard Thomsen and Jorgen Lindskov Knudsen. Taxonomy for programming languages with multisequential processes. The Journal of Systems and Software, 7(2):127–140, June 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tao:1991:FDV

[TK91] Yonglei Tao and Chenho Kung.

Formal definition and verification of data flow diagrams. The Journal of Systems and Software, 16(1):29–36, September 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tesch:1995:ISP

[TKS95] Deborah B. Tesch, Gary Klein, and Marion G. Sobol. Information system professionals’ attitudes: Development tools and concepts. The Journal of Systems and Software, 28(1):39–47, January 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tung:1993:MVE

[TKU93] Yukuh Tung, Amir A. Khwaja, and Joseph E. Urban. Multiple views of an executable software specification language. The Journal of Systems and Software, 21(3):305–319, June 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thoreson:1989:LMH

[TL89] Sharilyn A. Thoreson and Arlen N. Long. Locality, a memory hierarchy, and program restructuring in a dataflow environment. The Journal of Systems and Software, 9(4):245–252, May 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tang:1995:SLO

[TL95] S. M. Tang and R. Chow Louis. The study of letter-oriented hashing schemes. The Journal of Systems and Software, 28(2): 157–167, February 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tan:1996:CRD

[TL96] Hee Beng Kuan Tan and Tok Wang Ling. Components reuse for data-intensive business programs through an objectoriented architecture. The Journal of Systems and Software, 34 (1):3–20, July 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tan:1999:IDP

[TL99] Hee Beng Kuan Tan and Tok Wang Ling. Integrated design patterns for database applications. The Journal of Systems and Software, 47(2–3):159–172, July 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/47/2-3/
6151.pdf.

Lu:1989:SDI

[tLF89] Ming te Lu and Crumpton Farrell. Software development: an international perspective. The Journal of Systems and Software, 9(4):305–309, May 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tichy:1995:EEC

[TLPH95] Walter F. Tichy, Paul Lukowicz, Lutz Prechelt, and Ernst A. Heinz. Experimental evaluation in computer science: a quantitative study. The Journal of Systems and Software, 28(1):9– 18, January 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Troy:1997:ADD

[TM97] Douglas Troy and Robert Mc-Queen. An approach for developing domain specific CASE tools and its application to manufacturing process control. The Journal of Systems and Software, 38 (2):165–192, August 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tom:1998:ARO

[TM98] P. Ajith Tom and C. Siva Ram Murthy. Algorithms for reliabilityoriented module allocation in distributed computing systems. The Journal of Systems and Software, 40(2):125–138, February 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tomayko:1989:LLT

[Tom89] James E. Tomayko. Lessons learned teaching Ada in the context of software engineering. The Journal of Systems and Software, 10(4):281–283, November 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Torn:1990:MSA

[Tör90] Aimo A. Törn. Models of software accumulation. The Journal of Systems and Software, 12(1): 39–42, April 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Takahashi:1995:CSS

[TOYI95] Kenji Takahashi, Atsuko Oka, Shuichiro Yamamoto, and Sadahiro Isoda. A comparative study of structured and text-oriented analysis and design methodologies. The Journal of Systems and Software, 28(1):69–75, January 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tomayko:1989:SEG

[TR89] James E. Tomayko and David J. Rodjak. Is software engineering graduate-level material? The Journal of Systems and Software, 10(4):231–233, November 1989.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Treiber:1981:ITE

[Tre81] Arthur E. Treiber. Interoperability through effective information exchange standards. The Journal of Systems and Software, 2(4): 337–350, December 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tricot:1986:CA

[Tri86a] M. Tricot. Classification algorithms. The Journal of Systems and Software, 6(1–2):93, May 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tripathi:1986:DPS

[Tri86b] Satish K. Tripathi. On detecting parallelism in software. The Journal of Systems and Software, 6(1–2):133–135, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Titus:1989:FRS

[TS89] Gregory B. Titus and Allan M.

Stavely. Finding reachable states of finite-state concurrent systems. The Journal of Systems and Software, 9(4):253–272, May 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsuchiya:1985:AAD

[Tsu85] M. Tsuchiya. Availability analysis for the design of distributed processing networks. The Journal of Systems and Software, 5 (3):221–227, August 1985. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsai:1993:LMM

[TT93] Shang Rong Tsai and Lian-Jou Tsai. A logical machine monitor supporting an environment for development and execution of operating systems. The Journal of Systems and Software, 21(1): 27–39, April 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tian:1998:CMD

[TT98] Jeff Tian and Joel Troster. A comparison of measurement and defect characteristics of new and legacy software systems. The Journal of Systems and Software, 44(2):135– 146, December 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/2/6078.pdf.

Tian:1997:TSS

[TTP97] Jeff Tian, Joel Troster, and Joe Palma. Tool support for software measurement analysis and improvement. The Journal of Systems and Software, 39(2): 165–178, November 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Triantafyllos:1994:SMM

[TVK94] George Triantafyllos, Stamatis Vassiladis, and Walid Kobrosly. Software metrics for the microcode of computer systems. The Journal of Systems and Software, 26(3):221–232, September 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Triantafyllos:1995:PCI

[TVK95] George Triantafyllos, Stamatis Vassiliadis, and Walid Kobrosly. On the prediction of computer implementation faults via static error prediction models. The Journal of Systems and Software, 28(2):129–142, February 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thambu:1995:ETB

[TW95] Priyamvadha Thambu and Johnny Wong. An efficient token-based mutual exclusion algorithm in a distributed system. The Journal of Systems and Software, 28 (3):267–276, March 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tiemeyer:1998:TMA

[TW98] Matthew P. Tiemeyer and Johnny S. K. Wong. A task migration algorithm for heterogeneous distributed computing systems. The Journal of Systems and Software, 41(3): 175–188, June 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/41/3/6018.pdf.

Troy:1981:MQS

[TZ81] Douglas A. Troy and Stuart H. Zweben. Measuring the quality of structured designs. The Journal of Systems and Software, 2(2):113–120, June 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tian:1992:FPC

[TZ92] Jianhui Tian and Marvin V. Zelkowitz. A formal program complexity model and its application. The Journal of Systems and Software, 17(3):253– 266, March 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Uckan:1991:KRU

[Uck91] Yuksel Uckan. Knowledge representation using views in relational deductive data bases. The Journal of Systems and Software, 15(3):217–232, July 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Uckan:1986:OAP

[UH86] Yuksel Uckan and William D. Haseman. Optimal access path handling strategy in a database environment. The Journal of Systems and Software, 6(3):237– 249, August 1986. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Uchihira:1996:CAC

[UH96] Naoshi Uchihira and Shinichi Honiden. Compositional adjustment of concurrent programs to satisfy temporal logic constraints in MENDELS ZONE. The Journal of Systems and Software, 33 (3):207–221, June 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Urban:1994:DTO

[UhCLS94] Susan Darling Urban, Chiung hsun Chen Lai, and Sanjay Saxena. The design and translation of ORL: an object retrieval language. The Journal of Systems and Software, 24(2):187– 206, February 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ulusoy:1995:STT

[Ulu95] Özgür Ulusoy. A study of two transaction processing architectures for distributed real-time database systems. The Journal of Systems and Software, 31(2): 97–108, November 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ulusoy:1997:ENA

[Ulu97] Özgür Ulusoy. An evaluation of network access protocols for distributed real-time database systems. The Journal of Systems and Software, 37(1):49–60, April 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ulusoy:1998:TPD

[Ulu98]  Özgür Ulusoy.  Transaction

processing in distributed active real-time database systems. The Journal of Systems and Software, 42(3):247– 262, September 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/42/3/6041.pdf.

Ural:1990:SDS

[Ura90] Hasan Ural. Specifications of distributed systems in Prolog. The Journal of Systems and Software, 11(2):143–154, February 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Urban:1995:DCR

[UW95] Susan D. Urban and Anne M. Wang. The design of a constraint/rule language for an object-oriented data model. The Journal of Systems and Software, 28(3):203–224, March 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Varadharajan:1991:PNM

[Var91] Vijay Varadharajan. A Petri net model for system design and refinement. The Journal of Systems and Software, 15(3):239– 250, July 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

vonMayrhauser:1993:IPS

[vAW93] A. von Mayrhauser, K. Archie, and N. Weber. Incremental parsing for software maintenance tools. The Journal of Systems and Software, 23(3):235–243, December 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vaughn:1999:ICS

[VB99] Rayford B. Vaughn, Jr. and Julian E. Boggess III. Integration of computer security into the software engineering and computer science programs. The Journal of Systems and Software, 49(2– 3):149–153, December 30, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/26/30/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

vonStaa:1980:DPF

[vC80] Arndt von Staa and Donald D.

Cowan. Development proposal: the first step in software system construction. The Journal of Systems and Software, 1(2):107– 122, February 1980. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Verner:1997:PDY

[VC97] J. M. Verner and N. Cerpa. Prototyping: Does your view of its advantages depend on your job? The Journal of Systems and Software, 36(1):3–16, January 1997.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

vonMayrhauser:1993:SFA

[vD93] A. von Mayrhauser and Deepak Dube. Simple and fast approximations for generalized stochastic Petri nets. The Journal of Systems and Software, 21(2): 163–177, May 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

vandenBerg:1994:ADQ

[vdBK94] C. A. van den Berg and M. L. Kersten. Analysis of a dynamic query optimization technique for multijoin queries. The Journal of Systems and Software, 27(3): 233–241, December 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

vanEgmond:1989:IIS

[vEHvV89] S. van Egmond, F. C. Heeman, and J. C. van Vliet. INFORM: an interactive syntax-directed formulae editor. The Journal of Systems and Software, 9(3):169– 182, March 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Velasco:1987:MTD

[Vel87] F. R. D. Velasco. Method for test data selection. The Journal of Systems and Software, 7(2): 89–97, June 1987. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Verkamo:1989:PCD

[Ver89] A. Inkeri Verkamo. Performance comparison of distributive and Mergesort as external sorting algorithms. The Journal of Systems and Software, 10(3):187– 200, October 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vlahavas:1989:MLC

[VH89] J. Vlahavas and C. Halatsis. Lmachine: a low-cost personal sequential inference machine. The Journal of Systems and Software, 9(3):209–223, March 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Visaggio:1999:AMP

[Vis99a] Giuseppe Visaggio. Assessing the maintenance process through replicated, controlled experiments. The Journal of Systems and Software, 44(3):187– 197, January 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/44/3/6084.pdf.

Visaggio:1999:ARP

[Vis99b] Giuseppe Visaggio. Assessment of a renewal process experimented in the field. The Journal of Systems and Software, 45 (1):3–17, February 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/45/1/6091.pdf.

Vrbsky:1994:PAA

[VL94] Susan V. Vrbsky and Jane W. S. Liu. Producing approximate answers to set- and single-valued queries. The Journal of Systems and Software, 27(3):243–251, December 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vlahavas:1998:EPP

[Vla98] I. Vlahavas. Exploiting andor parallelism in Prolog: The OASys computational model and abstract architecture. The Journal of Systems and Software, 43 (1):45–57, October 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/43/1/6049.pdf.

Valett:1989:SSM

[VM89] Jon D. Valett and Frank E. Mc-Garry. A summary of software measurement experiences in the Software Engineering Laboratory. The Journal of Systems and Software, 9(2):137–148, February 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Voas:1993:SMS

[VM93] Jeffrey M. Voas and Keith W. Miller. Semantic metrics for software testability. The Journal of Systems and Software, 20 (3):207–216, March 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ventouris:1992:PAF

[VP92] K. P. Ventouris and P. E. Pintelas. A practical assessment of formal specification approaches for data abstractions. The Journal of Systems and Software, 17(2): 169–188, February 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Voas:1993:CCD

[VPM93] Jeffrey M. Voas, Jeffrey E.

Payne, and Keith W. Miller. Controversy corner: Designing programs that are less likely to hide faults. The Journal of Systems and Software, 20(1):93– 100, January 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

vanderPoel:1983:SMC

[vS83] Klaas G. van der Poel and Stephen R. Schach. Software metric for cost estimation and efficiency measurement in data processing system development. The Journal of Systems and Software, 3(3):187–191, September 1983. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

vanSlooten:1996:CIS

[vS96] Kees van Slooten and Bram Schoonhoven. Contingent information systems development. The Journal of Systems and Software, 33(2):153–161, May 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Verner:1987:MSS

[VT87] June M. Verner and Graham Tate. Model for software sizing. The Journal of Systems and Software, 7(2):173–177, June 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vrbsky:1998:STC

[VT98] S. V. Vrbsky and Saša Tomić. Satisfying timing constraints of real-time databases. The Journal of Systems and Software, 41 (1):63–73, April 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1998/41/1/6007.pdf.

Vrbsky:1999:STC

[VT99] Susan V. Vrbsky and Sasa Tomic. Satisfying temporal consistency constraints of real-time databases. The Journal of Systems and Software, 45(1):45– 60, February 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/jss/sub/
                                        1999/45/1/6094.pdf.

Veerasamy:1999:SCA

[VVS99] Jeyakesavan Veerasamy, S. Venkatesan, and Jay C. Shah. Spare capacity assignment in telecom networks using path restoration and further improvement using traffic splitting. The Journal of Systems and Software, 47 (1):27–33, May 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/47/1/6136.pdf.

Walters:1991:RSA

[Wal91] Neal Walters. Requirements specification for Ada software under DoD-STD-2167A. The Journal of Systems and Software, 15(2):173–183, May 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ward:1989:EBP

[War89] Paul T. Ward. Embedded behavior pattern languages: a contribution to a taxonomy of case languages. The Journal of Systems and Software, 9(2):109–128, February 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wetzel:1989:PPP

[WB89] Gregory F. Wetzel and William G.

Bulgren. Problems, problems, problems …. The Journal of Systems and Software, 9(4):297– 303, May 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Whitty:1990:MEP

[WBR90] Robin Whitty, Martin Bush, and Meg Russell. METKIT and the ESPRIT program. The Journal of Systems and Software, 12(3): 219–221, July 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Williford:1999:MFI

[WC99] Judson Williford and Andrew Chang. Modeling the FedEx IT division: a system dynamics approach to strategic IT planning. The Journal of Systems and Software, 46(2–3):203–211, April 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/46/2-3/
6128.pdf.

Wilde:1998:RES

[WCV^{+}98] Norman Wilde, Christopher Casey, Joe Vandeville, Gary Trio, and Dick Hotz. Reverse engineering of software threads: a design recovery technique for large multi-process systems. The Journal of Systems and Software, 43(1):11– 17, October 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/1/6046.pdf.

Wartik:1999:PRA

[WD99] Steven Wartik and Ted Davis. A phased reuse adoption model. The Journal of Systems and Software, 46(1):13–23, April 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/46/1/6112.
pdf.

Wood:1999:MMR

[WDMR99] Murray Wood, John Daly, James Miller, and Marc Roper. Multimethod research: an empirical investigation of object-oriented technology. The Journal of Systems and Software, 48(1): 13–26, August 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/29/11/37/17/
19/abstract.html;   http:/
/www.elsevier.nl/gej-ng/10/
29/11/37/17/19/article.pdf.

Weiss:1979:ESD

[Wei79] David M. Weiss. Evaluating software development by error analysis: The data from the Architecture Research Facility. The Journal of Systems and Software, 1(1):57–70, ???? 1979. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Weyuker:1999:ETI

[Wey99] Elaine J. Weyuker. Evaluation techniques for improving the quality of very large software systems in a cost-effective way. The Journal of Systems and Software, 47(2–3):97–103, July 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/47/2-3/
6144.pdf.

Wu:1996:DMH

[WFZ96] Jie Wu, Eduardo B. Fernandez, and Manxia Zhang. Design and modeling of hybrid fault-tolerant software with cost constraints. The Journal of Systems and Software, 35(2):141–149, November 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Werner:1991:IAD

[WH91a] Linda L. Werner and William E. Howden. An investigation of the applicability of data usage analysis. The Journal of Systems and Software, 15(3):205– 215, July 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wilde:1991:RTS

[WH91b] Norman Wilde and Ross Huitt. A reusable toolset for software dependency analysis. The Journal of Systems and Software, 14 (2):97–102, February 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Welzel:1997:PCS

[WH97] Dieter Welzel and Hans-Ludwig Hausen. Practical concurrent software evaluation for certification. The Journal of Systems and Software, 38(1):71–83, July 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Walczuch:1999:UIP

[WH99] Nikolaus Walczuch and Herbert Hoeger. Using individual prefixes in B^{+}-trees. The Journal of Systems and Software, 47 (1):45–51, May 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/47/1/6138.pdf.

Whale:1990:SMP

[Wha90] Geoff Whale. Software metrics and plagiarism detection. The Journal of Systems and Software, 13(2):131–138, October 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wheeler:1981:ESD

[Whe81] Thomas J. Wheeler. Embedded system design with Ada as the system design language. The Journal of Systems and Software, 2(1):11–21, February 1981. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:1989:GOP

[WHL89] H. S. Wang, S. R. Hsu, and J. C. Lin. A generalized optimal path-selection model for structural program testing. The Journal of Systems and Software, 10(1):55–63, July 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:1999:TSS

[WHMP99] W. Eric Wong, Joseph R. Horgan, Aditya P. Mathur, and Alberto Pasquini. Test set size minimization and fault detection effectiveness: a case study in a space application. The Journal of Systems and Software, 48(2):

79–89, October 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/37/23/18/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/37/23/abstract.
html.

Wick:1992:ESE

[Wic92] Michael R. Wick. Expert system explanation in retrospect: a case study in the evolution of expert system explanation. The Journal of Systems and Software, 19 (2):159–169, October 1992. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Williams:1989:CSM

[Wil89] Lloyd G. Williams. Comparative software methods: Introduction. The Journal of Systems and Software, 9(2):79–80, February 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:1999:DAM

[WJ99] Shiuh-Jeng Wang and Jinn-Ke Jan. A displacement addressing method for letter-oriented keys. The Journal of Systems and Software, 46(1):77– 88, April 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/1/6116.pdf.

Wiens:1988:EML

[WK88] Roger Wiens and Mohammad A.

Ketabchi. An efficient method lookup technique for secondary storage object-oriented systems. The Journal of Systems and Software, 8(4):275–282, September 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wilson:1994:AAA

[WKM94] Paul R. Wilson, Sheetal V.

Kakkad, and Shubhendu S. Mukherjee. Anomalies and adaptation in the analysis and development of prepaging policies. The Journal of Systems and Software, 27(2):147–153, November 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wernick:1999:SPW

[WL99] P. Wernick and M. M. Lehman.

Software process white box modelling for FEAST/1. The Journal of Systems and Software, 46(2– 3):193–201, April 15, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1999/46/2-3/6127.pdf.

Wang:1995:IEP

[WLC95] Hsiao-Hsi Wang, Pei-Ku Lu, and Ruei-Chuan Chang. An implementation of an external pager interface on BSD UNIX. The Journal of Systems and Software, 29(2):177–189, May 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wilkening:1995:RAS

[WLPL95] Daniel E. Wilkening, Joseph P. Loyall, Marc J. Pitarys, and Kenneth Littlejohn. A reuse approach to software reengineering. The Journal of Systems and Software, 30(1–2):117–125, July/August 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lam:1997:OCC

[wLyLlH97] Kwok wa Lam, Kam yiu Lam, and Sheung lun Hung. Optimistic concurrency control protocol for real-time databases. The Journal of Systems and Software, 38(2):119–131, August 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wernhart:1990:HEB

[WM90] Heidemarie Wernhart and Roland Mittermeir. The HIBOL-2 environment: a basis for visual programming of business objects. The Journal of Systems and Software, 12(2):157–165, May 1990. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:1995:RCM

[WM95] W. Eric Wong and Aditya P. Mathur. Reducing the cost of mutation testing: an empirical study. The Journal of Systems and Software, 31(3):185–196, December 1995. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:1996:NTA

[WM96] Johnny S. K. Wong and Soma Mitra. A nonblocking timed atomic commit protocol for distributed real-time database systems. The Journal of Systems and Software, 34(2):161– 170, August 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:1999:IMA

[WM99] Johnny S. K. Wong and Armin R. Mikler. Intelligent mobile agents in large distributed autonomous cooperative systems. The Journal of Systems and Software, 47(2–3):75–87, July 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.com/cas/tree/
store/jss/sub/1999/47/2-3/
6142.pdf.

Woodside:1986:SMP

[WNHM86] C. M. Woodside, E. Neron, E. D.-S. Ho, and B. Mondoux. ‘active server’ model for the performance of parallel programs written using rendezvous. The Journal of Systems and Software, 6(1–2):125–131, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Warren:1996:EES

[WNSC96] James R. Warren, A. F. Norcio, Jack W. Stott, and G. C. Canfield. Experimental evaluation of a simulation environment for information systems design. The Journal of Systems and Software, 32(1):3–20, January 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:1993:DCP

[Won93] Kam-Fai Wong. Design considerations of a parallel recovery scheme for the EDS data base server. The Journal of Systems and Software, 21(1):41–48, April 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Woodside:1980:MME

[Woo80] C. M. Woodside. A mathematical model for the evolution of software. The Journal of Systems and Software, 1(4):337–345, ???? 1980. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wohlin:1999:SIR

[WR99] Claes Wohlin and Björn Regnell. Strategies for industrial relevance in software engineering education. The Journal of Systems and Software, 49(2–3):125– 134, December 30, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/29/
11/46/26/27/article.pdf;
http://www.elsevier.nl/gej-
ng/10/29/11/46/26/abstract.
html.

Wessale:1993:LPE

[WRW93] W. Wessale, D. Reifer, and D. Weller. Large project experiences with object-oriented methods and reuse. The Journal of Systems and Software, 23(2): 151–161, November 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Woodfield:1981:SSM

[WSD81] S. N. Woodfield, V. Y. Shen, and H. E. Dunsmore. Study of several metrics for programming effort. The Journal of Systems and Software, 2(2):97–103, June 1981. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Welch:1995:RCB

[WSM^{+}95] Lonnie R. Welch, Antonio L.

Samuel, Michael W. Masters, Robert D. Harrison, Mark Wilson, and Joe Caruso. Reengineering computer-based systems for enhanced concurrency and layering. The Journal of Systems and Software, 30(1–2):45–70, July/ August 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Warren:1992:SSI

[WSN92] James R. Warren, Jack W. Stott, and A. F. Norcio. Stochastic simulation of information systems designs from data flow diagrams. The Journal of Systems and Software, 18(2):191– 199, May 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wileden:1983:BSS

[WSR^{+}83] Jack C. Wileden, John H. Sayler, William E. Riddle, Alan R. Segal, and Allan M. Stavely. Behavior specification in a software design system. The Journal of Systems and Software, 3(2):123– 135, June 1983. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wilde:1989:MAW

[WT89] Norman Wilde and Stephen M. Thebaut. The maintenance assistant: Work in progress. The Journal of Systems and Software, 9(1):3–17, January 1989. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:1995:FTA

[WTS95] Johnny Wong, Priyamvadha Thambu, and Rick Stoen. A fault-tolerant algorithm for mutual exclusion in a distributed system. The Journal of Systems and Software, 29(2):121– 134, May 1995. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See critical analysis [JM96].

Wang:1997:CAS

[WWC97] Yi-Min Wang, Hsiao-Hsi Wang, and Ruei-Chuan Chang. Clustered affinity scheduling on largescale NUMA multiprocessors. The Journal of Systems and Software, 39(1):61–70, October 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:1998:CAC

[WWC98] Yi-Min Wang, Hsiao-Hsi Wang, and Ruei-Chuan Chang. Classifying and alleviating the communication overheads in matrix computations on largescale NUMA multiprocessors. The Journal of Systems and Software, 44(1):17–29, December 1, 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/1/6068.pdf.

Wu:1994:UAS

[WWF94] J. Wu, Y. Wang, and E. B. Fernandez. A uniform approach to software and hardware fault tolerance. The Journal of Systems and Software, 26(2):117– 127, August 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Xie:1998:SSS

[XH98] M. Xie and G. Y. Hong. A study of the sensitivity of software release time. The Journal of Systems and Software, 44(2): 163–168, December 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/44/2/6081.pdf.

Xie:1999:SRP

[XHW99] M. Xie, G. Y. Hong, and C. Wohlin. Software reliability prediction incorporating information from a similar project.

The Journal of Systems and Software, 49(1):43–48, December 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/29/
                                        11/46/25/29/article.pdf;
                                        http://www.elsevier.nl/gej-
                                        ng/10/29/11/46/25/abstract.
                                        html.

Yang:1994:HMP

[Yan94] Wuu Yang. How to merge program texts. The Journal of Systems and Software, 27(2): 129–135, November 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yung-Chen:1992:LCS

[YCGH92] Hung Yung-Chen and Chen Gen-Huey. A layered communication system generator. The Journal of Systems and Software, 18(2): 157–170, May 1992. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yuen:1996:BSL

[YFY96] C. K. Yuen, M. D. Feng, and J. J. Yee. BaLinda suite of languages and implementations. The Journal of Systems and Software, 32 (3):251–267, March 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lam:1998:GEC

[yL98] Kam yiu Lam. Guest Editor’s corner. real-time active databases: Theory and practice. The Journal of Systems and Software, 42(3):189–191, September 1998. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Lam:1998:USC

[yLcY98] Kam yiu Lam and Wai cheong Yau. On using similarity for concurrency control in real-time database systems. The Journal of Systems and Software, 43(3): 223–232, November 1, 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.elsevier.
com/cas/tree/store/jss/sub/
1998/43/3/6063.pdf.

Yellen:1991:IWN

[YN91] Richard E. Yellen and Jay F.

Nunamaker. Increasing the willingness of novices to use computer application software. The Journal of Systems and Software, 15(3):279–285, July 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yu:1988:SIS

[YNDS88] T. J. Yu, B. A. Nejmeh, H. E.

Dunsmore, and V. Y. Shen. Smdc: an interactive software metrics data collection and analysis system. The Journal of Systems and Software, 8(1):39– 46, January 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yong:1994:CRR

[YP94] Deng Yong and E. K. Park.

Checkpointing and rollbackrecovery algorithms in distributed systems. The Journal of Systems and Software, 25(1):

59–71, April 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yau:1980:ATD

[YRN80] Stephen S. Yau, John L. Ramey, and Robin A. Nicholl. Assertion techniques for dynamic monitoring of linear list data structures. The Journal of Systems and Software, 1(4):319–336, ???? 1980. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yuasa:1990:RTG

[Yua90] Taiichi Yuasa. Real-time garbage collection on general-purpose machines. The Journal of Systems and Software, 11(3):181– 198, March 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yee:1993:TBE

[YY93] Jenn-Jong Yee and Chung-Kwong Yuen. Transputer-based emulation of a data-driven LISP machine: BIDDLE. The Journal of Systems and Software, 23 (1):51–63, October 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zelkowitz:1997:AIT

[ZC97] Marvin Zelkowitz and Barbara Cuthill. Application of an information technology model to software engineering environments. The Journal of Systems and Software, 37(1):27–40, April 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:1996:DMR

[ZCd96] Lichen Zhang and B. Chaibdraa. A design methodology for real-time systems to be implemented on multiprocessor machines. The Journal of Systems and Software, 33(1):37–56, April 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zaki:1988:ARM

[ZEB88] M. Zaki and M. M. El-Boraey.

Analysis of reliability models for interconnecting MIMD systems. The Journal of Systems and Software, 8(2):133–144, March 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zeil:1988:CET

[Zei88] Steven J. Zeil. Complexity of the equate testing strategy. The Journal of Systems and Software, 8(2):91–104, March 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zelkowitz:1988:RUD

[Zel88] Marvin V. Zelkowitz. Resource utilization during software development. The Journal of Systems and Software, 8(4):331–336, September 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zelkowitz:1996:MSE

[Zel96] Marvin V. Zelkowitz. Modeling software engineering environment capabilities. The Journal of Systems and Software, 35(1): 3–14, October 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zaki:1993:DID

[ZENA93] M. Zaki, M. Y. El-Nahas, and H. A. Allam. DPDP: An interactive debugger for parallel and distributed processing. The Journal of Systems and Software, 22(1):45–61, July 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhou:1997:FTS

[ZG97] Wanlei Zhou and Andrzej Goscinski. Fault-tolerant servers for the RHODOS system. The Journal of Systems and Software, 37(3):201–214, June 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhou:1993:DID

[Zho93] Wanlei Zhou. The design and implementation of a distributed program monitor. The Journal of Systems and Software, 22(1): 63–77, July 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhou:1994:RPS

[Zho94] Wanlei Zhou. A rapid prototyping system for distributed information system applications. The Journal of Systems and Software, 24(1):3–29, January 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zimmerman:1984:PMT

[Zim84] Rodney P. Zimmerman. Phases, methods, and tools: a triad of system development. The Journal of Systems and Software, 4 (2–3):207–218, July 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zaki:1985:MPD

[ZK85] M. Zaki and S. Koriem. A model for the protocol design of back-end storage networks. The Journal of Systems and Software, 5(4):249–265, November 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhu:1996:HPB

[ZLZ^{+}96] Jiang Zhu, T. G. Lewis, Lihua Zhao, Weldon Jackson, and Russel L. Wilson. HaRTS: Performance-based design of distributed hard real-time software. The Journal of Systems and Software, 32(2):143–156, February 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhou:1996:SMR

[ZM96] Wanlei Zhou and Brian Molinari. A system for managing remote procedure call transactions. The Journal of Systems and Software, 34(2):133– 149, August 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zaki:1999:TPS

[ZMAER99] M. Zaki, Y. B. Mahdy, S. A. Ali, and K. F. El-Reefy. Two practical systems for classification of three-dimensional objects. The Journal of Systems and Software, 48(1):59– 71, August 1, 1999. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/29/11/37/17/
                                        22/abstract.html;   http:/
                                        /www.elsevier.nl/gej-ng/10/
                                        29/11/37/17/22/article.pdf.

Zhao:1987:SIH

[ZR87] Wei Zhao and Krithi Ramamritham. Simple and integrated heuristic algorithms for scheduling tasks with time and resource constraints. The Journal of Systems and Software, 7(3):195–205, September 1987. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhou:1994:VFD

[ZR94] H. B. Zhou and L. Richter. Very fast distributed spreadsheet computing. The Journal of Systems and Software, 25(2):185– 192, May 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zaki:1988:LVI

[ZS88] M. Zaki and R. Salama. Linkage versus integration for binding data base and interactive graphics systems. The Journal of Systems and Software, 8(5):361–372, December 1988. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zand:1995:GEC

[ZS95] Mansour Zand and Mansur Samadzadeh. Guest Editors’ corner: Software reuse: Current status and trends. The Journal of Systems and Software, 30(3): 167–170, September 1995. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zand:1993:ILR

[ZSGS93] M. K. Zand, H. Saiedian, K. M.

George, and M. H. Samadzadeh. An interconnection language for reuse at the template/module level. The Journal of Systems and Software, 23(1):9–26, October 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zucconi:1990:CCDb

[Zuc90a] Lin Zucconi. Controversy corner: a debate, Zucconi vs. Reifer: Response to “Is Asia Really a Threat?”. The Journal of Systems and Software, 12(1): 77, April 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See also [Zuc90b, Rei90b].

Zucconi:1990:CCDa

[Zuc90b] Lin Zucconi. Controversy corner: a debate, Zucconi vs. Reifer: U.S. technology: We’re losing the edge in software. The Journal of Systems and Software, 12(1): 71–74, April 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See also [Zuc90a, Rei90b].

Zviran:1993:CMC

[Zvi93] Moshe Zviran. A comprehensive methodology for computer family selection. The Journal of Systems and Software, 22(1): 17–26, July 1993. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zweben:1990:RSS

[Zwe90] Stuart H. Zweben. On the resolution of the software science anomaly. The Journal of Systems and Software, 12(2):167– 170, May 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zupancic:1996:GEC

[ZWM96] Jože Zupančič, Stanisław Wrycza, and Ali Mili. Guest Editors’ corner. The Journal of Systems and Software, 33(2):111– 112, May 1996. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhuang:1994:DAS

[ZX94] W. J. Zhuang and M. Xie. Design and analysis of some faulttolerance configurations based on a multipath principle. The Journal of Systems and Software, 25 (1):101–108, April 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhou:1988:OML

[ZZ88] Songnian Zhou and Roberto Zicari. Object management in local distributed systems. The Journal of Systems and Software, 8(4): 283–295, September 1988. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).